A Complete Bibliography of Publications in Mathematics of Computation, 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 January 2025 Version 2.30

Title word cross-reference

#5578 [Wil79]. #8373 [BW96].

(a^{n} − 1)/(a − 1) [WS79a]. (modp) [Pol78]. (tu_{t})_{t} = u_{xx} [SS70]. +1 [LV79]. −1 [LV79]. $0.65 [P.71]. $0.65. [Rhe72]. 1, 000, 000 [Dut71]. 10^{n} ± 1 [AG74]. 12, 758 [DP74]. 125000 [Wag78]. 2 [BWW77, Dut71, JS78, RC78, Sha71c]. 2 − 10^{−12} < σ(N)/N < 2 + 10^{−12} [Kis78]. 200 < p < 521 [LM78]. 2^{m} ± 1 [BLS75]. 2A3^{n} + 1 [WZ72]. 2A3^{n} − 1 [WZ72]. 2 · 3 · 5 ⋯ p ± 1 [Bor72b]. 2m [BS71]. 2ξ(s)/s(s − 1) [Wal78]. 3 [DSW79, NS74a, SS73]. $3.00 [P.71, Rhe72]. 3 · 2^{n} + 1 [Gol76b]. 3x + 1 [Cra78]. 4 [DSW79]. $4.00 [I.75]. 4n ± 1 [BH78b]. _{2}F_{2}(1, α; ρ_{1}, ρ_{2}; z) [Kim72]. A [CS79, Dem79, Ise78, Jai72, Jel76, NS74b, Wd74]. A^{6} + B^{6} + C^{6} = D^{6} + E^{6} + F^{6} [Bru70a, Sha71a]. α [LV79]. Ax = λBx [Ste75]. Ax = y [DM70]. B [dD74]. \binom{2n}{n} [EGRS75]. \binom{n}{k} [EES74]. C^{1} [MS75]. χ^{2} [Fet79]. cosh z = az [Fet76]. E [EM78, Ern79]. e^{−t^{2}} [BKV72]. e^{x} − x − 1 [How77]. ϵ [Kah72b]. ≡ 1(mod 7) [Wil74]. F [AB72]. F^{7} [MB75]. F^{m}(ax) [Apo75]. F^{n}(bx) [Apo75]. F_{5} [Lon77]. \frac{∂ρ}{∂t} + \frac{∂}{∂x}(υ(ρ)ρ) = 0 [IN79]. G [MY79, Fie72, ABH73]. G, |G| < 10^{6} [FM78]. γ [Sha76a]. GL(n, Z) [PP77a, PP77b]. H^{2} [Win70]. h^{∗}(p) [LM78]. HNN [Ans76]. I_{0}(x) [Bla74]. I_{1}(x) [Bla74]. I_{MN} [ZE78]. ∫_{0}^{1}(−Logx) · x^{α} · f(x) · dx [Dan73]. ∫_{0}^{∞} E_{m}(x) · f(x) · dx [Dan73].

∫_{0}^{∞} ^{\frac{t^{2α−1}J_{ν}(x\sqrt{1+t^{2}})}{(1+t^{2})^{α+β−1}}}dt [Sch78b]. ∫_{0}^{∞} t^{2α−1}(1 + t^{2})^{1−α−β}J_{ν}(x\sqrt{1 + t^{2}})dt [Gla79b]. ∫_{0}^{∞} x^{α}e^{−x}f(x)dx [LV79]. ∫_{x}^{∞} e^{−t^{2}/2}t^{p}dt, x > 0, p [DiD78]. J_{−ν}(x) [CK79]. k! ± 1 [Bor72b]. K [Lau73, Wag72]. K_{n}(Z) [Par72, LG79]. K_{ν}(t)/I_{ν}(t) [LL72]. k · 2^{n} + 1 [Bai79, MW77]. Ki_{n}(x) [BEJ78a, BEJ78b]. L [LO79, Wag75, Wei75a, Bue77, Ger75]. L^{1} [Abd75]. L^{2} [de 76, Nic77]. L^{∞} [de 76, Sco76]. L^{p} [Wat73, Shr74]. L_{+}1 [AG79b]. l_{1} [FB78, Lew70]. L_{2} [CSW77]. l_{∞} [LR73, LR79a, DDW75]. L_{p} [Kah72c, Wol75]. [0, 1] [Ros71a]. [α, ∞) [Dun75a].

LDL^{T} [FP74]. LDV [GMS75]. log n [KR79]. M [McC78, Poo75, Mv77, Phi70, Wil76b]. M(23) [Hun74b]. M_{k,1/2} [Lau73]. GF(q) [Bea74, BW74a, BW76]. GF[2] [Mos72]. GF[q, x] [Web73]. IM(X) = 1 [Wun73]. QR [War76]. SOR [Ruh74]. μp [Joh73]. N [Kis78, Lyn76, WJ76a, Hud71, Per70]. N^{2} ± 1 [WJ76a]. n ≥ 5 [MS75]. ν [CH70a, CK79]. div υ = 0 [Fal76b]. p [Joh78, Wag75]. p < 8000 [Joh73]. ϕ(n) = ϕ(n + k) [LG72]. π [CDR71, Sal76a, Sha76a]. π_{3,2}(x) < π_{3,1}(x) [BH78a]. ψ(x) [Sch76b, RS75b]. Q(−p)^{1/2} [Coh79a]. Q(\sqrt{D}) [WB79]. Q(ζ^{16}) [Oja77]. QR [Rei71]. R^{1} [BR79]. R^{2} [Erd73]. R^{n} [Kea78]. S [Mor73, de 79]. sin z = az, cos z = az [Fet76]. ∑_{1}^{N}l/i [BE75]. t [KMR77]. T^{+m} [Str76a]. T_{+m} [Str76b]. τ [Spi73]. θ(x) [Sch76b, RS75b]. u^{t} + u^{x} = 0 [Hed75]. W_{k,1/2} [Lau73]. W_{k,m}(z) [WR71, Won73]. ℘ [Eck76]. x [BH78a, Mad72]. x^{3} + Dy^{3} = 1 [WH77]. x^{n} [McC77]. x^{n} − 1 [BW74a]. y [Mad72]. y′′ = f(x, y) [CM78, Jel78, Mic73]. Y^{2} + k = X^{5} [Bla76b]. ∥A^{−1}∥_{∞} [Dem79]. |G| < 10^{6} [MY79].

-Adic [Wag75]. -Algorithm [Kah72b]. -Bases [RC78]. -Class [Ger75]. -dimensional [de 79, Lyn76]. -Dirichlet [Spi73]. -distribution [Fet79]. -Divisibility [Joh78]. -Free [Wag72]. -Function [Eck76, Mor73, Fie72]. -Functions [Wag75, Wei75a, LO79, Bue77]. -Irregular [EM78, Ern79]. -Matrices [Poo75]. -Matrix [Mv77]. -Methods [Ruh74]. -norm [de 76]. -Rank [DSW79, NS74a]. -Sequences [Wil76b]. -Spline [dD74]. -Splines [ABH73]. -Stability [Ise78, Jel76]. -Stable [CS79, Wd74, Jai72, NS74b]. -Sylow [JS78, Sha71c]. -th [Phi70]. -Zeros [CH70a, CK79].

16 [Wim72a].

2 [S.78]. 2-Rank [EFO78]. 24 [Lun72, Sha71a, Sha71b]. 25 [BW96, Sha76a]. 27 [LR79a]. 29 [Hil76, Sha75b]. 2k [Hun75a]. 2n [Rio75]. 2nd [RT77, S.75].

31 [Gay79, Wil79]. 32 [BEJ78a, Ern79, God79].

46 [BW96].

50th [Gau94]. 55 [Wil79].

A. [P.71, Rhe72]. Abel [Bru74a, Wei72b]. Abramowitz [Str71]. Abscissa [Hud71]. Abscissas [PB74b, PB74a]. Absorbing [EM77]. Academic [Mac72b]. Accelerated [BR76, Had78]. Accelerating [JH70, SJR76]. acceleration [Ehr72]. accuracy [AG73, BS77, OD78, PR74]. Accurate [Cho73, CS73, Per70, RP70, ZA71]. Acknowledgement [Ano77a]. Adams [Hil76, Hil75]. Adaptive [Bra77, Hab75, SW79d]. added [Duf74]. Addendum [Ano75a, BEJ78a]. Addition [Ans76, Fet72, PB74b, Ans78, PB74a]. Adic [Wag75]. Adjusted [Lyn71a]. adjusting [Har78]. Advanced [Wd74]. Aerospace [Fet71a, Fet71b]. affine [Lyn79a]. Air [Fet71a, Fet71b]. Airy [BPD79, CH78, HB75b]. algebra [BK74, I.72]. Algebraic [But72a, Gol72, Lyn71c, NGF74, Tur75, CGZ72, PZ77, SR76, Wan76]. Algebraic-Numerical [NGF74]. algebras [Lon77]. Algorithm [Alo76, Atk75, Bas72, Boy73, Bro71, CDHW73, Cli72, Dun75b, GR79, Hen74, IN79, Kah72b, KLT78, Kyu79, New71, Nic77, PS79a, Rei71, Sco79, Ste71d, Swa79, War76, WB79, Bog74, BR74, CGZ72, Car72, Dun75c, EW76, Fet74, Gek72, How73a, How73b, Hud74, Lel75, McC77, Oja77, Pet73, SR76, Vau78, Wan78, Wil76a]. Algorithms [AG79b, BM70, Bra71, DGKS76, KD71, Kro70, LR73, Lyn79b, MPR77, OS72, Ore73, Ris73, Ruh79, Sam71, WJ76b, AG73, Hen77b, LR79a, Lyn79a, Moe77, Bre75c]. Aliasing [SF74]. Aliquot [Erd76, GS75]. Almost [Sal71a, Sal73]. Almost-Interpolatory [Sal73]. Alston [I.75]. Alternatives [Pru79]. Amicable [BLM70, Coh70, Hag70, Hag71, HL77, te 74, Bor72a]. Analog [Osl72a, Sim72]. Analogue [Osl72b, Sik79]. Analogues [Ste71b]. Analysis [BR79, BJM77, CDHW73, CP70, DDW78, KT78, New73, New74, P.71, Pai73, Pai79, Pau72, Ram71, Sni76a, Ste71d, BN70, BW74c, BD76, CH78, CR78, Fal76a, GK73, Hic78, NO76, Phi73, Skö79, I.75, Tha77]. Analytic [Cha73, Let73, RR70c, SP75]. Angles [BG73, RS75a]. Angstl [Bau77]. Anharmonic [Pau72]. Anniversary [Gau94]. Anomalous [Gau77]. any [JH78]. applicability [FHU77]. Application [Bel74, Bon71, Hou77, Ric71, Ris73, SF74, Als74, BK74, BR74, CR78, DM74, Kel75, LG79, Lew75, Wei74b, de 79, dW77]. Applications [Hen74, Hil73, Lyn71a, Poo75, Sha75a, Tur73, FT77, PS73, Sha75b, Sni76b, Usm78, Wil76a]. Applied [BMN70a, Nic78b, Skö79, Tha77]. Approach [Cha73, Mil75a, RV73]. Approximants [Chi73, CSW77, Sid77, ZE78]. Approximate [Mic73, GK73, NW74, SW79b, Ser75, WZ78]. Approximating [Hil73, BDT72, Pru73]. Approximation [BM70, Bel74, BKV72, Cus71, DiD78, Dun75a, Dun75b, Eck76, FB78, Gek74, Kah72c, KLT78, Kel75, Kut71, Lar70, LR73, Lew70, Nin70b, Osb75, Ros71a, Shr74, SF74, Wei72a, Wol75, Abd75, BDM78, BCSW78, Bel72, BK74, Bra73a, BS78, Bur78, Coh73, Cus72, Dun76, EW76, EF79, Fal74, Gir76, Gro75, Hen77b, JT75, Ker72, Kol79, LR79a, LG79, Lew75, Lus79, Net73, PS73, Sch78a, TW74a, Wil76a, Wil72, de 76]. Approximations [Car75b, CDR71, CH70b, CPT70, CHT71, CST73, Der77, DDW75, DLT73, EP75, FS71, For73, Hil73, JH70, Kra78, Lew72, Luk70, Nin70a, Ple72, Red76, Var70b, Win70, BBT77, Bey79, Bla74, Bla76a, BEJ76, BEJ78a, BEJ78b, Bow76, BO73, FT73, GA74, GT78, GKS72, Gus75, KT78, Kre72a, Mad72, Nev78, Osh72, Sha76a, Ste76, Tho77, Tho79, Wat73, Wei75b]. Approximative [Tur75]. Arbitrary [EP75]. arctangents [Car72]. Area [Kel79]. Argument [Mor79]. Arguments [Gre70b]. Arising [Cha73, Sac71, Var72, MM74, Sin72]. Arithmetic [Coh79a, GR79, GH79, Sal76a, Wei77, How73a, How73b, Tun75, Wag79a]. Arithmetic-Geometric [Sal76a]. arrangement [BH77]. artificial [Har78, Hic78, Hof79]. Artin [LO79]. Aspects [I.72, Ruh79, Nic79]. Associated [EES74, Ell71, Sam74, Wim62, LG74, Wim72a]. associativity [BvT78]. Asymmetric [Fia71]. Asymptotic [Ell71, Erd76, FT77, Kim72, Lyn71a, New76, Oka76, RR70a, Won73, dW73, CC72, Fet79, Fie72, Mac72a, Sch78a, SP75, Tem75]. Asymptotically [DM71, AO74]. attempt [SW79c]. August [Gau94]. Authors [P.71]. autonomous [Ral76]. Available [P.71, Rhe72]. averages [And75]. averaging [BS77, MW76]. avoid [SW79c].

B [BW96, I.72, Sin72]. B-splines [Sin72]. B. [KS79]. Back [Bur71b]. Backtrack [Knu75]. Backward [OS72, Tha79]. Bairstow [Bro75, Car73]. Banach [BS78]. Band [HP72, Ruh79, Tre74]. Banded [Dem79]. Base [Fet71a, Fet71b]. Based [ABH73, Hic78, Lyn74]. Bases [BMN70a, BMN70b, RC78, Sha71b, Var77]. basic [Alf79c]. Basis [MS75]. Bateman [SL72]. Bearings [Cry71]. behavior [FT77]. Ben [Bog76]. Ben-Israel [Bog76]. Bench [Lyn79b]. Bergman [Bur70, Bur71a]. Bernoulli [Cha79, Fra79, Joh74b]. Bessel [Wim72a, Amo74, BC79, BK78, Bla74, Gab79, GS78, Gla74b, LG79, Lin72, LK73, Lir71, Luk71a, Luk71b, Luk72, Nas74, Par72, PS77b, PS79b, Tha79, Wim62]. Best [CSW77, Kah72c, Lew70, Lew72, Nin70a, Win70, BCSW78, Wat73, WZ78]. beta [Tem75]. Between [Bre74, Cad71, Hud71, LV79, Ada79, BG73, Bre73, Tun75]. Beyer [Sha71b]. bicharacteristics [JP72]. Bickley [BEJ78a, BEJ78b]. Bicyclic [BWW77]. Bidiagonal [Cha74]. Bieberbach [Ros71b]. Bifurcation [Wei75b]. biharmonic [BR72]. bilinear [FHP78, Sch74]. Billevich [SR76]. Binary [Sta73a, Yan76]. Binomial [Mak70, Tak78]. Biquadrate [RB73]. Biquadrates [RB73]. Biquadratic [BWW77]. Bisecting [RS75a]. Bisection [Kea78, Sty79]. bisections [Sik79]. Bivariate [Dre78, GW78, Sal71b]. Block [BR72, Ris73, Var72, Wat78]. Block-Tridiagonal [Var72]. Blocks [Atk75]. Blue [Gau79b]. Book [Abr74, I.72, I.75, P.71, Rhe72, S.75, S.78, Tha77]. Books [Ano71e, Ano71f, Ano72e, Ano75d, Ano75e, Ano76d]. Bound [Hag73, Het72, Kea78, RS75a, Tuc73, de 76]. Boundaries [HS75a]. Boundary [BS71, BT72, Bra77, CP70, CS71, Cry71, DDW75, Hos73, Per70, Rob71, Sac71, Var70b, dS77, AG79a, AI75, Als74, ACR79, BG75, Bay78, Bey79, Bla76a, BDT72, BR74, EM77, FO77, Gir76, GA74, Gol77a, GT78, Gun77, GKS72, Gus75, Joh74a, Kel75, Kre72a, LP74, MM74, Oli74, Oli76, Osh72, Pru73, RV75, SW79b, Ser75, Skö79, Ste79, Ste76, Usm78, Var74, Wei74b, Wei75b, Zla75]. Boundary-Value [BS71, Bra77, CP70, Per70, Rob71, Var70b, BDT72, BR74, Oli74, Oli76, Wei74b]. Bounded [Fra71b]. Bounding [Dem79]. Bounds [Amo73, Boy70, Car70, FM71, Hag70, Het73, Let70, Let73, Mad73, RS75b, Sch70d, Wal72, WCJ72, Kah72a, Sch76b, Tun75]. British [Gau94]. Brouwer [Erd73]. Broyden [Den71, MT76]. Brudno [Sha71a]. Brun [SW74]. Bur [Str71]. Butcher [Nos73].

C [Fet71a, Fet71b, Sha76a, Str71]. Calculate [Die75]. Calculating [BK77, WZ78]. Calculation [BG70, Gus79, Lyn70, Lyn71b, Lyn71c, Sha75a, Spi71, Spi73, WB79, PWKW79, Wat73, Sha75b]. Calculations [RG70]. Calculators [Der77]. Can [LS72]. Capacitance [OW79, PW76]. Carasso [MS73b]. Cardinal [Kre72b, MSW71]. Carlo [Oka76, PWKW79, Pol78]. Cartesian [Zar76]. Case [Nic78a, PP77b, DC79, Hic78, Joh77]. Cases [PP77a]. Casilinear [Per70]. Caslin [Fet71a, Fet71b]. Catalan [CS73, te 73]. Catalogue [BCH73]. Cauchy [EP79, EF79, GB79]. CDS [Alf79a]. century [Gau94]. Certain [GH72, Ker70, LM70, Osl75, Rob75, Shr74, Var72, Wil77, Wil79, AL77, BC79, Bay78, CC72, HM77, Lak74, Lak75, Lun72, Osh72, SS72, Sty79, TE72]. Certaine [SF74]. Chains [Gup71, Gen72]. Chambers [BB74, Cox72]. changing [GW79]. Character [Hun74b, LLS70, Hun74a]. Characteristic [Wil76b, McN73]. Characterization [Jel78, DM74]. Characters [Hud73, McK76]. Charlier [BG70]. Chebyshev [BEJ78a, Abd77, BDM78, Bas70, Bla74, BEJ76, BEJ78b, Bog74, Bra73a, CH70b, CPT70, CHT71, CST73, CR72, Dun75a, Dun75b, Ehr72, EW76, GY74, Gau75, Hay73, Kah70, Kam73, MH77, NRRZ71, Nin70b, RR70b, Ros71a, RS75b, Sag70, Sal73, Sal76b, Sch76b, Sch78d, TW74a, Wil72]. Chebyshev-Type [GY74, Gau75, RR70b]. checking [Bau77]. chi [WB76]. choice [Cul79]. Choong [Sha76a]. chords [Rio75]. Chowla [LF71]. Christoffel [Gal71]. Christopherson [Cry71]. circle [Rio75, Yan79]. Circulant [Yan71, HM77]. Class [BR76, Bue76, BWW77, Coh79a, Ger75, GH72, Hen74, HP72, Kum74, Leh77, LM78, Lir71, McA71, Par78, Sha76b, Sta75, Wd74, Wil77, WS79b, AL77, Alf79c, BWB76, Bue77, But72b, Fal74, Gir76, Gri78, GK73, Hen75, Lak74, Lak75, Par77, SPB75, WB76, Wil79]. Class-Number [BWW77, Sta75, Wil77, WS79b, Wil79]. Classes [Cho71, Coh77, DMP74, TE72]. classical [Hud76, SS72]. Cliffs [Abr74, Bre75c]. Closed [GW71, Lel75, Pet73]. Clough [Cia78, Man78]. cm [I.75, P.71, Rhe72]. cm. [Abr74, S.75]. Co [S.75]. code [Mar75, Sha78]. Codes [Alt71]. Coefficient [Lyn71a, Mv77]. Coefficients [BG70, Der77, Lam70, Lyn70, Lyn71b, Lyn71c, Mak70, RR70c, Ros71b, Tak78, TE72, dD74, BO78, DD75a, FNC75, Pru73, Sch76a, Sch77]. Coerror [Amo73]. Collocation [Hen77a, AI75, ACR79, Bru74b, CR78, DD73, Hou77, HP78, LR79b, Sin72, Wei74b, Día79]. collocation-Galerkin [Día79]. Columbia [Gau94]. Combinatorial [BCH73, Str75, Dum79]. Combined [KLT78]. Comments [dS77]. Commerce [P.71, Rhe72]. Common [BK78]. Comp [BW96, BEJ78a, Ern79, Gay79, God79, Hil76, LR79a, Lun72, Sha71a, Sha71b, Sha75b, Sha76a, Wil79, Wim72a]. Compact [Osb75, Wei74a, CL75, CL78, Slo76]. Comparing [SW71, Hil75, Hil76]. Comparison [Hen77b, LR73, dS77, For72, Gal78, HKK75, LR79a, RV75, Var74]. Complementary [FCC73b, FCC73a, HR72]. Complete [Jel78]. completely [RH72]. Complex [FCC73a, Fet76, FK75, Kah71, Kol70, Kol72a, Par72, Sta75, Tha77, BDM78, Cio79, EW76, Lak74, Lak75, Let77, LG79]. complex-valued [BDM78]. complexity [PS77c]. Components [Sha73b]. Compound [FM71, RV72]. compression [Har78]. Computation [AMS84, AB72, Amo74, Bre77, Dre78, EFO78, ET73, Lak74, Lak75, LV79, Lew70, Lew72, Lin70, MR71, ND70, Pol78, Rei71, Ric71, Ros71b, Ruh79, Sal76a, WH77, BW74c, BC74, BC73, Chi72, GS78, Gau94, Har78, LK73, Mac73]. Computational [Gus70, KS78, Lyn74, Mor73, PS77c, Ser75, Tha77, BWB76, Gau94, WB76]. Computations [Jer77]. Compute [Gaf78]. Computer [KMR77, Pai79, Pex70, Sam71, SS78, KL79, PS73]. Computer-Generated [SS78]. Computers [War76]. Computing [Abd77, Bea74, BG73, Dix70, Erd73, Fet70, GMS75, KL79, Lin72, MPR77, Osl72b, SW73, Var70a, Win78, Car72, Fet74, Gab79, Gen72, Kro79, LO79, Pag73, PZ77, Ral76]. Conception [Le 77]. Concerning [JR70, RP71, Eil72, KS79, Sch74, Wre73]. Condition [Gau72, Gau79a, Hou72, Jel76, Bur78, Cio79, Gir76, Mah75, Var77]. conditional [Hud74]. Conditioned [Var70a]. Conditioning [Dor71, Rhe76, Sha70, SK70]. conditions [Bla76a, EM77, GA74, Ste76]. cones [Wil76a]. Confluent [Des74, SP75]. Conformal [Bur70, CA79, HKK72, Gut77, HKK75]. Congruence [Cho71, New75b, DMP74, Wag79b]. congruences [Joh77]. Congruent [AC74, God78, God79]. Congruential [BRW71, Nie72, BW96, Nie74, Nie76]. Conjecture [JR70, Mak70, Mal71, Ros71b, Ste71a, Zwi79, te 73, CS73]. conjectures [KS79]. Conjugate [BR76, Naz76]. Connected [Bur70, Bur71a]. Consecutive [Cad71, RJ70]. Conservation [ZA71, AG73, Hof79]. conserving [GP79]. consideration [OD78]. Consisting [GH79]. Constant [Boy70, SW74, BW74c, Bre77, LG74, PK72]. Constants [SL72, ZE78]. Constrained [BR76]. Constraint [FS71, Bel72]. Construct [Kyu79]. Constructed [RS75a]. Constructible [Yan71]. constructing [MS73b]. Construction [Dan73, Gau70, Smy78, Mon78b]. constructive [CMPY78, I.72]. contact [Har78]. Continued [Cus77, Hen74, Mil75b, WB79, ZP75, Ada79, Bre77, CGZ72, Dij77, Fie77, Fra79, Gau77, JT74, McC74, Sha76a, Vau78]. Continuity [JR71]. Continuous [Lyn70, Lyn71b, PS78]. Continuously [LV79]. contraction [Nev78]. Contractive [BT71]. Control [Gus70]. Converge [LS72, Sco79]. Convergence [Bro70, Bro71, CS71, Cli72, DM71, Dun75b, Dus73, Hd78, JH70, Kea78, Lam78, Lus79, Mad72, NW74, Nic77, Nic78a, SJR76, Sid79, TW74b, You70, Bey79, Bog76, BO73, But72b, Den71, DM74, Fie77, FT73, Gau77, Gus75, MT76, Nat77, Nev78, Nic78b, Sty79]. Convergent [Man71, Ral76]. convergents [Ada79]. Conversion [ZP75]. Converting [Sal76b]. Convex [MR78, Wil76a]. Coordinates [Boy73]. copy [P.71, Rhe72]. Correction [AL77, KLT78, Alf79c]. Corrections [Lon77, BDT72, FHU77]. Corrector [GH72]. Correspondence [McC78]. corresponding [Kam73]. Corrigenda [Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano74b, Ano75b, Ano75c, Ano76a, Ano76b, BW96, CS73, Ern79, Gay79, God79, JS78, LR79a, Sha75b]. Corrigendum [Ano70b, Ano71b, Ano71c, Ano72b, Ano72c, Ano72d, Ano73b, Ano74c, Ano76c, Ano77b, Ano77c, Ano77d, Lun72, Sha71a, Sha71b, Ste71c, Tuc77, Wim72a]. Cotes [Kam70]. Coulomb [CH70b, Ike75, Pex70, SG72]. Coupled [Car75a, Ehr72]. Covering [Cho71]. Coxeter [CDHW73]. Crank [For73]. Criteria [Wil76c, BLS75, GT78, Gro75, RV72]. Criterion [DM70, But72b]. critical [Bre79, LO79]. Cross [Let70]. Cross-Product [Let70]. crossings [Rio75]. Cubature [Let70, Let73, PH75, Sal77]. Cubatures [Fra71a, Hun75a]. Cubic [Ang76, EFO78, SK72, SL72, Sha74b, Wil77, WS79b, BWB76, Cio79, GK73, JH78, Ker72, Mac73, Sin72, Var77, Wil79]. cubics [Wah75]. Cumulative [AMS84]. Curiosity [Oli75]. Curvature [Bur71a]. Curve [New70a]. Curve-Fitting [New70a]. Curves [Bel74, BMP75, PP74, PP75, Tho76]. Cyclic [Coh79a, Sto78, New76]. Cyclic-Sixteen [Coh79a]. Cyclotomic [EM78, Ern79, EH79, Joh75, Leh77, LM78, McD74, New70b, Apo75, Oja77]. Cylinder [WR71].

D [Sha76a, Str71]. damping [Jel77]. Data [CS71, Kin74, MMO78, MPR77, BG75, TW74b]. Dawson [CPT70, Dij77, McC74]. Daykin [Sha76a]. Debye [ND70]. Decay [Bel74]. Deciding [BvT78]. Decimals [Dut71]. Decision [Ans76, Sac71]. Decomposition [Phi71, Ris73, Jes77]. Defective [fC77]. deferred [FHU77]. Defined [Chi73, Yan79]. Defining [SW73]. Definite [OSV79, Gla74b, Pag73]. deflation [OD78]. Degenerate [FB78, PS79b]. Degree [Erd73, Fri70, Gre70b, Hos71, MS75, PH75, Str76c].

Degrees [Mos72, McK76, Fet71a, Fet71b]. Dejon [I.72]. delay [Wie76]. Density [Wal72, WCJ72, McN73]. Department [P.71, Rhe72]. Dependent [GT76]. Derivative [Gre78, Gup78a, Hud71, Coh73, Ker72]. Derivative-Free [Gre78]. Derivatives [Gre72, Ike75, Kow71, Lau73, Lyn70, Ste71b, Bus74, Jel77, Naz76, Tho77, Bre75c]. Derived [Gol70]. Descent [GP71]. Descriptions [Ano71e, Ano71f, Ano72e, Ano75d, Ano75e, Ano76d]. Design [Ber75, Lew75]. Details [BH78a]. detecting [Bur75]. Determinants [LM70, GJ74, Lun72]. Determination [Bur71a, Sta73b, WJ76a]. Determining [JH78, BWB76, How78, PS77a]. Diagonal [DM71, AO74, BR72]. Diagonalization [WBH75]. Diaz [Sha76b, Sha76b]. Dickson [CS73, Lon77, te 73]. Difference [Bon71, BT71, Cha73, CS71, Cim71, Gek74, KP78, Kow71, Kre72a, Kut71, Liv75, McA71, Mir71, Red76, RP70, ST70, Ste71b, Var70b, Var72, Wim75, van74, AG73, Bey79, Bra73b, Bur78, CH78, CS75b, FT73, Gol73, GT78, GKS72, Gus75, Hed75, Hof79, KT78, LP74, Mad72, Nev78, Nic78b, Oli74, Oli76, Osh72, Skö79, TW74b, Tun75, WF74, Wei75b]. Differences [Cry71, DO74, Gup78b, Kro79]. different [Moc76]. Differential [Bru70b, Cha73, Fat78, For73, GH72, Hil73, KLT78, KD71, Kow71, Mic73, Mir71, Sha74a, SF74, Wim75, Alf79b, BO78, Bro79, Bru74b, Bur78, But73, CG73, EF79, FO77, FT73, Gir76, Gol77b, GP79, GW75, Gup79, GK73, Hed75, Hul72a, Jai72, Kre72a, MH77, Mos77, Nos73, PR74, Ral76, SD77, SL74, Sha73a, Sha78, Sni76b, SW79c, Wie76]. Differential-Difference [Cha73, Wim75]. Differentiation [GP70, Kro70, SW79d, Win70, Kah72a, PK72, Sal74]. difficulty [Dun75c]. Diffusion [Tod70]. Digit [Wil78]. digital [Lew75]. Digits [BMN70b, Sha71b]. Dilogarithm [Mor79]. Dimension [HS75a]. Dimensional [Boy73, Hay73, PP77a, PP77b, RV73, RP70, AG79a, Bar76, Hun74a, Lon77, Lyn76, ML79, OW79, Sik79, de 79]. Dimensions [GR70, Per70, Eil72]. Diophantine [Sha71a, Alt71, Bla76b, Bra71, Bru70a, Cus71, Cus72, Gri78]. Direct [LF72, Phi73]. Direction [BR76]. directions [Naz76]. Dirichlet [Bak73, BDT72, DD75b, Hud73, McA71, PPW77, Spi73, Ste71b, Wei75a]. discontinuities [Har78]. Discontinuous [CS71, RP70, BO78]. Discrete [AG79b, BM70, FB78, Hul72a, MS73a, Sam77, Usm78, Wil70, Win78, Wol75, Abd75, Gut77, Joh74a, Mad72, Sim72]. Discretization [Bru77]. discretizations [Zla75]. Disk [Boy70]. dispersion [CH78]. Displacements [FB78]. Dissipative [GT76]. Distinct [BE75, Cho71, Klo75, DMP74, Kis77, Kis78]. Distribution [AB72, BLM70, Bre74, Die71, Leh70, Nie72, SL72, AD73, Bre75b, Fet79, HB75b, Hen75, KMR77, Nie74, Nie76, NO76, Rio75]. distributions [For72]. Divergent [Boa77]. divided [Kro79]. Divisibility [Joh78, McD71]. Divisible [SS73, Kis77, Par77]. Division [Ste71d, Lon77]. divisor [HM73, HM75]. Divisors [Joh74b, McD74, WCJ72]. do [Fal76b]. Domains [Bab72, Bur70, Bur71a, SW78, SW79a, BDT72, Man78]. dominant [AL77, Alf79c]. Dominating [Ise78]. Dorothy [BW96]. Double [Chi73, GN78, Bay78]. Doubly [Bur71a]. Dover [I.75]. Drazin [SPB75]. Drives [GS75]. Dyadotropic [Coh76b, Coh79b]. Dynamic [Tur73, WBH75].

each [P.71]. Early [Lyn70]. ed [S.75]. edited [Str71]. edition [Mac72b]. effect [DD75a]. effective [Cul79, PZ77]. Effects [Wid71]. Efficiency [Knu75, Moe77]. Efficient [Kro70, PS73, Abd75, Var79]. Eigenfrequencies [TT73]. Eigenfunction [SJR76]. eigenfunctions [Rak75]. Eigenproblem [Kut71, Gol74a]. Eigensystem [Var70a]. Eigenvalue [Gek74, Kol79, OSV79, Red76, Ruh74, Tro73, BO78, BO73, Kre72a, Sim72, Ste75]. Eigenvalues [Lin70, Rei71, Ruh79, Rak75]. Eigenvectors [Gek74, Lin70]. eight [Hun74a]. eight-dimensional [Hun74a]. elasto [BJM77]. elasto-plastic [BJM77]. electromagnetic [PS77a]. Element [Bab72, Bab73, BR79, BZ70, Fal76b, Nic77, SW78, SW79a, Sco76, Zla74, Bak75, Bak77, Bla76a, BS77, BJM77, DD73, Fal76a, FHP78, HSK74, JT75, Kol79, LR79b, LG79, Lus79, Man78, RT77, SW77, Tho77, Zla75, Zla78]. elementary [McC77]. Elements [BZ70, Fia71, HS75a, Ker70, Klo75, Ste74, Bak77]. Elimination [Ste74, Duf74]. Ellipsoid [Kel79]. Elliptic [BMP75, BS71, Fet72, FGH^{+}74, GW71, Gla71, Luk70, McA71, PP74, PP75, Per70, Sch70c, Ste71b, TT73, Wri73, ZC70, Bak77, Dum79, Jes77, Joh74a, Nic78b, RT77, Sch76a, Sch77]. Elliptical [Tro73]. end [Fet74]. Endpoint [SK72]. Energy [GP79]. Energy-conserving [GP79]. engineering [Usm78]. Engineers [S.75]. Englewood [Abr74, Bre75c]. English [Mac72b]. Entropy [Le 77]. Envelopes [Mv78]. epsilon [BR74, Gek72]. Epstein [Sha75b, Sha75a]. Equal [Bru76, DSW79, RB73, Sal71a]. equality [Ans78]. Equally [Gre70b]. Equation [Alt71, Bla76b, Bru70a, Car74, Cha73, Cha71, Dus73, For73, IN79, LG72, Mad73, Mic73, Ric71, Sha74a, SS70, Tod70, Var70b, van74, Bar76, BR72, Bre75a, Bru74a, CL75, CL78, Día79, EF79, Gut77, Net73, Nic78b, OW79, Phi73, PW76, SD77, Sha78, Sha71a, Sko75, Toi77, Wei72b, WZ78]. Equations [BP70, Bra71, Bru70b, Car70, DM70, DDW78, EP75, Ehr79, Fal76b, Fat78, Fra71b, Gol72, GH72, HS75b, Hil73, KD71, Kow71, Le 77, Lin70, Lir71, Liv75, McA71, Mir71, Nic77, Pai73, Rhe76, Sch70b, ST70, SF74, Swa79, Tur75, Var72, Zla74, ZA71, Abd77, Alf79b, BO78, Bak75, Bak77, Bow76, BR74, Bro75, Bru74b, Bru77, Bur78, But73, CG73, CHK79, Den71, Dou79, DD73, Ehr72, Fal76a, FO77, FT73, Gal78, Gay77, Gay79, Gek72, Gir76, Gla79a, Gol77b, Gri78, Gro75, GP79, GW75, Gup76, Gup79, GK73, Hd78, HKK75, Hed75, HP78, Hul72a, Jai72, Jes77, JP74, Kre72a, Mad72, MH77, Mos77, NW74, Nos73, Oli74, Oli76, PS77b, PS77a, PS78, PS79b, PR74, Rak78, Ral76, RT77, Sag70, SL74, Sha73a, Sik79, Sin72]. equations [Slo76, Sni76b, SW79c, Tho79, Ven73, Wie76, Rhe72]. Equi [Hos71, Mac73]. Equi-Spaced [Hos71, Mac73]. equivalence [Hud74]. Equivalent [Ske79]. Erdélyi [Eil71]. Erdos [JR70, Mak70]. Errata [Ano71g, Ano71h, Ano72f, Ano75f, Ano75g, Ano76e, Wil79, Bre75c, Eil71, Fet71a, Fet71b, Hil76, Mac72b, Sha76a, Str71]. Erratum [Wre73]. Error [Bas70, BW74c, Car75b, CR78, DDW75, Fal74, FM71, FCC73b, Het72, Het73, JT75, Kam70, Kea78, Let70, Let73, Mad73, MR71, Mie77, New73, New74, NRRZ71, Pai73, Phi70, Phi73, RR70c, Ram71, SD77, Sch70d, SW71, Skö79, Sni76a, Ste71d, Bak73, BN70, BK74, BF77, BEJ76, Cia78, FCC73a, Fet74, GK73, HR72, Kah72a, Lyn76, McC74, PK72, Rak75, Rak78, Sch78d, Sim72]. Errors [Gus70, RR70c]. Estimate [Bon71, BK74, BF77, McC74, Rak75]. Estimates [Bas70, BNS75, DDW75, Fie77, Joh74a, Mie77, NS74c, SW78, SW79a, Sco76, Bak73, BO73, Cia78, Fal74, JT75, Rak78, SD77, SW77, Sim72, Tho79]. Estimating [Knu75, OSV79, RR70c]. Estimation [Gus70]. Estimators [SW71]. Euclid [Oja77]. Euclidean [Cio79]. Euler [BW74c, Bre77, Dus73, Hd78, LP73, NS74a, Wal72, de 79]. evaluate [McC77]. evaluating [JT74, WB76]. Evaluation [Cho73, CS73, GW71, Gus70, KR79, LG74, LL71, LL72, Lin78, New73, New74, New75a, Sch78b, SG72, BC79, DC79, GJ74, HR72, Kad76, ML79]. event [Bur75]. every [CS75b]. Evidence [RP71]. Exact [Die71, Bey79, How73a, How73b, SW79c]. Example [Dor71]. Existence [Bar76, DM70, Fri70, Gut77, Sal71a, Sim72, TW74a, Gay77, Gay79, Gro75]. Exit [RV72]. Expansion [Kim72, Lyn71a, LP73, Won73, Dij77, Fet79, Fie72, GJ74, Lyn76, McC74, PWKW79, Sch76a, Sch77]. Expansions [BMN70a, Des74, Ell71, FI75, Hos73, SJR76, Wim62, WR71, dW73, Fie77, Mac72a, Sch78d, SP75, Tem75, TE72, Wim72a, de 79]. Experiment [Lyn79b]. experiments [KS78]. Explicit [Fet72, LM70, MM75, Oli75, Sch70a, Wil76c, Alf79c, NS74b, Lun72]. Explicit-Implicit [MM75]. Exponential [Bel74, EP75, Pru79, Ros71a, SL74, BC74, Chi72, Kam73, Wag79b]. exponentials [Bra73a]. expression [RS78]. Expressions [LM70, Lun72]. Extended [Mon76, Mon78a, Mon78b]. extension [Cas78, FHU77]. Extensions [AD73, Coh76a, Hay73, Mon79]. Exterior [GN78]. Extrapolated [Lar73]. extrapolation [Gol77a, KS78, Sha73a]. Extremal [RC78, Pag73]. extreme [Bue77].

F [Eil71]. F. [Sha76b]. Factor

[New70b, Kam73, Pom75]. Factoring

[Ber70, Leh74, Ris73, WR75, Wan76, Moe77, MB75, Wan78]. Factorisation [Mil75a].

Factorization [BW74a, BW76, DGKS76, LL74, Zas78, BH74, MB75]. Factorizations [AG74, FP74, GGMS74, BLS75]. Factorized [Gol76a]. Factors [EES74, EGRS75, GMS75, HB75a, Leh77, LM78, Mig74, Shi78, Kis78, WJ76a]. Familiar [DLT73]. Families [Dia78b]. Family [Gol70, Ise77]. far [Cat77]. Fast [Ban78, GB79, BR72, BH74, PPW77, KL79, Pol71, Ram71]. faster [Sty79]. FCT [IN79]. Feit [Ste71a]. Fekete [BPW75]. Fermat [HB75a, Joh77, Joh78, Shi78]. Fewer [McK79]. Field [Pol71, BWB76, JH78, Oja77, PZ77, WB76]. Fields [Ang76, Ber70, BMP75, Bue76, BWW77, Coh79a, Dia78b, DSW79, EFO78, Ger75, NS74a, New70b, Par78, SS73, Sha74b, Sta75, Wil77, WS79b, Bue77, Cio79, Hen75, Lak74, Lak75, Par77, PS77c, Sha76b, SR76, Wan76, Wil79]. filter [Lew75]. Filtering [Sha75c]. find [LG79]. Finding [Atk75, Cha71, CMPY78, CH70a, Ste69, Ste71c, Abe73, Cox72, SR76]. Finite [Bab72, Bab73, BR79, Bak77, Ber70, BMP75, BZ70, Bus74, Cry71, DC79, Fal76b, Gek74, Gup78b, HS75a, KP78, Kow71, Kut71, Kyu79, LR79b, Nic77, PP77a, PP77b, Pol71, Red76, RV73, Ril71, Ros71a, SW73, SW78, SW79a, Sco76, Var72, Zla74, Zla75, van74, AG73, Bak75, Bey79, Bla76a, Bra73a, BS77, Bra73b, BJM77, DD73, DF77, Fal76a, FHP78, Gol73, HSK74, Hof79, JT75, Kol79, LP74, LG79, Lus79, PS77c, RT77, SW77, Skö79, Tho77, Yan79, Zla78]. Finite-Difference [Gek74, Kow71, Kut71, Red76, Var72, van74, Bra73b, Gol73]. Finite-Dimensional [RV73]. Finite-Element [BR79]. First [DM70, Hud71, New70b, ST70, Bak75, BH78a, Bow76, Bre73, Bru77, CG73, Coh73, Gla79a, Gro75, HKK75, Joh77, Ker72, LR79b, NW74]. First-Order [ST70, LR79b]. Fischer [Hun74b]. Fitting [Bel74, Ise77, New70a, SL74]. Five [PP77a, BR72, Kis78, Par77]. Fixed [Klo75, Kut71, Dun76, KY72]. flow [Car75a]. Flows [RP70, PWKW79, WF74]. Fluctuations [BH78b]. Fluid [Tur73, Mar75]. Force [Fet71a, Fet71b]. Form [Bai79, BH78b, Gau79a, GW71, JS78, MW77, Sha71c, WS79a, ZA71, How73a, How73b]. formedness [Bau77]. Forms [Cus71, LL74, Lyn71a, Ske79, WZ72, Cus72, Sch74]. Formula [Cha71, Fat78, Fet72, Gaf78, Lyn70, Lyn71b, Lyn71c, RG70, Spi71, Toi78, Bor72a, FNC75, Nos73, Sch77]. Formulae [Bar78, Gus70, Sal71a, SS74, Sil70, Bau77, Gup76]. Formulas [ABH73, Dan73, Dur71, Fri70, Gla74a, Kah70, Kam70, KP78, Kum74, KJ74, LS72, PB74b, Sal71b, Sal77, SK72, Ske79, Wil70, Bar76, Bar77, Hil75, Hil76, Kah72a, New76, PK72, PB74a, PH75, Sal76b, Str71, Str76c]. Forsythe [AD73]. found [Sha76b]. Four [Fia71, GT76, RB73, SS73, Shi78, te 74]. Fourier [KL79, Pol71, Ram71, FGH^{+}74, Lyn70, Lyn71a, Lyn71b, Lyn71c, MMO78, New73, Osl72b, PS77c, Sch78a, Shr74, Sko75, Win78]. Fourth [Kut71, Oli74, ZA71, Mac72b, Man78, Ste76, Str76c, WF74]. Fourth-Order [Kut71]. Fraction [Cus77, Hen74, WB79, Bre77, CGZ72, Dij77, Fie77, Fra79, Gau77, McC74, Vau78]. Fractional [DO74, BPD79]. Fractions [McC78, Mil75b, ZP75, Ada79, JT74, Sha76a]. Fraser [Dun75b]. Fredholm [HP78]. Free [Alo76, Cry71, Gre78, HS75a, Sac71, Sto78, Wag72, Bau77, New76, PWKW79]. Frequency [Pau72, PS77a]. Friedrich [Vai70]. Frobenius [How73a, How73b, RV73]. Frog [For73, AG79a]. full [ZE78]. Function [Bur70, CS75a, Car75b, CHT71, CST73, Coh76a, Eck76, EES74, Ell71, FCC73b, Gup78b, Hud71, Kim72, Kol70, Kol72b, LF71, Lal74, Leh70, Lin72, MR71, MSW71, Mor73, Mor79, Sha70, Spi71, Wal72, WCJ72, BC79, BPD79, BEJ76, BC74, Bre79, Coh73, Cox72, DC79, Dun76, FNC75, FCC73a, Fet74, Fie72, FK75, GS78, HR72, LG79, Lyn76, Par72, Tem75, Wre73]. Functional [Fra71b, Ise77, Lyn76, PWKW79]. Functionals [Gus79]. Functions [Amo73, BK78, Bus76, CH70a, Des74, DLT73, Fal76b, Fet72, Ike75, Jer77, JR71, Lar70, Lau73, Let73, Lir71, Luk71a, Luk71b, Luk72, Lyn70, Lyn71b, Lyn71c, MR71, MS73a, Mic73, Nas74, ND70, Osl75, Pex70, Phi79, RR70c, RS75b, Sam74, Sch70c, Sha75a, Smy78, Spi76, SG72, Tha79, Wag75, Wei75a, WJ76b, Wim62, WR71, Amo74, BDM78, Bel72, Bla74, BEJ78a, BEJ78b, Bue77, Bus74, CH78, Chi72, Coh77, Dum79, Fie77, Gab79, Gau77, Gla74b, HB75b, Kol72a, LO79, LK73, McN73, Pag73, PS77b, PS78, PS79b, Sch76a, Sch77, Sch76b, Sch78d, SP75, Str71, Tem75, TE72, Wim72a, Eil71, Sha75b, Wim72a]. Fundamental [I.72, JH78, PZ77]. Further [Lin78, Luk70, McC78].

G [Eil71]. Galerkin [Ste79, BBT77, Bow76, BNS75, CHK79, Día79, Dou79, DD75a, DD75b, DDW75, DDW78, Gun77, HS75b, Hul72a, Hul72b, Mos77, Nat77, NS74c, Sch74, Sch70c, Tho79, Wah75]. Galois [Sta73b]. Gamma [Kol70, Kol72b, Spi71, Wre73, Tem75]. GAP [Mar75]. Gaps [Bre74, Bre73]. gas [WBH75]. gas-dynamic [WBH75]. Gauss [JS78, PB74b, Bar76, Bar77, DC79, EP79, Kam70, Lar73, Mon78a, PB74a, Sha71c]. Gaussian [Bar78, CR72, Dan73, Duf74, Gau70, JR70, KN70, LV79, LTK75, Mon76, Mon78b, Ste74, Wun73]. General [Dur71, Kre72b, RV73, Swa79, Var70a, GJ74, OW79, PPW77, Ste76]. Generalization [For77, And75, Dij77, Vau78]. Generalizations [Hud76, Par74]. Generalized [BMN70a, Bra73b, Dur71, LTK75, Man71, McK72, New71, Nin70b, Osl75, Pai79, Phi70, Poo75, WJ76b, BC74, Bru74a, CH78, HB75b, Ste75, Wei72b]. Generated [Fia71, GR79, How77, Nie72, SS78, Bro79, Nie74, Nie76]. Generating [BDS74, Boy73, Bru70a, CS75a, MY79, Coh77, Mon79, Sha71a]. Generation [KP78, Naz76]. Geometric [Sal76a, PZ77, Tun75]. Geometries [BCH73]. geometry [Mad72]. Gershgorin [Ste75]. given [JH78, Wag79a]. Global [Bru74a, Kra78, dS77, FHP78, MT76, RV75, de 76]. Goldbach [Zwi79]. Good [Zar76]. grade [ZE78]. Gradshteyn [Mac72b]. Gram [BG70, DGKS76]. Graph [Pom79]. Graphs [BDS74, Fit74, McK79, Str71]. Greatest [Wag79a]. Greedy [GR79]. Green [MS73a]. Group [Atk75, Hun74b, New75b, Str75, Hun74a, Lak74, Lak75, McK76]. Groups [Alo76, Ans76, BNZ73, Bue76, Dix70, FM78, Ger75, MY79, Ril71, SW73, Sta73b, Sto78, Ans78, Cat77, Esp75, New76, Sha76b]. Growth [Boa77]. Grunsky [Ros71b]. Gutknecht [S.78].

Hadamard [Yan71, Yan79]. Half [Lir71, Gau94]. half-century [Gau94]. Hall [Abr74, Bre75c]. Hammerstein [HS75b]. Hamming [S.75]. Hand [Der77]. Handbook [Str71]. Hankel [CH70a, Fra79, Mac72a, Phi71, Ris73]. Hans [S.78]. Hans-Rudolf [S.78]. Hardy [Win70]. Harmonic [NGF74, Ehr72]. Harmonics [Fet70, Fet71a, Fet71b]. Hart [Dun75b]. Having [Lyn71c, Gay77, Gay79, LLS70, Wag79a]. Heat [Var70b, Bar76, Car75a]. Heinz [S.78]. Helmholtz [OW79, PW76]. Henrici [S.78, I.72, Tha77]. Hensel [Zas78]. Hermite [BG70, Bra71, Coh76a, Coh77, ET73, Kre72b, Wah75]. Hermitian [Gol74a, Gol74b, Rak75, Rei71]. Hessenberg [Lar73, SPB75]. High [PPW77, SG72, Tho77, van74, Gab79, Gup76]. High-Order [van74, Gup76]. Higher [AG73, Bla76a, BS77, CL75, GR70, Jer77, Jel77, Eil71]. Highly [Per70, MW76, Nat77]. Hilbert [Dot78, Lar70, Nev78]. Hill [Eil71, S.75, Jer77]. Homomorphisms [Ril71]. homotopy [CMPY78]. Householder [I.75, P.71, Rhe72]. Howland [LL71, Lin78]. Hsieh [Cia78]. Huddleston [Mer72]. Hurwitz [Spi76]. Hybrid [Oli76, Gup79, Har78, RT77]. hydrocode [Hic78]. Hydrodynamic [RP70]. Hyman [War76]. Hyperbolic [BT71, DDW78, Liv75, MM75, ST70, ZA71, AG73, Bak75, Bur78, CHK79, Dou79, Gol73, GA74, GT78, Gun77, GP79, Hou77, JP72, Oli74, Oli76, Skö79, Wah75]. Hypergeometric [Bus76, Kim72, Osl75, Wim75, DC79]. Hyperosculatory [Sal71b, Sal77]. hyperrectangles [Hun75a].

ibid [Sha76a]. ibn [Bor72a]. ideal [Hen75, Lak74, Lak75]. Identifying [Mos77]. Identity [Gla71, Osl75, Wri73]. Ignoring [Mil71]. II [Eil71, Fet71b, P.71, BMP75, Coh79b, Cus72, GKS72, HM75, How73b, Lak75, Luk71b, Lyn71b, Nie74, PP77b, Sch76b, Sha78]. III [BN70, BNZ73, Har78, Luk72, Lyn71c, Nie76]. Ill [Dor71, Fra74, Rhe76]. Ill-Conditioning [Dor71, Rhe76]. Ill-Posed [Fra74]. Imaginary [BWW77, Dia78b]. imbedding [Als74]. Implementation [CDHW73, Gal71, Ruh79, Var79]. Implementing [Gup78a]. Implicit [Fat78, MM75, Wat78, Bru74a, CL75, CL78, Var79, WF74, Wei74b]. important [DC79]. Improper [Hab75]. Improved [HKK72, Kin71, Mac73, Nin70a, Rak78, Alf79a, Kah72a, Wan78]. Improvement [Slo76]. Incomplete [Kol70, Kol72b, Kol72a, Tem75]. Indefinite [Nic78a, Sch74]. independent [GT78]. Index [AMS84, P.71, Pol78, Rhe72]. Indexing [Fit74]. Inductive [Kyu79]. Inequalities [Bus76, Ker70, Nas74, BF77, Fal74]. Inequality [Mig74]. Inertia [BK77]. Infinite [Bab72, Cry71, Gau75, KJ74, Lir71, Rob71, Als74, BC79, DR72, Kam73]. Infinitely [Bru70a, Sha71a]. infinity [HM77, Jel77]. Information [P.71]. Initial [CS71, Gou70, Hen77a, Hul72b, Kra78, MMO78, SS70, Var70b, AG79a, AL77, Alf79b, Alf79a, GT78, Gun77, GKS72, Gus75, Oli74, Oli76, Osh72, Skö79, dW77]. initial-boundary [GT78, Gun77, Skö79]. Initial-Value [SS70]. inner [Esp75]. Instability [For73]. Integer [Der77, LLS70, ND70, Sha73b, WB79, Ada79, FNC75]. Integers [BMN70b, Cho71, GR79, Leh74, Wag72, WR75, BH78a, Kis78, Sha71b]. Integral [Blu79, BNZ73, Car75b, Cha74, CPT70, Dre78, GW71, Gla71, Gla79b, Kin71, Lin70, Osl72a, Osl72b, Sch78b, Wal78, Wri73, BC74, Bru74a, Bru77, Chi72, Dij77, Don79, Gau79b, Gla79a, Gol77b, Gro75, HKK75, HP78, McC74, ML79, NW74, Net73, Phi73, Rak78, Sag70, Sch78a]. Integrals [Cho73, FM71, Hab75, KJ74, LL71, Lin78, Lin72, Luk70, Sam74, ZC70, BC79, BPD79, DR72, EP79, Gab79, Gla74b, Kad76, LK73, Mac72a, Mac72b, CS73]. integrand [Lyn76]. Integrating [Mir71]. Integration [ABH73, Bru70b, But72a, Fat78, Fri70, Hab75, Het73, Ise77, NRRZ71, Oka76, RR70a, RR70b, Win70, dW73, BN70, But72b, CR72, DR72, Let77, Nos73, Str76c, Wei72b, Zla78]. integro [Gol77b, GK73]. integro-differential [Gol77b, GK73]. Interesting [Wil78]. Interior [Bon71, BNS75, NS74c, SW77, Tho79]. internal [GA74]. Interpolating [Dun75a, Shr74, DD75a, Dun75c, MS73b, TW74a, Wil72]. Interpolation [Bur71b, Cia78, Gaf78, GW78, Gre70b, GR70, Gus70, Hos71, Hos73, Kre72b, Kro70, MR78, McK72, New70a, New71, Sal71b, Sch70d, Tho76, Win70, Wuy75, BG75, Bar76, Bar77, Bru74a, Sal76b].

Interpolations [MPR77]. Interpolatory [Sal73, SK72]. interpretation [Dum79]. Interprimed [Sha73b]. Interval [Jel78, Lyn71b, Rob71, Als74, Kam73]. Intervals [Cad71, Gau75, Boy78]. Introduction [Abr74]. invariance [Lyn79a]. Invariant [Var70a, Als74, Joh73]. Invariants [EM78, Joh75, SS73, Ern79]. Inverse [Alf79b, fC77, Fet72, Hal78, Ker70, Sch70a, BEJ76, Fet74, Fra79, SPB75]. Inverses [Dia78a, Dur71]. Inversion [Der72, Lyn70, Lyn71b, Lyn71c, Tre74, BH74, FHP78]. Inverting [Gol72]. Investigation [Pex70, CC72]. investigations [Ser75]. involution [RS78]. Involving [Ple72, Sam74, WR71, Gab79, LK73, PS77b, PS79b]. Irene [Str71]. Irrational [BMN70a]. Irreducible [Dix70, Mos72, PP77a, PP77b, Esp75, McK76]. Irregular [EM78, Joh74b, Joh75, Sco76, SG72, Wag78, Ern79]. Irregularities [Bre75b]. Isoclinal [Der72]. isoparametric [FHP78]. Israel [Bog76]. Iterated [Amo73, FHU77]. Iterates [Lal74]. Iterating [Smy78]. Iteration [Abe73, fC77, Dia78a, DM70, Kin71, Lar73, Nin70a, Par74, Bog76, How78, Ral76, Sag70, Slo76]. Iterations [Ric71]. Iterative [Mv77, Ple72, BD76, Dot78, Gol73, Nic78b]. Iwasawa [Joh73]. ix [S.75].

J [Lun72, Mal71, Sha71b, P.71]. Jacobi

[Ell71, Sam71]. Jacobi-like [Sam71]. Jacobian [Gay79, Dum79, Fet72, Gay77, Sch76a, Sch77, Sch70b, SW79c]. James [Fet71a, Fet71b, Gau79b]. January [JS78]. Jenkins [For77]. joining [Rio75]. Jonquiere [FK75]. Journal [Cry71].

Kelvin [Phi79]. Kernel [Lar70, Rak75]. kernels [Bar77]. Kind [BW74b, DM70, ZC70, Bow76, Bru77, Gla79a, Gla74b, Gro75, HKK75, HP78, NW74, Net73, Sag70]. Knot [Ril71, CR78]. Knots [Hos71, BCSW78, Ker72, Mac73]. known [Cat77, SPB75]. Kreiss [Gol77a]. Kummer [Gau77]. Kurrah [Bor72a]. Kutta [CS79, Oli75, SW71, Var79, Wei74b]. KWIC [P.71, Rhe72].

L [Gau79b, Lun72, WB76]. L. [Mal71]. Lab [Fet71a, Fet71b]. Labelled [Ber75]. Laboratory [P.71, Rhe72]. Lagrange [Fal76a, Kol79]. Lagrangian [Sal74]. Laguerre [Coh77]. Lanczos [Lyn74, PS79a, Ruh79, Sco79]. Laplace [PPW77, Won75]. Large [Ber70, Cad71, CP79, Leh74, PP74, Ruh74, Ruh79, Won73, te 74, Bre73]. Largest [OSV79, HM73, HM75, McK76, Pom75]. last [Joh77]. Lattice [BRW71, Die75, BW96]. Lattices [Kyu79, Zar76]. Lauchli [S.78]. Laurent [MS73b]. Lavoie [Lun72]. Law [ZA71, AG73, Hic78]. laws [Hof79]. Lawson [Cli72, EW76]. Lax [Eil72, Vai70]. Layer [GN78]. Leap [For73, AG79a]. Leap-Frog [For73, AG79a]. Least [BS71, BT72, Car74, Kin74, Pai79, PP75, Sam77, Wil70, Bak73, Bog76, Car75a, Jes77, Kis77, NW74, Ser75, Wag79a]. Least-Squares [Kin74, Ser75]. Legendre [Gau79b, Blu79, Kam70, Mon78a, Sam74]. Lehmer [Ste71c, Ste69, WJ76b]. Leibniz [Osl72a]. Level [Bra77]. like [Sam71]. Limiting [Sha74a, Sha78]. Line [Ore73, Wun73, Boy78]. Linear [AG79b, BDM78, BG73, Bra71, Bru72, BK77, Cus71, EW76, Fat78, Fra71b, Gol72, Gus79, Lam70, Le 77, Mv77, Mil75b, Nie72, Pai79, RV73, Rob71, Sha75c, Ste71b, dS77, AG73, Abd75, Abd77, Alf79b, Alf79c, Als74, And75, BK74, Bog74, Cas78, Cus72, NW74, Nie74, Nie76, Phi73, Pru73, RV75, SW79b, Toi77, Wat73, Wil76a, dW77]. Linearly [BR76]. lines [Hou77]. Liouville [Gek74, Hal78, Red76]. Listing [Wei74a]. Littlewood [BH78b]. Lobatto [PB74b, PB74a]. Local [GP71, Sha73a, BS77, Bra73b, Tho77, Var77]. Logarithmic [Phi70]. logarithms [Car72]. Long [GR79]. Longest [RS75a]. Low [Car75b, JR71, Oli75]. Low-Order [Car75b, JR71, Oli75]. Lower [Boy70, Hag70, Hag73, RS75a, Tuc73, Tun75]. Lucas [Mor75]. Lundquist [JP74].

M [Mac72b]. Maclaurin [LP73, de 79]. magnetohydrodynamics [JP74]. Magnus [Eil71]. Make [Sco79]. Manifolds [CSW77]. manipulation [PS73]. Many [Bru70a, Sha71a]. Mapping [CA79, HKK72, HKK75]. mappings [Gut77]. Maps [Bur70, CMPY78]. Marching [Ehr79]. Marginal [Bru70b]. Mark [Lyn79b]. Markoff [Ros76]. Markov [RP71]. Martin [S.78]. Math [BW96, BEJ78a, Ern79, Gay79, God79, Hil76, LR79a, Lun72, Sha71a, Sha71b, Sha75b, Sha76a, Wil79, Wim72a]. mathematical [Str71]. Mathematics [AMS84, Gau94, Gau94]. Mathematik [S.78]. Matrices [Bra71, BGW70, fC77, Der72, Gol72, HP72, Lar73, LM70, Man71, P.71, Par74, Phi71, Ple72, Poo75, Ris73, Ruh74, Ruh79, Tre74, Tur73, Yan71, Yan76, BR72, Don79, Duf74, Gol74a, RH72, RS78, SPB75, WBH75, Yan79, Lun72]. Matricial [Mir71, SL74]. Matrix [Cha74, GGMS74, GMS75, Hou72, Ker70, Mv77, OSV79, Pai73, Rei71, Ric71, Ris73, Sch70a, Toi78, Var70a, BH74, Fra79, Gay77, Gay79, Gol74b, HM77, How73a, How73b, OW79, PW76, Toi77]. Maximal [FM78, PP77a, PP77b, Yan76]. Maximum [BNS75, SW78, SW79a, Var70b, Bra73b, SW77]. Maximum-Norm [BNS75]. Maxwell [Hic78]. McGraw [Eil71, S.75]. McGraw-Hill [Eil71, S.75]. Mean [Sal76a, BCSW78, Tun75]. Means [Cry71, Gol70, Bur75]. Measures [Rhe76]. mechanism [Bau77]. Mehler [Coh76a]. Meijer [Fie72]. Membrane [Kut71, TT73]. Membranes [Tro73]. Mesh [BKO73, Cim71, Ven73, de 76]. Meshes [BR79, Sco76, Nat77]. Method [Bab72, Bab73, Bas70, BZ70, CM78, Cry71, Dus73, Fal76b, Fet70, For77, Fra74, FB78, GW78, Gre72, Gre78, GH72, Had78, Hal78, HKK72, HS75b, Kea78, Kin72, Kow71, Lam78, LL71, Lin72, MMO78, Mv77, Mv78, NGF74, Nie72, Pai73, Phi70, Sam77, SW78, SW79a, Sch70b, Sch70c, Sco76, Ste69, Ste71c, Tur75, War76, Wid71, Zas78, Abd75, AI75, AD73, Alf79a, And75, Bak73, Bak75, BF77, BB74, Bla76a, BS77, BJM77, Bro75, Car75a, Car73, Cas78, CG73, CHK79, CL78, Cox72, Den71, Día79, DD73, DD75b, Eil72, Fal76a, For72, Gol73, Gri78, Har78, Hic78, HSK74, Hou77, HP78, Jes77, KY72, KS78, Kol79, LP74, LG79, Lus79, MW76, Mon79, MT76, MB75, Mos77, MS73b, Nat77, Nic78b, Nie74, Nie76, PZ77]. method [PW76, Ral76, Sik79, Sko75, Ste79, Sty79, Tho77, WF74, Zla78]. Methods [BR76, BG73, BS71, BT72, BNS75, Bro70, Bru72, BK77, But72a, CS71, CS79, DM71, DDW78, FGH^{+}74, GGMS74, GMS75, Gol70, Gol72, GT76, Gou70, Gre70a, Gup78a, Hen77a, Hul72b, Jel76, Jel78, Kin74, Lam70, NS74c, Oli75, Pol78, Ruh74, S.75, ST70, SW71, Sha70, SK70, Wat78, Wd74, You70, Zla74, dS77, van74, Abe73, Alf79b, Alf79c, Bak77, BD76, BDT72, But72b, But73, CMPY78, CR78, DM74, Dou79, Gab79, Gal78, Gla79a, Gol76a, Gol77b, Gun77, GW75, GW79, Gup79, Hd78, Hud74, Hul72a, Jai72, Jel77, Kel75, KMR77, LR79b, McC77, Moc76, NS74b, Nic79, OW79, Oli74, Oli76, Phi73, PR74, RT77, RV75, Sag70, SL74, Sch74, SW77, Ser75, Usm78, Var74, Var79, Wei74b, Wie76, dW77]. Metric [Gol70, GW78, Gre70a, Ore73, Gol76a]. Metropolis [Sha71b]. Metsänkylä [Ern79]. Michaud [Lun72]. Microfiche [JS78, P.71, Rhe72]. Mildly [Car70, Lam70]. Milton [Str71]. Miniaturized [Luk71a, Luk71b, Luk72]. Minima [GP71]. Minimal [Fri70, JR71, MY79, PK72, RR70a, SW73, JT75]. Minimax [FS71, SW79b, Coh73]. Minimization [Bas72, BR76, Lyn79b, Ore73, Sha70, Bre75c, Lyn79a, Naz76]. Minimum [CH74, Dur71, Kah72a, RR70b, Rum79, Sch78c]. Minkowski [Don79]. minors [GJ74]. Mixed [Kow71, Ste71b, Var70b, ACR79, BJM77, GKS72, Gus75, Osh72, Ste76]. Möbius [Lyn70, Lyn71b, Lyn71c]. model [Nic78b]. Models [Hed75]. modes [Lus79]. Modification [FP74, Gol72, Sch70b]. Modified [Bru72, Gau70, Nas74, Vai70, Amo74, Bla74, Car73, GS78, KY72, Kro79, LG79, Nos73, Par72, Wah75]. Modifying [Gay77, Gay79, GGMS74, GMS75, Ste74]. Modular [Coh79a, New75b, How73a, How73b]. Moduli [Cho71]. modulo [DMP74]. Modulus [Bur71a, Wag79a]. molecular [PWKW79]. Mollifier [Hil73]. Moments [Gau70, LL72]. Monotone [Lew72, Man71, Wim72b]. monotonic [Dot78]. Monotonicity [Ple72, Str72]. Monte [Oka76, PWKW79, Pol78]. Mordell [Bre75a, Mal71]. Movable [GS70]. moving [GA74]. MR [BW96, Wil79]. Multi [Bra77, Hay73]. Multi-Dimensional [Hay73]. Multi-Level [Bra77]. Multiderivative [Jel76]. Multidimensional [BG70, Cus77, GP70, PS73].

Multigrid [Nic78a, Nic79]. Multiple

[Bar78, Gla74a, Hen77a, McD74, BN70, Car73]. multiplication [BH74, BvT78, Gen72, McC77]. Multiplicative [BRW71, BW96]. multipliers [Fal76a, Kol79]. Multiply [Ros75]. multipoint [LP74]. Multistep [Bru72, Gup78a, Jel76, Jel77, Jel78, Lam70, Ske79, Wd74, Alf79b, Alf79c, GW75, Gup76, GW79, Hil75, Hil76, SL74, Wie76, Zla75, dW77, Dou79]. Multistep-Galerkin [Dou79]. Multivariate [AB72, GW78, McK72, WR75, Hen77b, Wan76, Wan78].

N [Sha71b]. N.D. [PK72]. N.J [Abr74, Bre75c]. Nat [Str71]. Nathanson [Wag79b]. National [P.71, Rhe72]. natural [Hil75, Hil76, Mac73]. Near [Won75, Coh73, Joh74a, Lyn71b]. Nearest [WB79, Ada79, Wil76a]. Necessary [Jel76, Bur78, Mah75]. Neergaard [Sha71b]. Negative [Cha74]. Neumann [For72, Gir76, GN78, Hic78, Mon79]. neutral [CG73]. Neutron [Mad73]. Newton [Bro70, DM74, Gre72, Gre78, Kam70, Kin71, Lam78, Nin70a, Phi70, Sch78c, Sch70b, Sha70, SK70]. Nicolson [For73]. Niklaus [Abr74]. Nine [BDS74, PH75]. Nine-Point [BDS74]. No [GR79, Gre72, BEJ78a, Ern79, Gay79, God79, Wil79]. Nodal [MS75]. Node [Bar78]. nome [FNC75]. non [TW74b]. non-smooth [TW74b]. Nonabelian [FM78, MY79]. nonconforming [Bak77]. Nonelliptic [Gir76]. Nonexistence [Gau75, DR72, NS74b]. nonexpansive [Dot78]. Nonhomogeneous [Kin74]. Nonlinear [Bro71, Car70, CSW77, Fat78, For73, Lam78, MM75, Rhe72, Rhe76, Ric71, Sch70b, ST70, Bog76, BD76, Bro79, Den71, Dot78, DD75a, DD75b, Gay77, Gay79, Hou77, Kel75, LP74, Nev78, Sid79, Sik79, Sim72, SW79c, Ste76]. Nonnegative [dD74, RH72]. Nonnormal [Par74]. nonpolygonal [Man78]. Nonresidues [RJ70]. nonselfadjoint [BO73]. Nonseparable [Ehr79]. Nonsingularity [Hou72]. Nonsmooth [MMO78]. nonuniform [Nat77]. nonzeros [Duf74]. Nordsieck [Gup78a]. Norm [AG79b, BNS75, RR70b, SW78, SW79a, Var70b, Cul79, HM77, Kah72a, SW77, Sch78c, de 76].

Normal [Bra71, Car75b, Dre78, AD73, For72, Lus79]. Normality [Oka76]. Normalizer [BNZ73]. Norms [Kin72, LF72, Coh73, GP79]. Note [AC74, Bla76b, Bri72, Bru72, Bui79, CS73, Ger75, Gla79b, God78, Gou70, Het73, Hud73, Knu70, LF71, Leh70, Mal71, Mon76, Mor75, OS72, Phi79, PB74b, WS79b, te 73, BB74, CL78, Cox72, Eil72, God79, Gol73, HR72, LK73, PB74a]. Notice [Ano71d]. Number [BE75, BWW77, Cha73, Ger75, Hen74, McD71, Par78, Pom79, RG70, SW73, Sta75, Wil77, WS79b, BWB76, CS75b, Duf74, HM73, HM75, Hud76, Pet73, PZ77, Pom75, SR76, Var77, Wan76, WB76, Wil79]. Numbers [AC74, BE74, BLM70, Coh70, Die71, EH79, God78, Hag70, Hag71, Hag73, HL77, HB75a, How77, Joh74b, Leh77, LM78, Nie72, RP71, Ros76, Shi78, WZ72, Bor72a, Boy78, Bue77, CGZ72, Cha79, God79, Hen75, Kis77, Lel75, Nie74, Nie76, Par77, Tuc73]. Numerical [BO78, BPW75, BG73, BC74, Bru70b, Bur71a, Car70, CA79, CC72, CH70a, Dan73, Dor71, EF79, Fat78, GP70, GN78, HKK72, Het73, HSK74, I.75, JP74, JT74, Kah72b, Kro70, Le 77, Lin70, Lyn71a, Mad73, MM75, Mil71, NGF74, NRRZ71, P.71, Per70, Rak75, Rhe72, RP71, Sac71, Sni76b, SW79d, Tod70, Toi78, Web73, Wil72, AL77, Alf79b, Alf79a, BN70, BR72, Bru74b, But73, BC73, Cas78, Chi72, EM77, FO77, Gal78, HKK75, JP72, Let77, MM74, ML79, MH77, NO76, OD78, PK72, PW76, Rak78, Sko75, SW79c, Var74, S.75]. Numerische [S.78].

O.D.E. [Ise77]. Oak [P.71, Rhe72]. Oberhettinger [Eil71]. observation [Sch74]. Observations [GR70, WH78, Bur75]. observed [Nic78b]. obstacle [BF77]. Obtaining [Fra71a]. occurrence [Bre73]. occurring [Kad76]. oceans [Lus79]. Odd [Hag73, Kis77, Kis78, McD71, HM73, HM75, Pom75, Tuc73]. Off [Dia78a]. Office [Fet71a, Fet71b]. Ohio [Fet71a, Fet71b]. Olver [Cas78]. One [HS75a, Hen77a, Hul72b, Lew70, Lir71, RB73, Wat78, Wil77, WS79b, Bar76, Cat77, CMPY78, Hul72a, PR74, Wat73, Wil79]. one-dimensional [Bar76]. One-Half [Lir71]. One-Sided [Lew70, Wat73]. One-Step [Hen77a, Hul72b, Wat78, Hul72a, PR74]. Only [GH79]. Operational [BC79]. operator [CL78, NW74, Slo76, WZ78]. Operators [Bon71, Osb75, RV73, Bra73b, Dot78, Ral76]. Opposite [Hag70]. Optimal [BR79, Bru72, DDW75, Dur71, Fit74, Gaf78, Gen72, Lar70, PB74b, Sco76, SK70, McC77, PB74a]. optimally [Sal74]. Optimization [Toi78, Gol76a]. Optimizing [BH77]. Order [Alo76, BS71, Car75b, DO74, EP75, EH79, Jer77, JR71, Kin74, Kut71, Lau73, Lir71, Oli75, RP70, ST70, ZA71, van74, AG73, ACR79, Bak75, BCSW78, Bey79, Bla76a, BS77, BvT78, But73, CG73, CL75, Gla74b, Gup76, LP74, LR79b, Man78, Oli74, PPW77, RT77, Ste76, Tho77, Ven73, WF74]. Ordered [RV73]. Orders [Fet71a, Fet71b, ND70, Cat77, Ker72]. Ordinary [Bru70b, Fat78, KD71, Mir71, Alf79b, Bey79, Bru74b, But73, Esp75, GW75, Gup79, Hul72a, Jai72, Kre72a, Nos73, PR74, SL74, Sha73a]. ordinate [Mad72]. Origin [Won75]. Orthogonal [Fra71a, Gal71, Gau72, Hun75b, Hun74a, Sid77]. Orthogonalization [PS79a]. Oscillator [NGF74]. oscillatory [MW76]. Osculatory [Boy73, Sal77, Wuy75]. Other [Coh76a, Fra71a, MR71, For72]. overdetermined [Abd77, And75, Bog74]. Overrelaxation [Had78, Kin72, You70, KY72].

P [I.72]. Packing [Boy70, Boy73]. Padé [Sid77]. Pairs [CP79, Die71, Ise78, MY79, te 74, Rio75]. Paper [Mer72, Gau79b]. paperbound [I.75]. Parabolic [Bon71, BT72, CP70, Car70, DDW78, Kin74, Kow71, Le 79, Sac71, WR71, Zla74, BBT77, BF77, BC73, DD73, DD75a, EF79, FT73, Osh72, TW74b, Tho79, Ven73, Zla75]. Parallel [Ric71, Sam71, Swa79]. Parameter [EP75, PS78]. Parameters [Won73, Bus74, KY72]. parenthesis [Bau77]. parenthesis-free [Bau77]. Parity [Hag70]. parking [LG74]. Part [Dur71, Lyn70, Lyn71b, Lyn71c, SW78, SW79a]. Partial [Boa77, For73, BvT78, Bur78, Bus74, EF79, FT73, Gir76, GP79, Hed75, Sni76b]. Partially [RV73]. Partially-Ordered [RV73]. particular [ML79]. Partition [Gup78b, Wil74, FNC75]. Partitioning [Wei72a].

Partitions [CS75a, Knu70, SS78, CC72]. paths [Lel75, Pet73]. Patterns [Wil78]. Patterson [Fet71a, Fet71b]. Paul [JR70]. Peaceman [Wid71]. PECE [KD71]. Penalty [Bab73, KS78]. percentage [Fet79]. Perfect [Alt71, Hag73, McD71, HM73, HM75, Kis77, Pom75, Tuc73]. performed [Duf74]. periodic [Bur75, Pag73]. Periodicity [Jel78]. Permutation [Atk75]. Perron [RV73]. persymmetric [Gol74a, Gol74b]. Perturbation [Dor71, Pai79, KT78]. Perturbed [NGF74, BD76]. Peter [S.78, Tha77]. Phase [CH70b, Phi79]. PIC [Mar75]. PIC-type [Mar75]. Picard [DM70]. Piecewise [Hen77a, Hul72b, Lyn71b, MS75, Var77]. Pisot [Boy78]. Pivot [Ste74]. Planar [Fra71a, PH75]. Plane [SW78, SW79a, EW76, Joh74a]. plastic [BJM77]. Plateau [HSK74]. plates [BJM77]. plethysm [Esp75]. Point [BDS74, DDW75, Fat78, Liv75, Rob71, dS77, AI75, Als74, Kel75, Lyn76, RV75, Ste76, Var74, Wei75b, Wil76a]. Points [FCC73b, Hud71, BH77, Fet79, KT78, MM74, Rio75]. Poisson [Ban78, Día79, Lyn70, Lyn71b, Lyn71c, Sko75, van74]. Poles [Lyn71b]. Polygonal [SW78, SW79a, BDT72, Lel75, Pet73]. polyhedral [Wil76a]. polylogarithm [FK75]. Polynomial [BG75, Blu79, Coh73, CH74, FI75, Gup78a, Hen77a, Hul72b, Kro70, New74, New75a, Rum79, Sch70d, Ste69, Ste71c, Ste71d, Wim62, Abe73, Bro75, Car73, Gal78, Gau79b, Gen72, Gup79, How78, Moe77, OD78, Wan78, Wim72a]. Polynomials [BPW75, BW74b, BG70, Ber70, Coh76b, Coh79b, ET73, Ell71, Fra71a, Gal71, Gau72, Gau79a, Hay73, Hun75b, Kah71, KR79, McD74, Mig74, MS75, Mos72, NRRZ71, Ris73, Sta73a, WR75, Wim75, Apo75, Cha79, Coh77, CR72, Kam73, MH77, Sid77, SS72, Var77, Wan76, Yan79]. Poorly [Var70a]. portions [Joh74a]. Posed [Fra74]. Positive [OSV79, RV73, Ris73, Bar77, Pag73, Wil76a]. positive-definite [Pag73]. Positivity [Mon78a]. possibly [Gay77, Gay79]. Posteriori [Car70, Mad73, Mie77]. Potential [GN78]. Power [Chi73, Gau79a, McD71, RJ70, Tur75, ZP75, Fra79, Gen72]. Powers [Bru76, Cha74, CS75b]. pp [Abr74, I.75, P.71, Rhe72, S.75]. Practical [Bur71b, Nic79]. preceding [BC73, Gau79b]. Precision [Sha74a, SG72, Gab79, Hun75a, Sha78].

Predictor [GH72]. Predictor-Corrector [GH72]. Prentice [Abr74, Bre75c]. Prentice-Hall [Abr74, Bre75c]. Prescribed [LLS70]. presentation [SS72]. Preserving [MPR77]. Press [Mac72b]. Price [I.75, P.71, Rhe72]. Primal [RT77]. Primality [Mor75, WH78, BLS75, WJ76a]. Primary [P.71]. Prime [Bri72, CP79, EES74, EGRS75, Hag70, Joh74b, Leh77, McD71, McD74, New70b, Pom79, RG70, Wei74a, WZ72, WJ76b, CS75b, HM73, HM75, Kis78, Pom75]. Primes [AG77, Bai79, BH78b, Bre74, Cad71, EM78, GH79, Joh75, JR70, MW77, Wag78, Wei77, Wil74, Wil78, WS79a, Wun73, Zwi79, Bre73, Bre75b, DMP74, Ern79, Kis77, Wag79a]. Primitive [BW74b, Sta73a]. principal [EP79]. principles [Bra73b, I.75]. Printed [P.71, Rhe72]. Priori [Bon71]. Probabilistic [Cha73]. Probability [Car75b, CMPY78, McN73]. Problem [AG79b, BT72, Cra78, FB78, Gek74, GN78, Hal78, McA71, Rob71, Ruh74, Sac71, SS70, Var70b, Web73, Wuy75, Abd75, Ans78, Bak73, BDT72, DD75b, EF79, GJ74, HSK74, KT78, LG74, MW76, Oli74, Oli76, PPW77, Ste75, TW74a, Usm78]. Problems [Ans76, Ban78, BR76, BS71, Bra77, CP70, CS71, Cry71, Dor71, DDW75, FGH^{+}74, Fra74, GT76, Gou70, Hen74, Hen77a, Hul72b, Kin74, Kra78, Le 79, Mor73, Pai79, Per70, Red76, Ste71b, dS77, AG79a, AI75, AL77, Alf79b, Alf79a, Als74, ACR79, BO78, BBT77, Bay78, Bey79, Bog76, BR74, BC73, FO77, GT78, Gun77, GKS72, Gus75, Hou77, Joh74a, Kad76, Kel75, Kre72a, LP74, Man78, MM74, Osh72, Pru73, RV75, Sal74, SW79b, Ser75, Sim72, Skö79, Ste79, Ste76, Var74, Wah75, Wei74b, Wei75b, Zla75, dW77]. Procedure [Bui79, Bur70, Car74, Sni76a, SD77, Tuc73, Wah75]. Procedures [DM70, DD75a, Gro75]. process [Dot78]. Product [Let70, NS74a, Sto78, Wei72b, dW73, Gla74b]. Products [LF72, Zar76, DF77, Esp75, LK73, New76, PS73, Mac72b]. Programming [Abr74, AG79b]. Programs [Knu75]. Progression [Wei77]. Progressions [GR79, GH79, Wag79a]. Projection [BDT72, DDW78, Moc76]. projective [DF77]. Proof [Kea78]. proofs [Bak73]. Properties [Erd76, Gol76b, HP72, Hun75b, RR70a, Sch76a, Wim75, You70, Sid79]. Property [McC78]. Pseudo [BRW71, Die71, Nie72, BW96, FT73, Nie74, Nie76]. pseudo-parabolic [FT73]. Pseudo-Random [BRW71, Die71, Nie72, BW96, Nie74, Nie76]. pseudoinverse [Gol74b, SPB75]. Pseudoperfect [BE74]. Pseudospline [Tho76]. Psi [CST73]. Pure [EFO78, Wil77, WS79b, BWB76, Cio79, JH78, Wil79].

QR [DGKS76]. Quadratic [Bue76, Cha71, Dia78b, DSW79, Gup71, LLS70, LL74, MR78, NS74a, Par78, SS73, Sta75, Wil74, Bue77, FHP78, Hen75, Par77, Sha76b, WB76]. quadratically [Ral76]. Quadrature [Bar78, Bas70, Dan73, Dur71, FM71, Gau70, Gla79a, GS70, Gus70, Kah70, Kah71, Kah72b, Kam70, KN70, Kum74, KJ74, LS72, LTK75, Lyn71a, Mil71, Mon76, PB74b, RR70c, Sal71a, Sal73, SS74, Sil70, Str72, Tua71, Wil70, EP79, Lyn76, Mon78a, Mon78b, PB74a, de 79]. Quadratures [GY74, Gau75, Het72, JR71]. quartic [Cio79, Lak74, Lak75]. Quasi [Bro70, DDW78, Gre72, Gre78, HL77, Lam78, Le 77, Sch70b, Sha70, SK70, Ste71b, AG73, DM74, Sch78c]. Quasi-Amicable [HL77]. Quasi-Linear [Le 77, Ste71b, AG73]. Quasi-Newton [Bro70, Gre72, Gre78, Lam78, Sch70b, Sha70, SK70, DM74, Sch78c]. Quasi-Projection [DDW78]. quasilinear [DD73]. question [NO76]. questions [TW74a]. Quintic [Wil76c]. Quotient [Par74]. Quotients [Joh78].

R [BW96, KS79, Lun72, Sha71b, Sha76a]. R. [S.75]. Rachford [Wid71]. Ramanujan [Leh70, Spi73]. Random [BRW71, Die71, Nie72, AD73, BW96, Duf74, For72, Mon79, Nie74, Nie76]. range [Lew75]. Rank [Bas72, Bro70, DSW79, EFO78, NS74a, PP74, PP75]. Rate [BO73, Cli72, Gus75, Nic78b]. Rates [DM71, NW74, TW74b]. Rathbone [Sha76a]. ratio [de 76]. Rational [BM70, Bla74, BEJ76, BEJ78a, BEJ78b, Chi73, CDR71, Dun75b, Dun76, Eck76, KLT78, LR73, Nin70a, Nin70b, Rei71, Sha76a, Smy78, Wol75, Wuy75, Hen77b, LR79a]. Rationals [Dun75a, Dun75c, TW74a, Wil72]. Ratios [Amo73, Amo74, Gau77, GS78]. Rayleigh [Hal78, Par74, Sch70c]. Real [Ang76, DiD78, Mor79, Par77, Boy78, CGZ72, Hen75, PS78, WB76]. Reciprocal [How77]. Reconsidered [SL72]. Rectangles [Fra71a]. Rectangular [Ple72, Día79]. Recurrence [Kro79, OS72, Sch77, Cas78, Dum79, Sal76b]. Recurrences [Mil75b, Tha79, FT77]. Recursive [Bru74b, McK72]. reduced [Cia78, Zla78]. Reduction [Gla74a, Gol74a, Gol74b, JS78, Sha71c, Alf79a, Don79, How73a, How73b]. reductions [Esp75]. Refinement [BKO73]. Refinements [SW79a, Ven73]. reflection [Sni76b]. region [BH78a, Fet74]. Regions [Fra71a, Fri70, Día79, OW79, PPW77, PH75]. Regular [Ike75, Sha76a, Ada79, Bre77]. regularization [Cul79]. Regulator [WB79]. Related [Alt71, MR71, RG70, You70, Cha79, Chi72, Hul72a, Joh77, Mac72a, New76, NO76, Sch78d]. Relations [Hud71, SW73, Cas78, Kro79]. relationship [Ada79]. Relative [Par78]. Relatively [Hag70]. Relevant [Ros71b]. Remainder [Tua71]. Remark [Mak70, Mer72, Zas78]. remarks [Mon78b]. Remes [KLT78]. Remes-Differential [KLT78]. Remez [Hud74]. Reorthogonalization [DGKS76]. Representation [Gup78a, Gup79, Lyn74]. Representations [Dix70, Ric71, Bus74, DF77, Esp75]. Representing [Bri72]. Reproducing [Lar70]. Res [Fet71a, Fet71b]. Residuacity [Wil76c]. Residues [Gup71, RJ70, SL72]. Respect [Lau73, Bus74]. Restricted [Lew75, Sch78c]. resultant [Apo75, How78]. Results [BPW75, Bor72b, BGW70, Coh76a, Toi78, Web73, CC72, FHP78]. Retrospect [MSW71]. Review [Abr74, I.72, I.75, P.71, Rhe72, S.75, S.78, Tha77]. Reviews [Ano71e, Ano71f, Ano72e, Ano75d, Ano75e, Ano76d]. Revision [Gre78]. Richtmyer [Hic78]. Ridge [P.71, Rhe72]. Riemann [Bre79, CHT71, GB79, Kol70, Kol72a, RG70]. Ritz [Sch70c, BNS75, Hal78, HS75b, NS74c, Sch74]. Romberg [Str72]. Roof [BW96]. Root [Cha71, CH74, Dut71, Nin70a, Phi70, Rum79]. Roots [BMN70a, BMN70b, Fet76, Kin71, Lir71, PS77b, PS77a, PS78, PS79b, Sha71b]. Rosenbrock [Bui79]. Rounded [Dia78a]. Roundoff [Ram71, Kah72a]. Rudolf [S.78]. Rule [Osl72a, Gla79a, RV72]. Rules [Gau70, Gus70, Mon76, RR70a, RR70b, Tua71, DR72, EP79, Mon78a, Mon78b]. Runge [CS79, Oli75, SW71, Var79, Wei74b]. Russian [Mac72b]. Rutishauser [S.78]. Ryzhik [Mac72b].

S [I.75, Mac72b, P.71, Rhe72]. Saddle [FCC73b]. Salem [Boy78]. sampling [AD73, For72, KMR77]. Santos [KS79]. Satisfy [Fal76b]. satisfying [Ral76]. scalar [Alf79a, RS78]. scale [Lyn79a]. scales [BS78]. Scaling [Ore73, Wid71]. scattering [Kad76]. Scheme [GT78, Vai70, Bay78, Hof79]. Scheme-independent [GT78]. Schemes [ABH73, BT71, Cim71, Ise77, Liv75, MM75, Mir71, New75a, Ros75, RP70, ZA71, AG79a, CH78, CL75, Har78, Hed75, Skö79, TW74b]. Schett [Dum79]. Schmidt [DGKS76]. Schwarz [S.78]. Scientists [S.75]. Search [CP79, Ore73, PP74, Tuc73]. Second [Gup78a, Kin74, RP70, Tro73, BCSW78, Gla74b, HKK75, HP78, Hud74, Net73, Pom75, Sag70, Ven73]. Second-Derivative [Gup78a]. Second-Order [RP70]. Seidel [Lar73]. selection [CR78]. Selective [PS79a]. Self [Har78, Ore73]. Self-adjusting [Har78]. Self-Scaling [Ore73]. Semi [KJ74, Kam73]. Semi-Infinite [KJ74, Kam73]. Semianalytic [MM74]. Semicardinal [SS74]. Semidiscrete [BT72, Kin74, Tho79]. Semidiscrete-Least [BT72]. Semidiscretization [Le 79]. Semiexplicit [CS79]. semigroups [Nev78]. Separable [FGH^{+}74]. separably [Alf79a]. Separation [CH74, Rum79]. Sequence [Gol76b, GS75, LLS70, Sid79]. Sequences [Erd76, Fia71, Mor75, Wag72, Wil76b, Wim72b, Bro79, Yan79]. Serial [BMN70a]. Series [Boa77, Chi73, Gla74a, New73, Osl72b, Spi73, WR71, Wre73, ZP75, Gol77b, Mac72b, Sal76b, Sch76a, Sch77, Sid77, TE72]. Service [P.71]. Set [Cho71, Hag73, Klo75]. Sets [GR79]. Seven [Liv75, PP75, PP77a]. Seven-Point [Liv75]. Seventeen [Wei77]. several [SP75]. Shape [MPR77]. Sharper [RS75b, Sch76b]. SHASTA [IN79]. Shepard [GW78]. Shift [CH70b]. Shifting [Ste71d]. shocks [Har78]. shooting [Bay78]. Shortest [Die75]. Shrinking [Ber75]. Shuffling [Ros75]. Side [RS75a]. Sided [Lew70, Wat73]. significant [Fet71a, Fet71b]. Simcha [Sha71a]. Simple [BM70, FM78, Hun74b, MY79, RV73, Cat77, Vau78]. Simplest [Sha74b]. Simplex [LP73, DR72, de 79]. Simplexes [Sil70, Str76c]. Simplified [Bak73, CHK79]. Simplifying [Osl75]. Simply [Bur70]. Simpson [RV72]. simulation [EM77]. Simultaneous [BS78, WBH75]. simultaneously [Abe73]. sinc [Ste79]. sinc-Galerkin [Ste79]. Single [BBT77, Bro70]. Single-Rank [Bro70]. Singular [Dor71, FM71, Red76, Bay78, Bow76, Gay77, Gay79, KT78, Lyn76, Mad72, ML79, dW77]. Singularities [GS70, Lyn71c, Let77]. Singularity [Mil71]. Six [PP77b]. Sixteen [Coh79a, EH79]. Sixth [Bru76, McD71]. size [GW79]. Slightly [NGF74]. Slowly [Ber75, LS72, Sco79]. Small [Bre74, Bue77, Der77, Wei75a]. Smallest [Tro73]. Smith [Bra71]. smooth [TW74b]. smoothing [Cul79]. Smoothness [CSW77]. Sobolev [Hil73]. Sociable [Coh70]. Solid [Knu70]. Solution [BP70, Car70, DM70, Dor71, GP70, GN78, Gou70, KD71, MM75, Mv77, Mic73, Pai79, Per70, Sac71, Tod70, Tur75, Var72, Wag79b, WH77, Abd77, Alf79b, Alf79a, Als74, BR72, BF77, Bog74, BR74, Bro75, Bru74a, Bru74b, Cas78, FO77, Gek72, Gol77b, GP79, Gut77, GK73, HSK74, JP72, JP74, MH77, Net73, PW76, Rak75, Ser75, Sha73a, Sin72, Sko75, SW79c, Ste79, WZ78]. Solutions [Bra77, Bru70a, DDW75, Fra71b, Hil73, Kra78, Mad73, Gay77, Gay79, NW74, Rak78, SW79b, Sha71a, Sim72]. Solver [GB79, ACR79]. Solvers [Ban78, Sha74a, PPW77, Sha78]. Solving [Bro71, BK77, Cry71, Dus73, Gol72, Jel78, Nic77, Pai73, Pru73, Swa79, Bow76, Gal78, Gri78, GW75, Gup76, Gup79, Hou77, Jes77, PR74, Sik79]. Some [AG74, BPW75, BW74b, BMN70a, BMN70b, Bor72b, BGW70, BK77, Coh77, Dia78b, FHP78, Gla74b, GR70, GW75, Gup76, Hen74, Hil73, HP72, Hud71, Hun75b, Jai72, Kah71, MW77, McN73, Mon78b, Par74, RC78, RG70, RP71, Ros71b, Sal74, Sam74, SP75, Str76c, Tho79, Toi78, WZ72, WJ76b, WH78, Wil78, WS79a, Wim62, Wim72b, Kad76, KT78, Nic79, Pag73, Sha71b, Sid79, Var74, Wim72a].

SOR [Ehr72]. sound [Car75a]. Sources [Sha78]. Space [Tho76, Win70, AL77, Alf79c, Dot78]. Spaced [Gre70b, Hos71, Mac73]. Spaces [Hil73, Lar70, RV73, BS78, Moc76, Nev78]. Spacings [Cim71]. Sparse [BGW70, Bro71, Lam78, Ruh74, Ruh79, Sch70b, Toi78, Duf74, Toi77]. Sparsity [Ban78]. special [Alf79c, Bus74, DC79, Hic78]. spectra [PS77a]. Spectral [Osb75]. speed [Fie77]. Sphere [Boy73, BH77]. spherical [PS77b, PS77a, PS79b]. Spline [CSW77, Gre70b, Hos71, Hos73, MPR77, Mic73, Net73, Pru79, Sch70d, Sch70c, dD74, CR78, GK73, Hil75, Hil76]. Splines [ABH73, MR78, SK72, Shr74, dD74, BCSW78, Ker72, Mac73, MS73b, Nat77, Sin72, de 76]. Splitting [Man71]. Springfield [P.71, Rhe72]. Square [Alo76, BMN70a, BMN70b, Dut71, Nin70a, Sha71b]. Square-Free [Alo76]. Squares [Ber75, BS71, BT72, Bri72, Car74, Kin74, Pai79, Sam77, Wil70, Yan76, Bak73, Bog76, Car75a, Jes77, NW74, Ser75]. Stability [AG79a, Bro79, Bru70b, Dia78a, FT73, Fra71b, GKS72, HS75b, Hic78, Ise78, Jel76, Osh72, Vai70, Var70b, Wie76, AO74, BD76, Bur78, Gol73, GT78, Gun77, How78, JT74, PR74]. Stabilization [Bru70b]. Stable [BK77, Cim71, CS79, DGKS76, GA74, Gus79, KR79, Rei71, Wd74, Fet74, Jai72, NS74b, Sal74]. Standards [Str71]. standing [PS77a]. Starting [Nin70a, Hud74, Sha78]. States [Fet71a, Fet71b]. Stationary [Fal76b, Fal76a]. Statistical [BMN70b, Sac71, Sha71b]. Stegun [Str71]. Step [Hen77a, Hul72b, Wat78, BBT77, Eil72, GW79, Hul72a, PR74]. step-size [GW79]. Stiff [EP75, Fat78, GH72, Ise77, KD71, Mir71, SF74, Sni76a, AL77, Alf79a, Bru74b, FO77, Gup76, Jai72, MM74, MW76, Nos73, PR74, SD77, SW79c]. Stirling [Spi71]. Stokes [Fal76a, Fal76b]. strict [Abd77]. strip [Bre79, LO79]. Strongly [Fia71]. Structure [BRW71, Rob75, BW96]. Student [KMR77]. studies [MM74]. Study [BMN70b, Sha71b]. Sturm [Gek74, Hal78, Red76]. Subgroup [Sha71c, JS78]. Subgroups [FM78, New75b, PP77a, PP77b, Sto78]. Subject [FS71, Toi77]. Submatrices [Yan71]. Subsets [Ros71a, Bra73a]. Subspaces [BG73, Var70a, Sch78c]. substitution [GJ74]. Subsums [BE75]. Subtracting [Let77]. Successive [Bre74, Kin72, You70, Bre73, KY72]. sufficient [Mah75]. Suggested [Mil75a]. Sum [Bri72, Tak78, Tur75, WCJ72, Yan76, Bus74, CS75b, DC79, Fra79]. Summable [Wal78]. Summation [Lyn70, Lyn71b, Lyn71c]. Sums [Boa77, Bru76, DMP74, Klo75, LF72, Lir71]. Superconvergence [Zla78]. superlinear [DM74]. Supplement [PB74b]. Surface [Kel79, JT75]. Sylow [JS78, Sha71c]. Symbolic [KP78, Gen72]. Symmetric [BT71, BK77, Lin70, Mv77, Ruh79, Sil70, You70, ZC70, Esp75, HM77, McK76, PH75, Rak78, Sch78c, Toi77]. Symmetrization [Tur73, WBH75]. symmetrized [Esp75]. symplectic [RS78]. Symposium [Gau94]. Synthetic [Ste71d]. System [Sni76a, Bur78, Hof79]. Systematic [Abr74]. Systems [Ans76, BP70, BT71, Bro71, BK77, DM71, Fat78, GP70, Gol72, KD71, Lam78, MM75, Mv77, Mir71, Rob75, ST70, Var72, AG73, And75, AO74, Ans78, ACR79, Bog74, BR74, Bro79, Den71, Gay77, Gay79, Gek72, Gol73, GA74, Gun77, GP79, JP72, LR79b, Nos73, PR74]. Szekeres [Cus77].

T [Ern79]. Table

[Ang76, Ano71g, Ano71h, Ano72f, Ano75f, Ano75g, Ano76e, Bre75c, Eil71, Fet71a, Fet71b, Gre70b, Hil76, Hos71, Hun74b, LM78, Mac72b, Mos72, New70b, Sha76a, Str71, Hun74a, Lon77, Mac72b]. Tables [Ano71e, Ano71f, Ano72e, Ano75d, Ano75e, Ano76d, BPD79, BW74a, BW76, Esp75, LV79, Luk71a, Luk71b, Luk72, BvT78, CH78, Str71, Fet71a, Fet71b]. Tabulation [ZE78]. tail [Fet74]. tail-end [Fet74]. Tau [Leh70]. Taylor [Gol77b, Osl72b, Sch76a, Sch77]. Technical [P.71]. Technique [Ehr79, LL74, AL77, Alf79c, BWB76, GW79, WB76]. Techniques [CP70, CH70a, KS78, Lyn74]. ten [Kis77]. Tennessee [P.71, Rhe72]. Tension [Pru79]. Tensor [LF72, PS73]. terms [de 76]. Ternary [Cus71, JS78, Sha71c, Cus72]. Test [BMN70a, Moc76]. Testing [Mor75, WJ76b, WH78]. th [BS71, Phi70]. Thabit [Bor72a]. Their [Amo73, BLM70, Ike75, LM78, Amo74, McN73, de 79]. Theodorsen [Gut77].

Theorem [Gal71, I.72, Str75, Gol77a, Hud76, Joh77, NS74b]. theoretical [Nic79]. Theory [But72a, Cha73, Dur71, Gal71, Mor73, RV73, Sac71, Str75, BK74, CC72, GKS72, Hud76, Ste75]. theta [FNC75]. Third [BW74b, Fri70, Gre70b, Hos71, ZC70, ZA71]. Third-Degree [Gre70b, Hos71]. Thirteens [Lan73]. Thompson [Ste71a]. Three [Boy73, Lan73, PP75, RP70, Eil72, Lon77, OW79, PS79b, Sik79]. Three-Dimensional [Boy73, RP70, OW79, Sik79]. Throw [Bur71b]. Throw-Back [Bur71b]. Tikhonov [Fra74]. Time [GT76, KR79, Le 79]. Time-Dependent [GT76]. times [BC73]. Tocher [Cia78, Man78]. Todd [CDHW73]. Toeplitz [FGH^{+}74, Ris73, Tre74]. Toroidal [Fet70, Fet71a, Fet71b]. Totally [Ang76, Bar77]. Totient [Lal74]. totitives [KS79]. transcendency [Mah75]. Transcendental [Lir71, Chi72, Eil71, PS77b, PS77a, PS78, PS79b]. Transform [Win78, KL79, Pol71, Ram71]. Transformation [Str76b, Str76a]. Transformations [Wim72b, FHP78, Sid79]. Transforms [Osl72b, Shr74, Won75, Mac72a, PS77c]. Transitive [McK79]. translation [Mac72b]. Transport [Mad73]. Trapezoidal [Gou70]. Traub [For77]. Treatment [Rhe72, BO78]. Trial [Fal76b, Moc76]. triangle [Cia78]. Triangles [RS75a, BG75, Sty79]. Triangular [BZ70, BH74, CS75a, Phi71, Ris73]. Tricomi [Eil71]. Tridiagonal [Ker70, Rei71, Sch70a, Ste76, Swa79, Var72]. Tridiagonalization [RH72]. Trigonometric [JH70, KN70, New70a, LK73]. trillion [Cat77]. Trinomials [BW76]. Triple [Rob75]. Triples [Bru76]. trouble [Sha78]. Truncatable [AG77]. truncation [BF77, McC74]. turning [KT78, MM74]. Twenty [McK79]. Twin [CP79, Zwi79, Bre75b]. Two [AG79b, BM70, Bas72, BK78, Bri72, Des74, DDW75, EP75, Fat78, GT76, HB75a, KS79, Rob71, Sta75, Wre73, Yan76, dS77, AG79a, AI75, Als74, Chi72, CS75b, Eil72, Gla74b, Hof79, Hud74, Kel75, Kol72a, ML79, PS77b, PS77a, PS78, RV75, Ste76, Var74, Wei75b]. two-dimensional [AG79a, ML79]. Two-Four [GT76]. Two-Parameter [EP75]. Two-Point [DDW75, Fat78, Rob71, dS77, AI75, Als74, RV75, Ste76, Var74, Wei75b]. two-step [Eil72]. Type [GY74, Gau75, Kah70, RR70b, AG79a, CG73, EP79, HM77, Man78, PB74a, Sch78a, Ser75, PB74b, Mar75].

U [P.71, Rhe72]. Unconstrained

[Bas72, Ore73, Toi78, Gol76a, Naz76]. Uneven [Cim71]. Unicity [BCSW78]. unified [SS72]. Uniform [Bel72, Ell71, MR71, Nat77, Tem75, Wei72a, NO76]. uniformly [AO74]. Uniqueness [RP71, Ros76, Sid77]. unit [BH77, JH78, Yan79]. Unitary [Hag71, Lal74]. United [Fet71a, Fet71b]. units [PZ77, SR76]. unstable [Bay78]. Unsymmetric [You70]. updates [Sch78c]. Updating [DGKS76, Toi78, Toi77]. upper [Fet79, Tun75]. Use [Osl72b, WB79]. Useful [BKV72, Hil73, Man78]. Using [Bur70, Bur71a, Der77, Dur71, Fal76b, Fra71a, GN78, Gup78a, Hos71, LL74, Mor75, Sal76a, Toi78, WJ76b, Zwi79, AG79a, Bak77, BR74, EW76, Fal76a, How73a, How73b, Jel77, JP72, McC77, WJ76a].

Va [Rhe72]. Value

[BS71, BT72, Bra77, CP70, CS71, DDW75, Gou70, Hen77a, Hul72b, Kra78, Per70, Rob71, SS70, Var70b, dS77, AG79a, AI75, AL77, Alf79b, Alf79a, Als74, ACR79, Bay78, Bey79, BDT72, BR74, Dun76, EP79, FO77, GT78, Gun77, GKS72, Gus75, Kel75, LP74, MM74, Oli74, Oli76, Osh72, Pru73, RV75, SW79b, Ser75, Skö79, Ste79, Ste76, Usm78, Var74, Wei74b, Wei75b, Zla75, dW77]. Valued [Bel74, BDM78, Bel72]. Values [Hud71, LV79, Bue77, Gay77, Gay79]. Vancouver [Gau94]. Vandermonde [BP70, GP70]. vanishing [Dun76, Joh73]. Variable [Gol70, Gre70a, Ore73, Won73, BCSW78, Gol76a, LP74, Usm78]. Variable-Metric [Gol70, Gre70a]. Variables [Des74, Mon79, SP75]. Variance [Dur71]. Variational [DM71, Gol70, BF77, Fal74]. Variations [Gre70a]. Various [BMN70a, BMN70b, Sha71b]. Varying [Lam70, LV79]. Vector [Ans76, Ans78, Bel74, War76, Bel72, FT77, KL79].

Vector-Valued [Bel74, Bel72]. Vectors [BRW71, Die75, Sha73b, BW96]. version [Alf79a]. Vertices [Fit74, McK79]. vii [P.71, Rhe72]. viii [P.71]. Virginia [P.71]. Virtual [FB78]. viscosity [Hic78, Hof79]. viscous [WF74]. Vogelaere [CM78]. Vol [Eil71]. Vols [S.78]. Volterra [Bow76, Bru77, Gla79a, Gol77b, GK73, Net73]. Volume [P.71, P.71]. Volumes [AMS84]. Vorlesungen [S.78]. Vortex [Dus73, Hd78].

W [BW96, Eil71, S.75, Sha71b]. Waerden [SS78]. Waring [Web73]. Washington [Str71]. Wave [Car74, Ike75, Pex70, SG72, CL75, CL78]. waves [EM77, PS77a]. Weakly [FM71, Bow76]. Weierstrass [Eck76]. weight [Dun76, Kam73]. Weights [Sal71a, Mon78a]. Weird [BE74]. Weisner [Coh76a]. well [Bau77]. well-formedness [Bau77]. Wendroff [Eil72, Vai70]. Weyl [BPD79]. Which [Mv77, Fal76b, Kis78]. Whittaker [Lau73, MSW71]. whose [Lyn70]. Wiener [CS73, Cho73]. Williams [Dun75c]. Williamson [BW96]. winding [Lel75, Pet73]. Wirth [Abr74]. without [Bre75c, KT78, Naz76, Ore73]. WONDY [Hic78]. wreath [DF77]. Wright [Fet71a, Fet71b]. Wright-Patterson [Fet71a, Fet71b]. Wynn [Gek72].

x [I.75].

York [Eil71, I.75, Mac72b, S.75].

Z [P.71]. zero [Cox72, Dun76, Gla74b]. zeroes [CMPY78]. Zeros [BK78, CH70a, Ike75, Kol70, Kol72b, Spi76, Ste69, Ste71c, Ste71d, Wag75, Wei75a, Abe73, Bre79, Car73, CK79, FCC73a, FK75, HB75b, Kol72a, LG79, Par72, Ral76]. Zeta [CHT71, Kol70, Sha75a, Spi76, Bre79, Kol72a, Sha75b].

References

Amos:1972:CMD

[AB72] D. E. Amos and W. G. Bulgren. Computation of a multivariate F distribution. Mathematics of Computation, 26 (117):255–264, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abdelmalek:1975:EMD

[Abd75] Nabih N. Abdelmalek. An efficient method for the discrete linear L^{1} approximation problem. Mathematics of Computation, 29 (131):844–850, July 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abdelmalek:1977:CSC

[Abd77] Nabih N. Abdelmalek. Computing the strict Chebyshev solution of overdetermined linear equations. Mathematics of Computation, 31(140):974–983, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aberth:1973:IMF

[Abe73] Oliver Aberth. Iteration methods for finding all zeros of a polynomial simultaneously. Mathematics of Computation, 27(122):339–344, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Andria:1973:IFS

[ABH73] George D. Andria, George D. Byrne, and David R. Hill. Integration formulas and schemes based on g-splines. Mathematics of Computation, 27(124):831–838, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abrahams:1974:SPI

[Abr74] Paul Abrahams. Book review: Niklaus Wirth, Systematic Programming: An Introduction, Prentice-Hall, Inc., Englewood Cliffs, N.J., 1973, 167 pp., 24 cm. Mathematics of Computation, 28(127):881–883, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alter:1974:NCN

[AC74] Ronald Alter and Thaddeus B. Curtz. A note on congruent numbers. Mathematics of Computation, 28(125):303–305, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ascher:1979:CSM

[ACR79] U. Ascher, J. Christiansen, and R. D. Russell. A collocation solver for mixed order systems of boundary value problems. Mathematics of Computation, 33(146): 659–679, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ahrens:1973:EFM

[AD73] J. H. Ahrens and U. Dieter. Extensions of Forsythe’s method for random sampling from the normal distribution. Mathematics of Computation, 27(124):927–937, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adams:1979:RBC

[Ada79] William W. Adams. On a relationship between the convergents of the nearest integer and regular continued fractions. Mathematics of Computation, 33(148):1321– 1331, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abarbanel:1973:HOA

[AG73] S. Abarbanel and D. Gottlieb. Higher order accuracy finite difference algorithms for quasi-linear, conservation law hyperbolic systems. Mathematics of Computation, 27(123):505–523, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Angell:1974:SF

[AG74] I. O. Angell and H. J. Godwin. Some factorizations of 10^{n} ± 1. Mathematics of Computation, 28(125):307–308, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Angell:1977:TP

[AG77] I. O. Angell and H. J. Godwin. On truncatable primes. Mathematics of Computation, 31(137):265–267, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abarbanel:1979:STD

[AG79a] Saul Abarbanel and David Gottlieb. Stability of two-dimensional initial boundary value problems using leap-frog type schemes. Mathematics of Computation, 33 (148):1145–1155, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Armstrong:1979:TLP

[AG79b] Ronald D. Armstrong and James P. Godfrey. Two linear programming algorithms for the linear discrete L_{+}1 norm problem. Mathematics of Computation, 33 (145):289–300, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ahlberg:1975:CMT

[AI75] J. H. Ahlberg and T. Ito. A collocation method for two-point boundary value problems. Mathematics of Computation, 29(131):761–776, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alfeld:1977:CDS

[AL77] P. Alfeld and J. D. Lambert. Correction in the dominant space: a numerical technique for a certain class of stiff initial value problems. Mathematics of Computation, 31(140):922–938, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alfeld:1979:IVR

[Alf79a] Peter Alfeld. An improved version of the reduction to scalar CDS method for the numerical solution of separably stiff initial value problems. Mathematics of Computation, 33(146):535–539, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alfeld:1979:ILM

[Alf79b] Peter Alfeld. Inverse linear multistep methods for the numerical solution of initial value problems of ordinary differential equations. Mathematics of Computation, 33(145):111–124, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alfeld:1979:SCE

[Alf79c] Peter Alfeld. A special class of explicit linear multistep methods as basic methods for the correction in the dominant space technique. Mathematics of Computation, 33(148):1195–1212, October 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alonso:1976:GSF

[Alo76] J. Alonso. Groups of square-free order, an algorithm. Mathematics of Computation, 30(135):632–637, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alspaugh:1974:AII

[Als74] Dale W. Alspaugh. The application of invariant imbedding to the solution of linear two-point boundary value problems on an infinite interval. Mathematics of Computation, 28(128):1005–1015, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alter:1971:DER

[Alt71] Ronald Alter. On a Diophantine equation related to perfect codes. Mathematics of Computation, 25(115):621–624, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Amos:1973:BIC

[Amo73] D. E. Amos. Bounds on iterated coerror functions and their ratios. Mathematics of Computation, 27(122):413–427, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Amos:1974:CMB

[Amo74] D. E. Amos. Computation of modified Bessel functions and their ratios. Mathematics of Computation, 28(125):239–251, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

AMS:1984:CIM

[AMS84] AMS. Cumulative Index to Mathematics of Computation, Volumes 24–43, 1970–1984. Mathematics of Computation. American Mathematical Society, Providence, RI, USA, October 1984. CODEN MCM-PAF. ISBN 0-8218-0093-0. ISSN 0025-5718 (print), 1088-6842 (electronic). vii + 503 pp.

Anderson:1975:GMA

[And75] Ned Anderson. A generalization of the method of averages for overdetermined linear systems. Mathematics of Computation, 29(130):607–614, April 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Angell:1976:TTR

[Ang76] I. O. Angell. A table of totally real cubic fields. Mathematics of Computation, 30(133):184–187, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1970:Ca

[Ano70a] Anonymous. Corrigenda. Mathematics of Computation, 24(109):243, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1970:Cb

[Ano70b] Anonymous. Corrigendum. Mathematics of Computation, 24(112):1001, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1971:Cb

[Ano71a] Anonymous. Corrigenda. Mathematics of Computation, 25(114):409, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1971-25-114/S0025-5718-71-99717-1.

Anonymous:1971:Cc

[Ano71b] Anonymous. Corrigendum. Mathematics of Computation, 25(115):637, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1971:Cd

[Ano71c] Anonymous. Corrigendum. Mathematics of Computation, 25(116):947, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1971:N

[Ano71d] Anonymous. Notice. Mathematics of Computation, 25(114):411, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-71-99716-X.

Anonymous:1971:RDTa

[Ano71e] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 25(114):393–404, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-71-99719-5.

Anonymous:1971:RDTb

[Ano71f] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 25(115):627–633, July 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/1971-25-
115/S0025-5718-71-99712-2;   http:
//www.ams.org/journals/mcom/1971-
25-115/S0025-5718-71-99712-2/S0025-
5718-71-99712-2.pdf.

Anonymous:1971:TEa

[Ano71g] Anonymous.  Table errata.  Math-

ematics of Computation, 25(114):405– 408, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-71-99718-3.

Anonymous:1971:TEb

[Ano71h] Anonymous. Table errata. Mathematics of Computation, 25(115):635, July 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/1971-25-
115/S0025-5718-71-99711-0;   http:
//www.ams.org/journals/mcom/1971-
25-115/S0025-5718-71-99711-0/S0025-
5718-71-99711-0.pdf.

Anonymous:1972:Ca

[Ano72a] Anonymous. Corrigenda. Mathematics of Computation, 26(117):309, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1972:Cb

[Ano72b] Anonymous. Corrigendum. Mathematics of Computation, 26(118):601, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1972:Cc

[Ano72c] Anonymous. Corrigendum. Mathematics of Computation, 26(119):815, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1972:Cd

[Ano72d] Anonymous. Corrigendum. Mathematics of Computation, 26(120):1031, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1972:RDT

[Ano72e] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 26(117):287–303, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1972-26-117/S0025-5718-72-99110-7;
 http://www.ams.org/journals/mcom/
1972-26-117/S0025-5718-72-99110-7/
S0025-5718-72-99110-7.pdf.

Anonymous:1972:TE

[Ano72f] Anonymous. Table errata. Mathematics of Computation, 26(117):305–307, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1972-26-117/S0025-5718-72-99109-0;
 http://www.ams.org/journals/mcom/
1972-26-117/S0025-5718-72-99109-0/
S0025-5718-72-99109-0.pdf.

Anonymous:1973:Cb

[Ano73a] Anonymous. Corrigenda. Mathematics of Computation, 27(123):683, July 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1973:Ca

[Ano73b] Anonymous. Corrigendum. Mathematics of Computation, 27(122):453, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1974:Cb

[Ano74a] Anonymous. Corrigenda. Mathematics of Computation, 28(127):887, July 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1974:Cc

[Ano74b] Anonymous. Corrigenda. Mathematics of Computation, 28(128):1183–1184, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1974:Ca

[Ano74c] Anonymous. Corrigendum. Mathematics of Computation, 28(125):343, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1975:A

[Ano75a] Anonymous. An addendum to 12,758. Mathematics of Computation, 29(130):675, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1975:Ca

[Ano75b] Anonymous. Corrigenda. Mathematics of Computation, 29(130):673–674, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1975:Cb

[Ano75c] Anonymous. Corrigenda. Mathematics of Computation, 29(132):1167, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1975:RDTa

[Ano75d] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 29(130):648–669, May 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1975-29-130/
S0025-5718-75-99678-7/S0025-5718-
75-99678-7.pdf.

Anonymous:1975:RDTb

[Ano75e] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 29(132):1152–1165, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1975-29-132/S0025-5718-75-99674-X;
 http://www.ams.org/journals/mcom/
1975-29-132/S0025-5718-75-99674-X/
S0025-5718-75-99674-X.pdf.

Anonymous:1975:TEa

[Ano75f] Anonymous. Table errata. Mathematics of Computation, 29(130):670, May 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1975:TEb

[Ano75g] Anonymous. Table errata. Mathematics of Computation, 29(132):1166, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1975-29-132/S0025-5718-75-99673-8;
 http://www.ams.org/journals/mcom/
1975-29-132/S0025-5718-75-99673-8/
S0025-5718-75-99673-8.pdf.

Anonymous:1976:Ca

[Ano76a] Anonymous. Corrigenda. Mathematics of Computation, 30(133):198, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1976:Cc

[Ano76b] Anonymous. Corrigenda. Mathematics of Computation, 30(136):900, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1976:Cb

[Ano76c] Anonymous. Corrigendum. Mathematics of Computation, 30(135):679, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1976:RDT

[Ano76d] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 30(134):371–380, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1976-30-134/
                                              S0025-5718-76-99667-8.

Anonymous:1976:TE

[Ano76e] Anonymous. Table errata. Mathematics of Computation, 30(134):381, April 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/journals/mcom/
                                              1976-30-134/S0025-5718-76-99666-6.

Anonymous:1977:A

[Ano77a] Anonymous. Acknowledgement. Mathematics of Computation, 31(137):331, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1977:Ca

[Ano77b] Anonymous. Corrigendum. Mathematics of Computation, 31(137):330, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1977:Cb

[Ano77c] Anonymous. Corrigendum. Mathematics of Computation, 31(138):617, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1977:Cc

[Ano77d] Anonymous. Corrigendum. Mathematics of Computation, 31(139):808, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anshel:1976:DPG

[Ans76] Michael Anshel. Decision problems for HNN groups and vector addition systems. Mathematics of Computation, 30 (133):154–156, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anshel:1978:VGE

[Ans78] Michael Anshel. Vector groups and the equality problem for vector addition systems. Mathematics of Computation, 32(142):614–616, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anderssen:1974:SUA

[AO74] R. S. Anderssen and B. J. Omodei. On the stability of uniformly asymptotically diagonal systems. Mathematics of Computation, 28(127):719–730, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Apostol:1975:RCP

[Apo75] Tom M. Apostol. The resultant of the cyclotomic polynomials F^{m}(ax) and F^{n}(bx). Mathematics of Computation, 29(129):1–6, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkinson:1975:AFB

[Atk75] M. D. Atkinson. An algorithm for finding the blocks of a permutation group. Mathematics of Computation, 29(131):911–913, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babuska:1972:FEM

[Bab72] Ivo Babuška. The finite element method for infinite domains. I. Mathematics of Computation, 26(117):1–11, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babuska:1973:FEM

[Bab73] Ivo Babuška. The finite element method with penalty. Mathematics of Computation, 27(122):221–228, April 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baillie:1979:NPF

[Bai79] Robert Baillie. New primes of the form k ·2^{n} +1. Mathematics of Computation, 33 (148):1333–1336, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1973:SPE

[Bak73] Garth A. Baker. Simplified proofs of error estimates for the least squares method for Dirichlet’s problem. Mathematics of Computation, 27(122):229–235, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1975:FEM

[Bak75] Garth A. Baker. A finite element method for first order hyperbolic equations. Mathematics of Computation, 29 (132):995–1006, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1977:FEM

[Bak77] Garth A. Baker. Finite element methods for elliptic equations using nonconforming elements. Mathematics of Computation, 31(137):45–59, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Banegas:1978:FPS

[Ban78] Alexandra Banegas. Fast Poisson solvers for problems with sparsity. Mathematics of Computation, 32(142):441–446, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrow:1976:EGI

[Bar76] David L. Barrow. Existence of Gauss interpolation formulas for the one-dimensional heat equation. Mathematics of Computation, 30(133):24–34, January 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrow:1977:GIF

[Bar77] David L. Barrow. Gauss interpolation formulas and totally positive kernels. Mathematics of Computation, 31(140):984–993, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrow:1978:MNG

[Bar78] David L. Barrow. On multiple node Gaussian quadrature formulae. Mathematics of Computation, 32(142):431–439, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Basu:1970:EEC

[Bas70] N. K. Basu. Error estimates for a Chebyshev quadrature method. Mathematics of Computation, 24(112):863–867, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bass:1972:RTA

[Bas72] Ronald Bass. A rank two algorithm for unconstrained minimization. Mathematics of Computation, 26(117):129–143, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bauer:1977:AMC

[Bau77] F. L. Bauer. Angstl’s mechanism for checking well-formedness of parenthesis-free formulae. Mathematics of Computation, 31 (137):318–320, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bayliss:1978:DSS

[Bay78] Alvin Bayliss. A double shooting scheme for certain unstable and singular boundary value problems. Mathematics of Computation, 32(141):61–71, January 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blackburn:1974:NCM

[BB74] J. A. Blackburn and Y. Beaudoin. A note on Chambers’ method. Mathematics of Computation, 28(126):573–574, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1977:SSG

[BBT77] Garth A. Baker, James H. Bramble, and Vidar Thomee. Single step Galerkin approximations for parabolic problems. Mathematics of Computation, 31(140): 818–847, October 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buzbee:1973:NCP

[BC73] B. L. Buzbee and Alfred Carasso. On the numerical computation of parabolic problems for preceding times. Mathematics of Computation, 27(122):237–266, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Breig:1974:NCG

[BC74] W. F. Breig and A. L. Crosbie. Numerical computation of a generalized exponential integral function. Mathematics of Computation, 28(126):575–579, April 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babb:1979:OEC

[BC79] Stanley E. Babb, Jr. and James W. Cafky.

Operational evaluation of certain infinite Bessel function integrals. Mathematics of Computation, 33(147):1033–1039, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blackburn:1973:CCG

[BCH73] John E. Blackburn, Henry H. Crapo, and Denis A. Higgs. A catalogue of combinatorial geometries. Mathematics of Computation, 27(121):155–166, January 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrow:1978:UBM

[BCSW78] D. L. Barrow, C. K. Chui, P. W. Smith, and J. D. Ward. Unicity of best mean approximation by second order splines with variable knots. Mathematics of Computation, 32(144):1131–1143, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boggs:1976:SAP

[BD76] Paul T. Boggs and J. E. Dennis, Jr. A stability analysis for perturbed nonlinear iterative methods. Mathematics of Computation, 30(134):199–215, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0395209-4.

Barrodale:1978:LCA

[BDM78] I. Barrodale, L. M. Delves, and J. C. Mason. Linear Chebyshev approximation of complex-valued functions. Mathematics of Computation, 32(143):853–863, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1974:GNP

[BDS74] H. H. Baker, A. K. Dewdney, and A. L. Szilard. Generating the nine-point graphs. Mathematics of Computation, 28(127): 833–838, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1972:PMD

[BDT72] James H. Bramble, Todd Dupont, and Vidar Thomee. Projection methods for Dirichlet’s problem in approximating polygonal domains with boundary-value corrections. Mathematics of Computation, 26(120):869–879, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Benkoski:1974:WPN

[BE74] S. J. Benkoski and P. Erdős. On weird and pseudoperfect numbers. Mathematics of Computation, 28(126):617–623, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bleicher:1975:NDS

[BE75] M. N. Bleicher and P. Erdős. The number of distinct subsums of ∑_{1}^{N}l/i. Mathematics of Computation, 29(129):29–42, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beard:1974:C

[Bea74] Jacob T. B. Beard, Jr. Computing in GF(q). Mathematics of Computation, 28 (128):1159–1166, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blair:1976:RCA

[BEJ76] J. M. Blair, C. A. Edwards, and J. H.

Johnson. Rational Chebyshev approximations for the inverse of the error function. Mathematics of Computation, 30 (136):827–830, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blair:1978:ARC

[BEJ78a] J. M. Blair, C. A. Edwards, and J. H.

Johnson. Addendum: “Rational Chebyshev approximations for the Bickley functions Ki_{n}(x)” (math. comp. 32 (1978), no. 143, 876–886). Mathematics of Computation, 32(143):Microfiche supplement, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [BEJ78b].

Blair:1978:RCA

[BEJ78b] J. M. Blair, C. A. Edwards, and J. H.

Johnson. Rational Chebyshev approximations for the Bickley functions Ki_{n}(x). Mathematics of Computation, 32(143): 876–886, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See addendum [BEJ78a].

Belford:1972:UAV

[Bel72] Geneva G. Belford. Uniform approximation of vector-valued functions with a constraint. Mathematics of Computation, 26(118):487–492, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Belford:1974:VVA

[Bel74] Geneva G. Belford. Vector-valued approximation and its application to fitting exponential decay curves. Mathematics of Computation, 28(125):179–184, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berlekamp:1970:FPL

[Ber70] E. R. Berlekamp. Factoring polynomials over large finite fields. Mathematics of Computation, 24(111):713–735, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berlekamp:1975:DSS

[Ber75] Elwyn R. Berlekamp. The design of slowly shrinking labelled squares. Mathematics of Computation, 29(129):25–27, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beyn:1979:EOC

[Bey79] Wolf-Jürgen Beyn. The exact order of convergence for finite difference approximations to ordinary boundary value problems. Mathematics of Computation, 33 (148):1213–1228, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1977:EET

[BF77] Alan E. Berger and Richard S. Falk. An error estimate for the truncation method for the solution of parabolic obstacle variational inequalities. Mathematics of Computation, 31(139):619–628, July 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berkowitz:1970:CMH

[BG70] S. Berkowitz and F. J. Garner. The calculation of multidimensional Hermite polynomials and Gram–Charlier coefficients. Mathematics of Computation, 24 (111):537–545, July 1970. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bjorck:1973:NMC

[BG73] Åke Björck and Gene H. Golub. Numerical methods for computing angles between linear subspaces. Mathematics of Computation, 27(123):579–594, July 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnhill:1975:PIB

[BG75] R. E. Barnhill and J. A. Gregory. Polynomial interpolation to boundary data on triangles. Mathematics of Computation, 29 (131):726–735, July 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brayton:1970:SRS

[BGW70] Robert K. Brayton, Fred G. Gustavson, and Ralph A. Willoughby. Some results on sparse matrices. Mathematics of Computation, 24(112):937–954, October 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunch:1974:TFI

[BH74] James R. Bunch and John E. Hopcroft.

Triangular factorization and inversion by fast matrix multiplication. Mathematics of Computation, 28(125):231–236, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berman:1977:OAP

[BH77] Joel Berman and Kit Hanes. Optimizing the arrangement of points on the unit sphere. Mathematics of Computation, 31 (140):1006–1008, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bays:1978:DFR

[BH78a] Carter Bays and Richard H. Hudson. Details of the first region of integers x with π_{3,2}(x) < π_{3,1}(x). Mathematics of Computation, 32(142):571–576, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bays:1978:FLP

[BH78b] Carter Bays and Richard H. Hudson. On the fluctuations of Littlewood for primes of the form 4n±1. Mathematics of Computation, 32(141):281–286, January 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brezzi:1977:AMF

[BJM77] F. Brezzi, C. Johnson, and B. Mercier. Analysis of a mixed finite element method for elasto-plastic plates. Mathematics of Computation, 31(140):809–817, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Belford:1974:AAT

[BK74] Geneva G. Belford and E. H. Kaufman, Jr. An application of approximation theory to an error estimate in linear algebra. Mathematics of Computation, 28(127):711–712, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunch:1977:SSM

[BK77] James R. Bunch and Linda Kaufman. Some stable methods for calculating inertia and solving symmetric linear systems. Mathematics of Computation, 31 (137):163–179, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Benton:1978:CZT

[BK78] T. C. Benton and H. D. Knoble. Common zeros of two Bessel functions. Mathematics of Computation, 32(142):533–535, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Browning:1973:MR

[BKO73] Gerald Browning, Heinz-Otto Kreiss, and Joseph Oliger. Mesh refinement. Mathematics of Computation, 27(121):29–39, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bellman:1972:UA

[BKV72] Richard Bellman, B. G. Kashef, and R. Vasudevan. A useful approximation to e^{−t^{2}}. Mathematics of Computation, 26 (117):233–235, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blair:1974:RCA

[Bla74] J. M. Blair. Rational Chebyshev approximations for the modified Bessel functions I_{0}(x) and I_{1}(x). Mathematics of Computation, 28(126):581–583, April 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blair:1976:HOA

[Bla76a] J. J. Blair. Higher order approximations to the boundary conditions for the finite element method. Mathematics of Computation, 30(134):250–262, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0398123-3.

Blass:1976:NDE

[Bla76b] Josef Blass. A note on Diophantine equation Y^{2} + k = X^{5}. Mathematics of Computation, 30(135):638–640, July 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bratley:1970:ANT

[BLM70] Paul Bratley, Fred Lunnon, and John McKay. Amicable numbers and their distribution. Mathematics of Computation, 24(110):431–432, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1975:NPC

[BLS75] John Brillhart, D. H. Lehmer, and J. L. Selfridge. New primality criteria and factorizations of 2^{m} ± 1. Mathematics of Computation, 29(130):620–647, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blue:1979:LPI

[Blu79] James L. Blue. A Legendre polynomial integral. Mathematics of Computation, 33(146):739–741, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrodale:1970:TSA

[BM70] I. Barrodale and J. C. Mason. Two simple algorithms for discrete rational approximation. Mathematics of Computation, 24(112):877–891, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beyer:1970:GST

[BMN70a] W. A. Beyer, N. Metropolis, and J. R. Neergaard. The generalized serial test applied to expansions of some irrational square roots in various bases. Mathematics of Computation, 24(111):745–747, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beyer:1970:SSD

[BMN70b] W. A. Beyer, N. Metropolis, and J. R. Neergaard. Statistical study of digits of some square roots of integers in various bases. Mathematics of Computation, 24(110):455–473, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://links.jstor.org/sici?
sici=0025-5718%28197004%2924%3A110%
3C455%3ASSODOS%3E2.0.CO%3B2-I. See

corrigendum [Sha71b].

Borosh:1975:ECF

[BMP75] I. Borosh, C. J. Moreno, and H. Porta. Elliptic curves over finite fields. II. Mathematics of Computation, 29(131):951–964, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnhill:1970:EAN

[BN70] Robert E. Barnhill and Gregory M. Nielson. An error analysis for numerical multiple integration. III. Mathematics of Computation, 24(110):301–314, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1975:MNI

[BNS75] James H. Bramble, Joachim A. Nitsche, and Alfred H. Schatz. Maximum-norm interior estimates for Ritz–Galerkin methods. Mathematics of Computation, 29 (131):677–688, July 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1973:IGI

[BNZ73] H. Brown, J. Neubuser, and H. Zassenhaus. On integral groups. III: Normalizer. Mathematics of Computation, 27 (121):167–182, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1973:RCE

[BO73] J. H. Bramble and J. E. Osborn. Rate of convergence estimates for nonselfadjoint eigenvalue approximations. Mathematics of Computation, 27(123):525–549, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babuska:1978:NTE

[BO78] I. Babuška and J. E. Osborn. Numerical treatment of eigenvalue problems for differential equations with discontinuous coefficients. Mathematics of Computation, 32(144):991–1023, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boas:1977:GPS

[Boa77] R. P. Boas, Jr. Growth of partial sums of divergent series. Mathematics of Computation, 31(137):257–264, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boggs:1974:NAC

[Bog74] Paul T. Boggs. A new algorithm for the Chebyshev solution of overdetermined linear systems. Mathematics of Computation, 28(125):203–217, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boggs:1976:CBI

[Bog76] Paul T. Boggs. The convergence of the Ben-Israel iteration for nonlinear least squares problems. Mathematics of Computation, 30(135):512–522, July 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bondesson:1971:IPE

[Bon71] Magnus Bondesson. An interior a priori estimate for parabolic difference operators and an application. Mathematics of Computation, 25(113):43–58, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borho:1972:TIK

[Bor72a] Walter Borho. On Thabit ibn Kurrah’s formula for amicable numbers. Mathematics of Computation, 26(118):571–578, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borning:1972:SR

[Bor72b] Alan Borning. Some results for k! ± 1 and 2 · 3 · 5 ⋯ p ± 1. Mathematics of Computation, 26(118):567–570, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bownds:1976:SWS

[Bow76] John M. Bownds. On solving weakly singular Volterra equations of the first kind with Galerkin approximations. Mathematics of Computation, 30(136):747–757, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1970:LBD

[Boy70] David W. Boyd. Lower bounds for the disk packing constant. Mathematics of Computation, 24(111):697–704, July 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1973:AGS

[Boy73] David W. Boyd. An algorithm for generating the sphere coordinates in a threedimensional osculatory packing. Mathematics of Computation, 27(122):369–377, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1978:PSN

[Boy78] David W. Boyd. Pisot and Salem numbers in intervals of the real line. Mathematics of Computation, 32(144):1244–1260, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bjorck:1970:SVS

[BP70] Ake Björck and Victor Pereyra. Solution of Vandermonde systems of equations. Mathematics of Computation, 24(112):893–903, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Balazs:1979:TWF

[BPD79] N. L. Balazs, H. C. Pauli, and O. B. Dabbousi. Tables of Weyl fractional integrals for the Airy function. Mathematics of Computation, 33(145):353–358, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bateman:1975:SNR

[BPW75] Paul T. Bateman, George B. Purdy, and Samuel S. Wagstaff, Jr. Some numerical results on Fekete polynomials. Mathematics of Computation, 29(129):7–23, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bauer:1972:BFD

[BR72] Louis Bauer and Edward L. Reiss. Block five diagonal matrices and the fast numerical solution of the biharmonic equation. Mathematics of Computation, 26(118): 311–326, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brezinski:1974:SSE

[BR74] Claude Brezinski and Alain C. Rieu. The solution of systems of equations using the epsilon — algorithm, and an application to boundary-value problems. Mathematics of Computation, 28(127):731–741, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Best:1976:CAC

[BR76] Michael J. Best and Klaus Ritter. A class of accelerated conjugate direction methods for linearly constrained minimization problems. Mathematics of Computation, 30(135):478–504, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babuska:1979:AOF

[BR79] I. Babuška and W. C. Rheinboldt. Analysis of optimal finite-element meshes in R^{1}. Mathematics of Computation, 33(146): 435–463, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bradley:1971:AHS

[Bra71] Gordon H. Bradley. Algorithms for Hermite and Smith normal matrices and linear Diophantine equations. Mathematics of Computation, 25(116):897–907, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Braess:1973:CAE

[Bra73a] Dietrich Braess. Chebyshev approximation by exponentials on finite subsets. Mathematics of Computation, 27(122):327–331, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brandt:1973:GLM

[Bra73b] Achi Brandt. Generalized local maximum principles for finite-difference operators. Mathematics of Computation, 27 (124):685–718, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brandt:1977:MLA

[Bra77] Achi Brandt. Multi-level adaptive solutions to boundary-value problems. Mathematics of Computation, 31(138):333–390, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1973:FOL

[Bre73] Richard P. Brent. The first occurrence of large gaps between successive primes. Mathematics of Computation, 27(124): 959–963, October 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1974:DSG

[Bre74] Richard P. Brent. The distribution of small gaps between successive primes. Mathematics of Computation, 28(125):315–324, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bremner:1975:EM

[Bre75a] Andrew Bremner. An equation of Mordell. Mathematics of Computation, 29(131): 925–928, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1975:IDP

[Bre75b] Richard P. Brent. Irregularities in the distribution of primes and twin primes. Mathematics of Computation, 29(129):43– 56, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1975:TEA

[Bre75c] Richard P. Brent. Table errata: Algorithms for minimization without derivatives (Prentice-Hall, Englewood Cliffs, N.J., 1973). Mathematics of Computation, 29(132):1166, October 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1975-29-132/
S0025-5718-1975-0371062-9;   http:
//www.ams.org/journals/mcom/1975-
29-132/S0025-5718-1975-0371062-9/
S0025-5718-1975-0371062-9.pdf; https:
//www.ams.org/mathscinet-getitem?
mr=371062;   https://www.ams.org/
mathscinet/search/authors.html?authorName=
Brent%2C%20Richard%20P..

Brent:1977:CRC

[Bre77] Richard P. Brent. Computation of the regular continued fraction for Euler’s constant. Mathematics of Computation, 31 (139):771–777, July 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1979:ZRZ

[Bre79] Richard P. Brent. On the zeros of the Riemann zeta function in the critical strip. Mathematics of Computation, 33 (148):1361–1372, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1972:NRP

[Bri72] John Brillhart. Note on representing a prime as a sum of two squares. Mathematics of Computation, 26(120):1011–1013, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Broyden:1970:CSR

[Bro70] C. G. Broyden. The convergence of singlerank quasi-Newton methods. Mathematics of Computation, 24(110):365–382, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Broyden:1971:CAS

[Bro71] C. G. Broyden. The convergence of an algorithm for solving sparse nonlinear systems. Mathematics of Computation, 25(114):285–294, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0297122-5.

Brodlie:1975:BMS

[Bro75] Kenneth W. Brodlie. On Bairstow’s method for the solution of polynomial equations. Mathematics of Computation, 29(131):816–826, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1979:SSG

[Bro79] R. Leonard Brown. Stability of sequences generated by nonlinear differential systems. Mathematics of Computation, 33(146):637–645, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brudno:1970:GIM

[Bru70a] Simcha Brudno. On generating infinitely many solutions of the Diophantine equation A^{6} + B^{6} + C^{6} = D^{6} + E^{6} + F^{6}. Mathematics of Computation, 24(110):453–454, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1970:MSS

[Bru70b] H. Brunner. Marginal stability and stabilization in the numerical integration of ordinary differential equations. Mathematics of Computation, 24(111):635–646, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1972:NMO

[Bru72] H. Brunner. A note on modified optimal linear multistep methods. Mathematics of Computation, 26(119):625–631, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1974:GSG

[Bru74a] H. Brunner. Global solution of the generalized Abel integral equation by implicit interpolation. Mathematics of Computation, 28(125):61–67, January 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1974:RCN

[Bru74b] H. Brunner. Recursive collocation for the numerical solution of stiff ordinary differential equations. Mathematics of Computation, 28(126):475–481, April 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brudno:1976:TSP

[Bru76] Simcha Brudno. Triples of sixth powers with equal sums. Mathematics of Computation, 30(135):646–648, July 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1977:DVI

[Bru77] Hermann Brunner. Discretization of Volterra integral equations of the first kind. Mathematics of Computation, 31 (139):708–716, July 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beyer:1971:LSM

[BRW71] W. A. Beyer, R. B. Roof, and Dorothy Williamson. The lattice structure of multiplicative congruential pseudorandom vectors. Mathematics of Computation, 25(114):345–363, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/journals/
mcom/1971-25-114/S0025-5718-1971-
0309263-4;   http://www.jstor.org/
stable/pdfplus/2004931.pdf. See cor-

rigenda [BW96].

Bramble:1971:LSM

[BS71] J. H. Bramble and A. H. Schatz. Least squares methods for 2m th order elliptic boundary-value problems. Mathematics of Computation, 25(113):1–32, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1977:HOL

[BS77] J. H. Bramble and A. H. Schatz. Higher order local accuracy by averaging in the finite element method. Mathematics of Computation, 31(137):94–111, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1978:SAS

[BS78] James H. Bramble and Ridgway Scott. Simultaneous approximation in scales of Banach spaces. Mathematics of Computation, 32(144):947–954, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brenner:1971:CDS

[BT71] Philip Brenner and Vidar Thomee. Contractive difference schemes for symmetric hyperbolic systems. Mathematics of Computation, 25(114):205– 217, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0297151-1.

Bramble:1972:SLS

[BT72] James H. Bramble and Vidar Thomee. Semidiscrete-least squares methods for a parabolic boundary value problem. Mathematics of Computation, 26(119):633–648, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buell:1976:CGQ

[Bue76] Duncan A. Buell. Class groups of quadratic fields. Mathematics of Computation, 30(135):610–623, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buell:1977:SCN

[Bue77] Duncan A. Buell. Small class numbers and extreme values of L-functions of quadratic fields. Mathematics of Computation, 31 (139):786–796, July 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bui:1979:NRP

[Bui79] T. D. Bui. A note on the Rosenbrock procedure. Mathematics of Computation, 33 (147):971–975, July 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burbea:1970:PCM

[Bur70] J. Burbea. A procedure for conformal maps of simply connected domains by using the Bergman function. Mathematics of Computation, 24(112):821–829, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burbea:1971:NDM

[Bur71a] J. Burbea. A numerical determination of the modulus of doubly connected domains by using the Bergman curvature. Mathematics of Computation, 25(116):743–756, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1971:PTB

[Bur71b] F. D. Burgoyne. Practical throw-back interpolation. Mathematics of Computation, 25(113):115–117, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burr:1975:DPE

[Bur75] Stefan A. Burr. On detecting a periodic event by means of periodic observations. I. Mathematics of Computation, 29(129): 57–65, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burns:1978:NCS

[Bur78] Anne M. Burns. A necessary condition for the stability of a difference approximation to a hyperbolic system of partial differential equations. Mathematics of Computation, 32(143):707–724, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buschman:1974:FSR

[Bus74] R. G. Buschman. Finite sum representations for partial derivatives of special functions with respect to parameters. Mathematics of Computation, 28(127):817–824, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buschman:1976:IHF

[Bus76] R. G. Buschman. Inequalities for hypergeometric functions. Mathematics of Computation, 30(134):303–305, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1976-30-134/
                                              S0025-5718-1976-0390306-1.

Butcher:1972:ATI

[But72a] J. C. Butcher. An algebraic theory of integration methods. Mathematics of Computation, 26(117):79–106, January 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1972:CCC

[But72b] J. C. Butcher. A convergence criterion for a class of integration methods. Mathematics of Computation, 26(117):107–117, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1973:ONM

[But73] J. C. Butcher. The order of numerical methods for ordinary differential equations. Mathematics of Computation, 27 (124):793–806, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunting:1978:DAP

[BvT78] Paul W. Bunting, Jan van Leeuwen, and Dov Tamari. Deciding associativity for partial multiplication tables of order 3. Mathematics of Computation, 32(142): 593–605, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beard:1974:FT

[BW74a] Jacob T. B. Beard, Jr. and Karen I.

West. Factorization tables for x^{n} − 1 over GF(q). Mathematics of Computation, 28 (128):1167–1168, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beard:1974:SPP

[BW74b] Jacob T. B. Beard, Jr. and Karen I. West.

Some primitive polynomials of the third kind. Mathematics of Computation, 28 (128):1166–1167, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beyer:1974:EAC

[BW74c] W. A. Beyer and M. S. Waterman. Error analysis of a computation of Euler’s constant. Mathematics of Computation, 28(126):599–604, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beard:1976:FTT

[BW76] Jacob T. B. Beard, Jr. and Karen I.

West. Factorization tables for trinomials over GF(q). Mathematics of Computation, 30(133):179–183, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beyer:1996:CLS

[BW96] W. A. Beyer and W. W. Wood. Corrigenda: “The lattice structure of multiplicative congruential pseudo-random vectors” [Math. Comp. 25 (1971), 345– 363, MR 46 #8373] by W. A. Beyer, R. B. Roof, and Dorothy Williamson. Mathematics of Computation, 65(213): 445–446, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://www.jstor.
org/stable/pdfplus/2153872.pdf. See

[BRW71].

Barrucand:1976:CTD

[BWB76] Pierre Barrucand, H. C. Williams, and L. Baniuk. A computational technique for determining the class number of a pure cubic field. Mathematics of Computation, 30(134):312–323, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0392913-9.

Buell:1977:IBB

[BWW77] D. A. Buell, H. C. Williams, and K. S. Williams. On the imaginary bicyclic biquadratic fields with class-number 2. Mathematics of Computation, 31(140): 1034–1042, October 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1970:TEF

[BZ70] James H. Bramble and Milos̆ Zlámal. Triangular elements in the finite element method. Mathematics of Computation, 24(112):809–820, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chakravarthy:1979:NCM

[CA79] Sukumar Chakravarthy and Dale Anderson. Numerical conformal mapping. Mathematics of Computation, 33(147):953–969, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cadwell:1971:LIB

[Cad71] J. H. Cadwell. Large intervals between consecutive primes. Mathematics of Computation, 25(116):909–913, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carasso:1970:PBN

[Car70] Alfred Carasso. A posteriori bounds in the numerical solution of mildly nonlinear parabolic equations. Mathematics of Computation, 24(112):785–792, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1972:ACL

[Car72] B. C. Carlson. An algorithm for computing logarithms and arctangents. Mathematics of Computation, 26(118):543–549, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carrano:1973:MBM

[Car73] F. M. Carrano. A modified Bairstow method for multiple zeros of a polynomial. Mathematics of Computation, 27 (124):781–792, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carasso:1974:LSP

[Car74] Alfred Carasso. A least squares procedure for the wave equation. Mathematics of Computation, 28(127):757–767, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carasso:1975:CSH

[Car75a] Alfred Carasso. Coupled sound and heat flow and the method of least squares.

Mathematics of Computation, 29(130): 447–463, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carta:1975:LOA

[Car75b] David G. Carta. Low-order approximations for the normal probability integral and the error function. Mathematics of Computation, 29(131):856–862, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cash:1978:EOM

[Cas78] J. R. Cash. An extension of Olver’s method for the numerical solution of linear recurrence relations. Mathematics of Computation, 32(142):497–510, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cato:1977:OKS

[Cat77] Craig Cato. The orders of the known simple groups as far as one trillion. Mathematics of Computation, 31(138):574–577, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cheema:1972:NIC

[CC72] M. S. Cheema and W. E. Conway. Numerical investigation of certain asymptotic results in the theory of partitions. Mathematics of Computation, 26(120):999–1005, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cannon:1973:IAT

[CDHW73] John J. Cannon, Lucien A. Dimino, George Havas, and Jane M. Watson. Implementation and analysis of the Todd– Coxeter algorithm. Mathematics of Computation, 27(123):463–490, July 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Choong:1971:RA

[CDR71] K. Y. Choong, D. E. Daykin, and C. R.

Rathbone. Rational approximations to π. Mathematics of Computation, 25(114): 387–392, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0300981-0. See errata

[Sha76a].

Castleton:1973:FOM

[CG73] R. N. Castleton and L. J. Grimm. A first order method for differential equations of neutral type. Mathematics of Computation, 27(123):571–577, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cantor:1972:CFA

[CGZ72] David G. Cantor, Paul H. Galyean, and Horst G. Zimmer. A continued fraction algorithm for real algebraic numbers. Mathematics of Computation, 26(119):785–791, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cochran:1970:NTF

[CH70a] James Alan Cochran and Judith N. Hoffspiegel. Numerical techniques for finding ν-zeros of Hankel functions. Mathematics of Computation, 24(110):413–422, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1970:CAC

[CH70b] W. J. Cody and K. E. Hillstrom. Chebyshev approximations for the Coulomb phase shift. Mathematics of Computation, 24(111):671–677, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Collins:1974:MRS

[CH74] George E. Collins and Ellis Horowitz. The minimum root separation of a polynomial. Mathematics of Computation, 28(126):589–597, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chin:1978:DAD

[CH78] R. C. Y. Chin and G. W. Hedstrom. A dispersion analysis for difference schemes: tables of generalized Airy functions. Mathematics of Computation, 32(144):1163– 1170, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chambers:1971:QFF

[Cha71] Ll. G. Chambers. A quadratic formula for finding the root of an equation. Mathematics of Computation, 25(114): 305–307, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0295559-1.

Chamayou:1973:PAD

[Cha73] Jean-Marie-François Chamayou. A probabilistic approach to a differential-difference equation arising in analytic number theory. Mathematics of Computation, 27 (121):197–203, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chatterjee:1974:NIP

[Cha74] Gurudas Chatterjee. Negative integral powers of a bidiagonal matrix. Mathematics of Computation, 28(127):713–714, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Charalambides:1979:BRP

[Cha79] Ch. A. Charalambides. Bernoulli related polynomials and numbers. Mathematics of Computation, 33(146):794–804, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chipman:1972:NCT

[Chi72] D. M. Chipman. The numerical computation of two transcendental functions related to the exponential integral. Mathematics of Computation, 26(117):241–249, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chisholm:1973:RAD

[Chi73] J. S. R. Chisholm. Rational approximants defined from double power series. Mathematics of Computation, 27(124):841–848, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chin:1979:SGM

[CHK79] R. C. Y. Chin, G. W. Hedstrom, and K. E. Karlsson. A simplified Galerkin method for hyperbolic equations. Mathematics of Computation, 33(146):647–658, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Choi:1971:CSI

[Cho71] S. L. G. Choi. Covering the set of integers by congruence classes of distinct moduli. Mathematics of Computation, 25(116): 885–895, October 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chorin:1973:AEW

[Cho73] Alexandre Joel Chorin. Accurate evaluation of Wiener integrals. Mathematics of Computation, 27(121):1–15, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [CS73].

Cody:1971:CAR

[CHT71] W. J. Cody, K. E. Hillstrom, and Henry C. Thacher, Jr. Chebyshev approximations for the Riemann zeta function. Mathematics of Computation, 25(115):537–547, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ciarlet:1978:IEE

[Cia78] Philippe G. Ciarlet. Interpolation error estimates for the reduced Hsieh–Clough– Tocher triangle. Mathematics of Computation, 32(142):335–344, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ciment:1971:SDS

[Cim71] Melvyn Ciment. Stable difference schemes with uneven mesh spacings. Mathematics of Computation, 25(114):219– 227, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0300470-3.

Cioffari:1979:ECP

[Cio79] Vincent G. Cioffari. The Euclidean condition in pure cubic and complex quartic fields. Mathematics of Computation, 33(145):389–398, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Conde:1979:Z

[CK79] S. Conde and S. L. Kalla. The ν-zeros of J_{−ν}(x). Mathematics of Computation, 33(145):423–426, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ciment:1975:HOC

[CL75] Melvyn Ciment and Stephen H. Leventhal.

Higher order compact implicit schemes for the wave equation. Mathematics of Computation, 29(132):985–994, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ciment:1978:NOC

[CL78] Melvyn Ciment and Stephen H. Leventhal.

A note on the operator compact implicit method for the wave equation. Mathematics of Computation, 32(141):143–147, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cline:1972:RCL

[Cli72] A. K. Cline. Rate of convergence of Lawson’s algorithm. Mathematics of Computation, 26(117):167–176, January 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coleman:1978:VM

[CM78] John P. Coleman and Julie Mohamed.

On de Vogelaere’s method for y′′ = f(x, y). Mathematics of Computation, 32 (143):751–762, July 1978. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chow:1978:FZM

[CMPY78] Shui Nee Chow, John Mallet-Paret, and James A. Yorke. Finding zeroes of maps: homotopy methods that are constructive with probability one. Mathematics of Computation, 32(143):887–899, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1970:ASN

[Coh70] Henri Cohen. On amicable and sociable numbers. Mathematics of Computation, 24(110):423–429, April 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1973:PAF

[Coh73] Edgar A. Cohen, Jr. Polynomial approximation of a function and its first derivative in near minimax norms. Mathematics of Computation, 27(124):817–827, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1976:EMW

[Coh76a] M. E. Cohen. Extensions of the Mehler– Weisner and other results for the Hermite function. Mathematics of Computation, 30 (135):553–564, July 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1976:DP

[Coh76b] Harvey Cohn. Dyadotropic polynomials. Mathematics of Computation, 30(136): 854–862, October 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1977:SCG

[Coh77] M. E. Cohen. Some classes of generating functions for the Laguerre and Hermite polynomials. Mathematics of Computation, 31(138):511–518, April 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1979:CSC

[Coh79a] Harvey Cohn. Cyclic-sixteen class fields for Q(−p)^{1/2} by modular arithmetic. Mathematics of Computation, 33(148):1307– 1316, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1979:DPI

[Coh79b] Harvey Cohn. Dyadotropic polynomials. II. Mathematics of Computation, 33(145):

359–367, January 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cox:1972:NCM

[Cox72] M. G. Cox. A note on Chambers’ method for finding a zero of a function. Mathematics of Computation, 26(119):749–750, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carasso:1970:ABV

[CP70] Alfred Carasso and Seymour V. Parter. An analysis of “boundary-value techniques” for parabolic problems. Mathematics of Computation, 24(110):315–340, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crandall:1979:SLT

[CP79] R. E. Crandall and M. A. Penk. A search for large twin prime pairs. Mathematics of Computation, 33(145):383–388, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1970:CAD

[CPT70] W. J. Cody, Kathleen A. Paciorek, and Henry C. Thacher, Jr. Chebyshev approximations for Dawson’s integral. Mathematics of Computation, 24(109):171–178, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Curtis:1972:GIC

[CR72] A. R. Curtis and P. Rabinowitz. On the Gaussian integration of Chebyshev polynomials. Mathematics of Computation, 26(117):207–211, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Christiansen:1978:EAS

[CR78] J. Christiansen and R. D. Russell. Error analysis for spline collocation methods with application to knot selection. Mathematics of Computation, 32(142):415–419, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crandall:1978:P

[Cra78] R. E. Crandall. On the “3x + 1” problem. Mathematics of Computation, 32 (144):1281–1292, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cryer:1971:MCS

[Cry71] Colin W. Cryer. The method of christopherson for solving free boundary problems for infinite journal bearings by means of finite differences. Mathematics of Computation, 25(115):435–443, July 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chartres:1971:CDM

[CS71] Bruce Chartres and Robert Stepleman. Convergence of difference methods for initial and boundary value problems with discontinuous data. Mathematics of Computation, 25(116):729–732, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chorin:1973:CAE

[CS73] Alexandre J. Chorin and Daniel Shanks. Corrigenda: Accurate Evaluation of Wiener Integrals and A Note on the Catalan–Dickson Conjecture. Mathematics of Computation, 27(124):1011–1012, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlitz:1975:GFT

[CS75a] L. Carlitz and Richard Scoville. A generating function for triangular partitions. Mathematics of Computation, 29(129):67– 77, January 1975. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1975:ENS

[CS75b] Fred Cohen and J. L. Selfridge. Not every number is the sum or difference of two prime powers. Mathematics of Computation, 29(129):79–81, January 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooper:1979:SSR

[CS79] G. J. Cooper and A. Sayfy. Semiexplicit A-stable Runge–Kutta methods. Mathematics of Computation, 33(146):541–556, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1973:CAP

[CST73] W. J. Cody, Anthony J. Strecok, and Henry C. Thacher, Jr. Chebyshev approximations for the psi function. Mathematics of Computation, 27(121):123–127, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1977:SBA

[CSW77] Charles K. Chui, Philip W. Smith, and Joseph D. Ward. On the smoothness of best L_{2} approximants from nonlinear spline manifolds. Mathematics of Computation, 31(137):17–23, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cullum:1979:ECS

[Cul79] Jane Cullum. The effective choice of the smoothing norm in regularization. Mathematics of Computation, 33(145):149–170, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cusick:1971:DAT

[Cus71] T. W. Cusick. Diophantine approximation of ternary linear forms. Mathematics of Computation, 25(113):163–180, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cusick:1972:DAT

[Cus72] T. W. Cusick. Diophantine approximation of ternary linear forms. II. Mathematics of Computation, 26(120):977–993, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cusick:1977:SMC

[Cus77] T. W. Cusick. The Szekeres multidimensional continued fraction. Mathematics of Computation, 31(137):280–317, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Danloy:1973:NCG

[Dan73] Bernard Danloy. Numerical construction of Gaussian quadrature formulas for ∫_{0}^{1}(−Logx) · x^{α} · f(x) · dx and ∫_{0}^{∞} E_{m}(x) · f(x) · dx. Mathematics of Computation, 27(124):861–869, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Detrich:1979:FSE

[DC79] John Detrich and Robert W. Conn. Finite sum evaluation of the Gauss hypergeometric function in an important special case. Mathematics of Computation, 33(146):788–791, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1973:FEC

[DD73] Jim Douglas, Jr. and Todd Dupont. A finite element collocation method for quasilinear parabolic equations. Mathematics of Computation, 27(121):17–28, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1974:SNS

[dD74] C. de Boor and James W. Daniel. Splines with nonnegative B-spline coefficients. Mathematics of Computation, 28(126): 565–568, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1975:EIC

[DD75a] Jim Douglas, Jr. and Todd Dupont. The effect of interpolating the coefficients in nonlinear parabolic Galerkin procedures. Mathematics of Computation, 29(130): 360–389, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1975:GMN

[DD75b] Jim Douglas, Jr. and Todd Dupont. A Galerkin method for a nonlinear Dirichlet problem. Mathematics of Computation, 29 (131):689–696, July 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1975:OEE

[DDW75] Jim Douglas, Jr., Todd Dupont, and Lars Wahlbin. Optimal L_{∞} error estimates for Galerkin approximations to solutions of two-point boundary value problems. Mathematics of Computation, 29(130):475–483, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1978:QPA

[DDW78] Jim Douglas, Jr., Todd Dupont, and Mary F. Wheeler. A quasi-projection analysis of Galerkin methods for parabolic and hyperbolic equations. Mathematics of Computation, 32(142):345–362, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1976:BNA

[de 76] Carl de Boor. A bound on the L^{∞}- norm of L^{2} — approximation by splines in terms of a global mesh ratio. Mathematics of Computation, 30(136):765–771, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deDoncker:1979:NEM

[de 79] Elise de Doncker. New Euler–Maclaurin expansions and their application to quadrature over the s-dimensional simplex. Mathematics of Computation, 33 (147):1003–1018, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demko:1979:BB

[Dem79] Stephen Demko. On bounding ∥A^{−1}∥_{∞} for banded A. Mathematics of Computation, 33(148):1283–1288, October 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dennis:1971:CBM

[Den71] J. E. Dennis, Jr. On the convergence of Broyden’s method for nonlinear systems of equations. Mathematics of Computation, 25(115):559–567, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Derr:1972:IIM

[Der72] Leroy J. Derr. Inversion of isoclinal matrices. Mathematics of Computation, 26 (119):719–721, July 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Derenzo:1977:AHC

[Der77] Stephen E. Derenzo. Approximations for hand calculators using small integer coefficients. Mathematics of Computation, 31(137):214–222, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Deshpande:1974:CEF

[Des74] V. L. Deshpande. Confluent expansions for functions of two variables. Mathematics of Computation, 28(126):605–611, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Durbin:1977:PRF

[DF77] John R. Durbin and K. Bolling Farmer. On projective representations of finite wreath products. Mathematics of Computation, 31(138):527–535, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Daniel:1976:RSA

[DGKS76] J. W. Daniel, W. B. Gragg, L. Kaufman, and G. W. Stewart. Reorthogonalization and stable algorithms for updating the Gram–Schmidt QR factorization. Mathematics of Computation, 30(136):772–795, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Diamond:1978:SRI

[Dia78a] Harold G. Diamond. Stability of rounded off inverses under iteration. Mathematics of Computation, 32(141):227–232, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiazyDiaz:1978:SFI

[Dia78b] F. Diaz y Diaz. On some families of imaginary quadratic fields. Mathematics of Computation, 32(142):637–650, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Diaz:1979:CGM

[Día79] Julio César Díaz. A collocation-Galerkin method for Poisson’s equation on rectangular regions. Mathematics of Computation, 33(145):77–84, January 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiDonato:1978:AR

[DiD78] A. R. DiDonato. An approximation for ∫_{x}^{∞} e^{−t^{2}/2}t^{p}dt, x > 0, p real. Mathematics of Computation, 32(141):271–275, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dieter:1971:PRN

[Die71] U. Dieter. Pseudo-random numbers: The exact distribution of pairs. Mathematics of Computation, 25(116):855–883, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dieter:1975:HCS

[Die75] U. Dieter. How to calculate shortest vectors in a lattice. Mathematics of Computation, 29(131):827–833, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dijkstra:1977:CFE

[Dij77] D. Dijkstra. A continued fraction expansion for a generalization of Dawson’s integral. Mathematics of Computation, 31(138):503–510, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dixon:1970:CIR

[Dix70] John D. Dixon. Computing irreducible representations of groups. Mathematics of Computation, 24(111):707–712, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dutt:1973:NAF

[DLT73] J. E. Dutt, T. K. Lin, and L. C. Tao. New approximations to familiar functions.

Mathematics of Computation, 27(124): 939–942, October 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Diaz:1970:IPE

[DM70] J. B. Diaz and F. T. Metcalf. On iteration procedures for equations of the first kind, Ax = y, and Picard’s criterion for the existence of a solution. Mathematics of Computation, 24(112):923–935, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delves:1971:CRV

[DM71] L. M. Delves and K. O. Mead. On the convergence rates of variational methods. I: Asymptotically diagonal systems. Mathematics of Computation, 25(116):699–716, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dennis:1974:CSC

[DM74] J. E. Dennis, Jr. and Jorge J. More.

A characterization of superlinear convergence and its application to quasi-Newton methods. Mathematics of Computation, 28(126):549–560, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dressler:1974:SDP

[DMP74] Robert E. Dressler, Andrzej Makowski, and Thomas Parker. Sums of distinct primes from congruence classes modulo 12. Mathematics of Computation, 28(126): 651–652, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Diaz:1974:DFO

[DO74] J. B. Diaz and T. J. Osler. Differences of fractional order. Mathematics of Computation, 28(125):185–202, January 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Donaldson:1979:MRI

[Don79] John L. Donaldson. Minkowski reduction of integral matrices. Mathematics of Computation, 33(145):201–216, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dorr:1971:EIC

[Dor71] Fred W. Dorr. An example of illconditioning in the numerical solution of singular perturbation problems. Mathematics of Computation, 25(114):271– 283, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0297142-0.

Dotson:1978:IPN

[Dot78] W. G. Dotson, Jr. An iterative process for nonlinear monotonic nonexpansive operators in Hilbert space. Mathematics of Computation, 32(141):223–225, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dougalis:1979:MGM

[Dou79] Vassilios A. Dougalis. Multistep-Galerkin methods for hyperbolic equations. Mathematics of Computation, 33(146):563–584, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dressler:1974:XXX

[DP74] Robert E. Dressler and Thomas Parker. 12, 758. Mathematics of Computation, 28 (125):313–314, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Davis:1972:NSI

[DR72] P. J. Davis and P. Rabinowitz. On the nonexistence of simplex integration rules for infinite integrals. Mathematics of Computation, 26(119):687–688, July 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Drezner:1978:CBN

[Dre78] Z. Drezner. Computation of the bivariate normal integral. Mathematics of Computation, 32(141):277–279, January 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1977:CCG

[dS77] Carl de Boor and Blair Swartz. Comments on the comparison of global methods for linear two-point boundary value problems. Mathematics of Computation, 31(140):916–921, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiazyDiaz:1979:QFR

[DSW79] F. Diaz y Diaz, Daniel Shanks, and H. C.

Williams. Quadratic fields with 3-rank equal to 4. Mathematics of Computation, 33(146):836–840, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duff:1974:NNA

[Duf74] I. S. Duff. On the number of nonzeros added when Gaussian elimination is performed on sparse random matrices. Mathematics of Computation, 28(125):219–230, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dumont:1979:CIS

[Dum79] Dominique Dumont. A combinatorial interpretation for the Schett recurrence on the Jacobian elliptic functions. Mathematics of Computation, 33(148):1293–1297, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dunham:1975:CAI

[Dun75a] Charles B. Dunham. Chebyshev approximation by interpolating rationals on [α, ∞). Mathematics of Computation, 29(130):549–551, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dunham:1975:CFH

[Dun75b] Charles B. Dunham. Convergence of the Fraser–Hart algorithm for rational Chebyshev approximation. Mathematics of Computation, 29(132):1078–1082, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dunham:1975:DWA

[Dun75c] Charles B. Dunham. A difficulty in Williams’ algorithm for interpolating rationals. Mathematics of Computation, 29(130):552–553, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dunham:1976:RAV

[Dun76] Charles B. Dunham. Rational approximation with a vanishing weight function and with a fixed value at zero. Mathematics of Computation, 30(133):45–47, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duris:1971:OQF

[Dur71] C. S. Duris. Optimal quadrature formulas using generalized inverses. part I: General theory and minimum variance formulas. Mathematics of Computation, 25(115): 495–504, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dushane:1973:CVM

[Dus73] Theodore E. Dushane. Convergence for a vortex method for solving Euler’s equation. Mathematics of Computation, 27 (124):719–728, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dutka:1971:SRD

[Dut71] Jacques Dutka. The square root of 2 to 1, 000, 000 decimals. Mathematics of Computation, 25(116):927–930, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deHoog:1973:AEP

[dW73] Frank de Hoog and Richard Weiss. Asymptotic expansions for product integration. Mathematics of Computation, 27(122): 295–306, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deHoog:1977:ALM

[dW77] Frank R. de Hoog and Richard Weiss. The application of linear multistep methods to singular initial value problems. Mathematics of Computation, 31(139):676–690, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eckhardt:1976:RAW

[Eck76] Ulrich Eckhardt. A rational approximation to Weierstrass’ ℘-function. Mathematics of Computation, 30(136):818–826, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ecklund:1974:NFA

[EES74] E. F. Ecklund, Jr., P. Erdős, and J. L.

Selfridge. A new function associated with the prime factors of \binom{n}{k}. Mathematics of Computation, 28(126):647–649, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ewing:1979:NAC

[EF79] Richard E. Ewing and Richard S. Falk. Numerical approximation of a Cauchy problem for a parabolic partial differential equation. Mathematics of Computation, 33 (148):1125–1144, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eisenbeis:1978:CRP

[EFO78] H. Eisenbeis, G. Frey, and B. Ommerborn. Computation of the 2-rank of pure cubic fields. Mathematics of Computation, 32(142):559–569, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdos:1975:PF

[EGRS75] P. Erdős, R. L. Graham, I. Z. Ruzsa, and E. G. Straus. On the prime factors of \binom{2n}{n}. Mathematics of Computation, 29(129):83– 92, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Evans:1979:CNO

[EH79] Ronald J. Evans and Jay Roderick Hill. The cyclotomic numbers of order sixteen. Mathematics of Computation, 33(146): 827–835, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ehrlich:1972:CHE

[Ehr72] L. W. Ehrlich. Coupled harmonic equations, SOR, and Chebyshev acceleration. Mathematics of Computation, 26(118): 335–343, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ehrlich:1979:MTN

[Ehr79] Louis W. Ehrlich. A marching technique for nonseparable equations. Mathematics of Computation, 33(147):881–890, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eilbeck:1971:TEH

[Eil71] J. C. Eilbeck. Table errata: Higher transcendental functions. Vol. I, II (McGraw-Hill, New York, 1953) by A. Erdélyi, W. Magnus, F. Oberhettinger and F. G. Tricomi. Mathematics of Computation, 25 (115):635, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-115/
S0025-5718-1971-0414948-5;   http:
//www.ams.org/journals/mcom/1971-
25-115/S0025-5718-1971-0414948-5/
S0025-5718-1971-0414948-5.pdf; https:
//www.ams.org/mathscinet-getitem?
mr=414948;   https://www.ams.org/
mathscinet/search/authors.html?authorName=
Eilbeck%2C%20J.%20C..

Eilon:1972:NCT

[Eil72] B. Eilon. A note concerning the twostep Lax–Wendroff method in three dimensions. Mathematics of Computation, 26(117):41–43, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1971:UAE

[Ell71] David Elliott. Uniform asymptotic expansions of the Jacobi polynomials and an associated function. Mathematics of Computation, 25(114):309–315, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0294737-5.

Engquist:1977:ABC

[EM77] Björn Engquist and Andrew Majda. Absorbing boundary conditions for the numerical simulation of waves. Mathematics of Computation, 31(139):629–651, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1978:CII

[EM78] R. Ernvall and T. Metsänkylä. Cyclotomic invariants and E-irregular primes. Mathematics of Computation, 32(142):617–629, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigenda [Ern79].

Ehle:1975:TPA

[EP75] Byron L. Ehle and Zdenek Picel. Twoparameter, arbitrary order, exponential approximations for stiff equations. Mathematics of Computation, 29(130):501–511, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1979:GTQ

[EP79] David Elliott and D. F. Paget. Gauss type quadrature rules for Cauchy principal value integrals. Mathematics of Computation, 33(145):301–309, January 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdelsky:1973:CBD

[Erd73] P. J. Erdelsky. Computing the Brouwer degree in R^{2}. Mathematics of Computation, 27(121):133–137, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdos:1976:APA

[Erd76] P. Erdős. On asymptotic properties of aliquot sequences. Mathematics of Computation, 30(135):641–645, July 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1979:CCI

[Ern79] Reijo Ernvall. Corrigenda: “Cyclotomic invariants and E-irregular primes” (Math. Comp. 32 (1978), no. 142, 617–629) by Ernvall and T. Metsänkylä. Mathematics of Computation, 33(145):433, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [EM78].

Esper:1975:TRS

[Esp75] N. Esper. Tables of reductions of symmetrized inner products (‘inner plethysms’) of ordinary irreducible representations of symmetric groups. Mathematics of Computation, 29(132):1150– 1151, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eisenhart:1973:CHP

[ET73] Laurance C. Eisenhart and George E.

Trapp, Jr. Computation of Hermite polynomials. Mathematics of Computation, 27 (123):625–632, July 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ellacott:1976:LCA

[EW76] S. Ellacott and Jack Williams. Linear Chebyshev approximation in the complex plane using Lawson’s algorithm. Mathematics of Computation, 30(133):35–44, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Falk:1974:EEA

[Fal74] Richard S. Falk. Error estimates for the approximation of a class of variational inequalities. Mathematics of Computation, 28(128):963–971, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Falk:1976:AFE

[Fal76a] Richard S. Falk. An analysis of the finite element method using Lagrange multipliers for the stationary Stokes equations. Mathematics of Computation, 30(134): 241–249, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1976-30-134/
                                              S0025-5718-1976-0403260-0.

Falk:1976:FEM

[Fal76b] Richard S. Falk. A finite element method for the stationary Stokes equations using trial functions which do not have to satisfy div υ = 0. Mathematics of Computation, 30(136):698–702, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fatunla:1978:ITP

[Fat78] Simeon Ola. Fatunla. An implicit twopoint numerical integration formula for linear and nonlinear stiff systems of ordinary differential equations. Mathematics of Computation, 32(141):1–11, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fraser:1978:MVD

[FB78] W. Fraser and J. M. Bennett. A method of virtual displacements for the degenerate discrete l_{1} approximation problem. Mathematics of Computation, 32(142):421–430, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1977:IID

[fC77] Nai fu Chen. Inverse iteration on defective matrices. Mathematics of Computation, 31 (139):726–732, July 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1973:CZE

[FCC73a] Henry E. Fettis, James C. Caslin, and Kenneth R. Cramer. Complex zeros of the error function and of the complementary error function. Mathematics of Computation, 27(122):401–407, April 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1973:SPC

[FCC73b] Henry E. Fettis, James C. Caslin, and Kenneth R. Cramer. Saddle points of the complementary error function. Mathematics of Computation, 27(122):409–412, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1970:NMC

[Fet70] Henry E. Fettis. A new method for computing toroidal harmonics. Mathematics of Computation, 24(111):667–670, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1971:TEBa

[Fet71a] Henry E. Fettis. Table errata: Tables of toroidal harmonics. I: Orders 0 5, all significant degrees (Aerospace Res. Lab., Office of Aerospace Res., United States Air Force, Wright-Patterson Air Force Base, Ohio, 1969) by Fettis and James C. Caslin. Mathematics of Computation, 25(114):405–407, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0400627-7.

Fettis:1971:TEBb

[Fet71b] Henry E. Fettis. Table errata: Tables of toroidal harmonics, II: Orders 5 10, all significant degrees (Aerospace Res. Lab., Office of Aerospace Res. United States Air Force, Wright-Patterson Air Force Base, Ohio, 1969) by Fettis and James C. Caslin. Mathematics of Computation, 25(114):407–408, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0400628-9.

Fettis:1972:EAF

[Fet72] Henry E. Fettis. An explicit addition formula for the inverse Jacobian elliptic functions. Mathematics of Computation, 26 (120):965–967, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1974:SAC

[Fet74] Henry E. Fettis. A stable algorithm for computing the inverse error function in the ‘tail-end’ region. Mathematics of Computation, 28(126):585–587, April 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1976:CR

[Fet76] Henry E. Fettis. Complex roots of sin z = az, cos z = az, and cosh z = az. Mathematics of Computation, 30(135):541–545, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1979:AEU

[Fet79] Henry E. Fettis. An asymptotic expansion for the upper percentage points of the χ^{2}-distribution. Mathematics of Computation, 33(147):1059–1064, July 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fischer:1974:FTM

[FGH^{+}74] D. Fischer, G. Golub, O. Hald, C. Leiva, and O. Widlund. On Fourier–Toeplitz methods for separable elliptic problems.

Mathematics of Computation, 28(126): 349–368, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frey:1978:SRG

[FHP78] A. E. Frey, C. A. Hall, and T. A.

Porsching. Some results on the global inversion of bilinear and quadratic isoparametric finite element transformations. Mathematics of Computation, 32(143): 725–749, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frank:1977:EAI

[FHU77] Reinhard Frank, Joerg Hertling, and Christoph W. Ueberhuber. An extension of the applicability of iterated deferred corrections. Mathematics of Computation, 31(140):907–915, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fields:1975:PE

[FI75] Jerry L. Fields and Mourad E. H. Ismail. Polynomial expansions. Mathematics of Computation, 29(131):894–902, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fiala:1971:SAS

[Fia71] Frantisek Fiala. Strongly asymmetric sequences generated by four elements. Mathematics of Computation, 25(113):155–162, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fields:1972:AEM

[Fie72] Jerry L. Fields. The asymptotic expansion of the Meijer G-function. Mathematics of Computation, 26(119):757–765, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Field:1977:ESC

[Fie77] David A. Field. Estimates of the speed of convergence of continued fraction expansions of functions. Mathematics of Computation, 31(138):495–502, April 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

FitzGerald:1974:OIV

[Fit74] Carl H. FitzGerald. Optimal indexing of the vertices of graphs. Mathematics of Computation, 28(127):825–831, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fornberg:1975:CZJ

[FK75] B. Fornberg and K. S. Kolbig. Complex zeros of the Jonquiere or polylogarithm function. Mathematics of Computation, 29(130):582–599, April 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Feldstein:1971:EBC

[FM71] Alan Feldstein and Richard K. Miller. Error bounds for compound quadrature of weakly singular integrals. Mathematics of Computation, 25(115):505–520, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fischer:1978:NSG

[FM78] J. Fischer and J. McKay. The nonabelian simple groups G, |G| < 10^{6} — maximal subgroups. Mathematics of Computation, 32(144):1293–1302, October 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ferguson:1975:PFI

[FNC75] Helaman Rolfe Pratt Ferguson, Dale E. Nielsen, and Grant Cook. A partition formula for the integer coefficients of the theta function nome. Mathematics of Computation, 29(131):851–855, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Flaherty:1977:NSB

[FO77] Joseph E. Flaherty and R. E. O’Malley, Jr.

The numerical solution of boundary value problems for stiff differential equations. Mathematics of Computation, 31(137):66– 93, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Forsythe:1972:NCM

[For72] George E. Forsythe. Von Neumann’s comparison method for random sampling from the normal and other distributions. Mathematics of Computation, 26(120):817–826, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fornberg:1973:ILF

[For73] B. Fornberg. On the instability of leapfrog and Crank–Nicolson approximations of a nonlinear partial differential equation. Mathematics of Computation, 27(121):45– 57, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ford:1977:GJT

[For77] J. A. Ford. A generalization of the Jenkins–Traub method. Mathematics of Computation, 31(137):193–203, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fletcher:1974:MF

[FP74] R. Fletcher and M. J. D. Powell. On the modification of LDL^{T} factorizations. Mathematics of Computation, 28(128): 1067–1087, October 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franke:1971:OCR

[Fra71a] Richard Franke. Obtaining cubatures for rectangles and other planar regions by using orthogonal polynomials. Mathematics of Computation, 25(116):803–817, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franklin:1971:SBS

[Fra71b] Joel N. Franklin. Stability of bounded solutions of linear functional equations. Mathematics of Computation, 25(115): 413–424, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franklin:1974:TMI

[Fra74] Joel N. Franklin. On Tikhonov’s method for ill-posed problems. Mathematics of Computation, 28(128):889–907, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frame:1979:HPS

[Fra79] J. S. Frame. The Hankel power sum matrix inverse and the Bernoulli continued fraction. Mathematics of Computation, 33(146):815–826, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fritsch:1970:ERM

[Fri70] F. N. Fritsch. On the existence of regions with minimal third degree integration formulas. Mathematics of Computation, 24(112):855–861, October 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fike:1971:MAS

[FS71] C. T. Fike and P. H. Sterbenz. Minimax approximations subject to a constraint. Mathematics of Computation, 25(114):295–298, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0303176-X.

Ford:1973:SCD

[FT73] William H. Ford and T. W. Ting. Stability and convergence of difference approximations to pseudo-parabolic partial differential equations. Mathematics of Computation, 27(124):737–743, October 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Feldstein:1977:ABV

[FT77] Alan Feldstein and J. F. Traub. Asymptotic behavior of vector recurrences with applications. Mathematics of Computation, 31(137):180–192, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1974:SAH

[GA74] M. Goldberg and S. Abarbanel. Stable approximations for hyperbolic systems with moving internal boundary conditions. Mathematics of Computation, 28(126):413–447, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gabutti:1979:HPM

[Gab79] Bruno Gabutti. On high precision methods for computing integrals involving Bessel functions. Mathematics of Computation, 33(147):1049–1057, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gaffney:1978:COI

[Gaf78] P. W. Gaffney. To compute the optimal interpolation formula. Mathematics of Computation, 32(143):763–777, July 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Galant:1971:ICT

[Gal71] David Galant. An implementation of Christoffel’s theorem in the theory of orthogonal polynomials. Mathematics of Computation, 25(113):111–113, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Galantai:1978:CNM

[Gal78] Aurel Galantai. The comparison of numerical methods for solving polynomial equations. Mathematics of Computation, 32(142):391–397, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1970:CGQ

[Gau70] Walter Gautschi. On the construction of Gaussian quadrature rules from modified moments. Mathematics of Computation, 24(110):245–260, April 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1972:COP

[Gau72] Walter Gautschi. The condition of orthogonal polynomials. Mathematics of Computation, 26(120):923–924, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1975:NCT

[Gau75] Walter Gautschi. Nonexistence of Chebyshev-type quadratures on infinite intervals. Mathematics of Computation, 29(129):93–99, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1977:ACC

[Gau77] Walter Gautschi. Anomalous convergence of a continued fraction for ratios of Kummer functions. Mathematics of Computation, 31(140):994–999, October 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1979:CPP

[Gau79a] Walter Gautschi. The condition of polynomials in power form. Mathematics of Computation, 33(145):343–352, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1979:PPL

[Gau79b] Walter Gautschi. On the preceding paper “A Legendre polynomial integral” by James L. Blue. Mathematics of Computation, 33(146):742–743, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gay:1977:MSV

[Gay77] David Gay. Modifying singular values: existence of solutions to systems of nonlinear equations having a possibly singular Jacobian matrix. Mathematics of Computation, 31(140):962–973, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigenda [Gay79].

Gay:1979:CMS

[Gay79] David M. Gay. Corrigenda: “Modifying singular values: existence of solutions to systems of nonlinear equations having a possibly singular Jacobian matrix” (math. comp. 31 (1977), no. 140, 962–973). Mathematics of Computation, 33(145):432–433, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Gay77].

Ghil:1979:FCR

[GB79] Michael Ghil and Ramesh Balgovind. A fast Cauchy–Riemann solver. Mathematics of Computation, 33(146):585–635, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gekeler:1972:SSE

[Gek72] E. Gekeler. On the solution of systems of equations by the epsilon algorithm of Wynn. Mathematics of Computation, 26(118):427–436, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gekeler:1974:EFD

[Gek74] Eckart Gekeler. On the eigenvectors of a finite-difference approximation to the Sturm–Liouville eigenvalue problem. Mathematics of Computation, 28(128): 973–979, October 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gentleman:1972:OMC

[Gen72] W. Morven Gentleman. Optimal multiplication chains for computing a power of a symbolic polynomial. Mathematics of Computation, 26(120):935–939, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gerth:1975:NCG

[Ger75] Frank Gerth, III. A note on l-class groups of number fields. Mathematics of Computation, 29(132):1135–1137, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gill:1974:MMM

[GGMS74] P. E. Gill, G. H. Golub, W. Murray, and M. A. Saunders. Methods for modifying matrix factorizations. Mathematics of Computation, 28(126):505–535, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guderley:1972:PCM

[GH72] Karl G. Guderley and Chen-Chi Hsu. A predictor-corrector method for a certain class of stiff differential equations. Mathematics of Computation, 26(117):51–69, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grosswald:1979:APC

[GH79] Emil Grosswald and Peter Hagis, Jr.

Arithmetic progressions consisting only of primes. Mathematics of Computation, 33 (148):1343–1352, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Girault:1976:NAC

[Gir76] V. Girault. Nonelliptic approximation of a class of partial differential equations with Neumann boundary condition. Mathematics of Computation, 30(133):68–91, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gentleman:1974:EDE

[GJ74] W. M. Gentleman and S. C. Johnson.

The evaluation of determinants by expansion by minors and the general problem of substitution. Mathematics of Computation, 28(126):543–548, April 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guzek:1973:NEA

[GK73] Joseph A. Guzek and Gene A. Kemper. A new error analysis for a cubic spline approximate solution of a class of Volterra integro-differential equations. Mathematics of Computation, 27(123):563–570, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gustafsson:1972:STD

[GKS72] Bertil Gustafsson, Heinz-Otto Kreiss, and Arne Sundström. Stability theory of difference approximations for mixed initial boundary value problems. II. Mathematics of Computation, 26(119):649–686, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1971:EII

[Gla71] M. L. Glasser. An elliptic integral identity. Mathematics of Computation, 25 (115):533–534, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1974:RFM

[Gla74a] M. L. Glasser. Reduction formulas for multiple series. Mathematics of Computation, 28(125):265–266, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1974:SDI

[Gla74b] M. L. Glasser. Some definite integrals of the product of two Bessel functions of the second kind: (order zero). Mathematics of Computation, 28(126):613–615, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gladwin:1979:QRM

[Gla79a] Charles J. Gladwin. Quadrature rule methods for Volterra integral equations of the first kind. Mathematics of Computation, 33(146):705–716, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1979:NI

[Gla79b] M. L. Glasser. A note on the integral ∫_{0}^{∞} t^{2α−1}(1 + t^{2})^{1−α−β}J_{ν}(x\sqrt{1 + t^{2}})dt. Mathematics of Computation, 33(146): 792–793, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gill:1975:MCM

[GMS75] Philip E. Gill, Walter. Murray, and Michael A. Saunders. Methods for computing and modifying the LDV factors of a matrix. Mathematics of Computation, 29 (132):1051–1077, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Giroire:1978:NSE

[GN78] J. Giroire and J. C. Nedelec. Numerical solution of an exterior Neumann problem using a double layer potential. Mathematics of Computation, 32(144):973–990, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Godwin:1978:NCN

[God78] H. J. Godwin. A note on congruent numbers. Mathematics of Computation, 32 (141):293–295, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigenda [God79].

Godwin:1979:CNC

[God79] H. J. Godwin. Corrigenda: “A note on congruent numbers” (math. comp. 32 (1978), no. 141, 293–295). Mathematics of Computation, 33(146):847, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [God78].

Goldfarb:1970:FVM

[Gol70] Donald Goldfarb. A family of variablemetric methods derived by variational means. Mathematics of Computation, 24(109):23–26, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldfarb:1972:MMI

[Gol72] D. Goldfarb. Modification methods for inverting matrices and solving systems of linear algebraic equations. Mathematics of Computation, 26(120):829–852, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1973:NSI

[Gol73] Moshe Goldberg. A note on the stability of an iterative finite-difference method for hyperbolic systems. Mathematics of Computation, 27(121):41–44, January 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldstein:1974:REH

[Gol74a] Marvin J. Goldstein. Reduction of the eigenproblem for Hermitian persymmetric matrices. Mathematics of Computation, 28(125):237–238, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldstein:1974:RPH

[Gol74b] Marvin J. Goldstein. Reduction of the pseudoinverse of a Hermitian persymmetric matrix. Mathematics of Computation, 28(127):715–717, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldfarb:1976:FVM

[Gol76a] Donald Goldfarb. Factorized variable metric methods for unconstrained optimization. Mathematics of Computation, 30 (136):796–811, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Golomb:1976:PS

[Gol76b] Solomon W. Golomb. Properties of the sequence 3·2^{n}+1. Mathematics of Computation, 30(135):657–663, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1977:BET

[Gol77a] Moshe Goldberg. On a boundary extrapolation theorem by Kreiss. Mathematics of Computation, 31(138):469–477, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldfine:1977:TSM

[Gol77b] Alan Goldfine. Taylor series methods for the solution of Volterra integral and integro-differential equations. Mathematics of Computation, 31(139):691–707, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gourlay:1970:NTM

[Gou70] A. R. Gourlay. A note on trapezoidal methods for the solution of initial value problems. Mathematics of Computation, 24(111):629–633, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Galimberti:1970:NDS

[GP70] G. Galimberti and V. Pereyra. Numerical differentiation and the solution of multidimensional Vandermonde systems. Mathematics of Computation, 24(110):357–364, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldstein:1971:DLM

[GP71] A. A. Goldstein and J. F. Price. On descent from local minima. Mathematics of Computation, 25(115):569–574, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gunzburger:1979:ECN

[GP79] Max D. Gunzburger and Robert J. Plemmons. Energy-conserving norms for the solution of hyperbolic systems of partial differential equations. Mathematics of Computation, 33(145):1–10, January 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guenther:1970:SOI

[GR70] R. B. Guenther and E. L. Roetman. Some observations on interpolation in higher dimensions. Mathematics of Computation, 24(111):517–522, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gerver:1979:SIN

[GR79] Joseph L. Gerver and L. Thomas Ramsey. Sets of integers with no long arithmetic progressions generated by the greedy algorithm. Mathematics of Computation, 33 (148):1353–1359, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greenstadt:1970:VVM

[Gre70a] J. Greenstadt. Variations on variablemetric methods. Mathematics of Computation, 24(109):1–22, January 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greville:1970:TTD

[Gre70b] T. N. E. Greville. Table for third-degree spline interpolation with equally spaced arguments. Mathematics of Computation, 24(109):179–183, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greenstadt:1972:QNM

[Gre72] John Greenstadt. A quasi-Newton method with no derivatives. Mathematics of Computation, 26(117):145–166, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greenstadt:1978:RDF

[Gre78] John Greenstadt. Revision of a derivativefree quasi-Newton method. Mathematics of Computation, 32(141):201–221, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grinstead:1978:MSC

[Gri78] Charles M. Grinstead. On a method of solving a class of Diophantine equations. Mathematics of Computation, 32 (143):936–940, July 1978. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Groetsch:1975:ECA

[Gro75] C. W. Groetsch. On existence criteria and approximation procedures for integral equations of the first kind. Mathematics of Computation, 29(132):1105–1108, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goodrich:1970:MSQ

[GS70] R. F. Goodrich and F. Stenger. Movable singularities and quadrature. Mathematics of Computation, 24(110):283–300, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guy:1975:WDA

[GS75] Richard K. Guy and J. L. Selfridge. What drives an aliquot sequence? Mathematics of Computation, 29(129):101–107, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1978:CMB

[GS78] Walter Gautschi and Josef Slavik. On the computation of modified Bessel function ratios. Mathematics of Computation, 32 (143):865–875, July 1978. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gottlieb:1976:DTF

[GT76] David Gottlieb and Eli Turkel. Dissipative two-four methods for time-dependent problems. Mathematics of Computation, 30(136):703–723, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1978:SIS

[GT78] Moshe Goldberg and Eitan Tadmor. Scheme-independent stability criteria for difference approximations of hyperbolic initial-boundary value problems. I. Mathematics of Computation, 32(144):1097– 1107, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gunzburger:1977:SGM

[Gun77] Max D. Gunzburger. On the stability of Galerkin methods for initial-boundary value problems for hyperbolic systems. Mathematics of Computation, 31(139): 661–675, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gupta:1971:CQR

[Gup71] Hansraj Gupta. Chains of quadratic residues. Mathematics of Computation, 25(114):379–382, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0297684-8.

Gupta:1976:SNH

[Gup76] G. K. Gupta. Some new high-order multistep formulae for solving stiff equations. Mathematics of Computation, 30 (135):417–432, July 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gupta:1978:ISD

[Gup78a] G. K. Gupta. Implementing secondderivative multistep methods using the Nordsieck polynomial representation. Mathematics of Computation, 32(141):13– 18, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gupta:1978:FDP

[Gup78b] Hansraj Gupta. Finite differences of the partition function. Mathematics of Computation, 32(144):1241–1243, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gupta:1979:PRH

[Gup79] G. K. Gupta. A polynomial representation of hybrid methods for solving ordinary differential equations. Mathematics of Computation, 33(148):1251–1256, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gustafson:1970:CEC

[Gus70] Sven-Åke Gustafson. Control and estimation of computational errors in the evaluation of interpolation formulae and quadrature rules. Mathematics of Computation, 24(112):847–854, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gustafsson:1975:CRD

[Gus75] Bertil Gustafsson. The convergence rate for difference approximations to mixed initial boundary value problems. Mathematics of Computation, 29(130):396–406, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gustafson:1979:SCL

[Gus79] Sven-Åke Gustafson. On stable calculation of linear functionals. Mathematics of Computation, 33(146):694–704, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gutknecht:1977:ESD

[Gut77] Martin H. Gutknecht. Existence of a solution to the discrete Theodorsen equation for conformal mappings. Mathematics of Computation, 31(138):478–480, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1971:CFE

[GW71] M. L. Glasser and V. E. Wood. A closed form evaluation of the elliptic integral. Mathematics of Computation, 25 (115):535–536, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gupta:1975:SNM

[GW75] G. K. Gupta and C. S. Wallace. Some new multistep methods for solving ordinary differential equations. Mathematics of Computation, 29(130):489–500, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gordon:1978:SMM

[GW78] William J. Gordon and James A. Wixom. Shepard’s method of “metric interpolation” to bivariate and multivariate interpolation. Mathematics of Computation, 32(141):253–264, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gupta:1979:NSS

[GW79] G. K. Gupta and C. S. Wallace. A new step-size changing technique for multistep methods. Mathematics of Computation, 33(145):125–138, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1974:CTQ

[GY74] Walter Gautschi and Hiroki Yanagiwara.

On Chebyshev-type quadratures. Mathematics of Computation, 28(125):125–134, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1975:AII

[Hab75] Seymour Haber. Adaptive integration and improper integrals. Mathematics of Computation, 29(131):806–809, July 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hadjidimos:1978:AOM

[Had78] Apostolos Hadjidimos. Accelerated overrelaxation method. Mathematics of Computation, 32(141):149–157, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1970:LBR

[Hag70] Peter Hagis, Jr. Lower bounds for relatively prime amicable numbers of opposite parity. Mathematics of Computation, 24(112):963–968, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1971:UAN

[Hag71] Peter Hagis, Jr. Unitary amicable numbers. Mathematics of Computation, 25 (116):915–918, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1973:LBS

[Hag73] Peter Hagis, Jr. A lower bound for the set of odd perfect numbers. Mathematics of Computation, 27(124):951–953, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hald:1978:ISL

[Hal78] Ole H. Hald. The inverse Sturm–Liouville problem and the Rayleigh–Ritz method. Mathematics of Computation, 32(143): 687–705, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Harten:1978:ACM

[Har78] Amiram Harten. The artificial compression method for computation of shocks and contact discontinuities. III. self-adjusting hybrid schemes. Mathematics of Computation, 32(142):363–389, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hays:1973:MDE

[Hay73] Richard O. Hays. Multi-dimensional extensions of the Chebyshev polynomials. Mathematics of Computation, 27(123): 621–624, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hallyburton:1975:TNF

[HB75a] John C. Hallyburton, Jr. and John Brillhart. Two new factors of Fermat numbers. Mathematics of Computation, 29 (129):109–112, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Headley:1975:DZG

[HB75b] V. B. Headley and V. K. Barwell. On the distribution of the zeros of generalized Airy functions. Mathematics of Computation, 29(131):863–877, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hald:1978:CVM

[Hd78] Ole Hald and Vincenza Mauceri del Prete.

Convergence of vortex methods for Euler’s equations. Mathematics of Computation, 32(143):791–809, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hedstrom:1975:MDS

[Hed75] G. W. Hedstrom. Models of difference schemes for u^{t} +u^{x} = 0 by partial differential equations. Mathematics of Computation, 29(132):969–977, October 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hendy:1974:ACF

[Hen74] M. D. Hendy. Applications of a continued fraction algorithm to some class number problems. Mathematics of Computation, 28(125):267–277, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hendy:1975:DIC

[Hen75] M. D. Hendy. The distribution of ideal class numbers of real quadratic fields. Mathematics of Computation, 29(132): 1129–1134, October 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hennart:1977:OSP

[Hen77a] J. P. Hennart. One-step piecewise polynomial multiple collocation methods for initial value problems. Mathematics of Computation, 31(137):24–36, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Henry:1977:CAM

[Hen77b] Jackson N. Henry. Comparison of algorithms for multivariate rational approximation. Mathematics of Computation, 31(138):485–494, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hetherington:1972:EBQ

[Het72] J. H. Hetherington. An error bound for quadratures. Mathematics of Computation, 26(119):695–698, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hetherington:1973:NEB

[Het73] J. H. Hetherington. Note on error bounds for numerical integration. Mathematics of Computation, 27(122):307–316, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hicks:1978:SAW

[Hic78] D. L. Hicks. Stability analysis of WONDY (a hydrocode based on the artificial viscosity method of von Neumann and Richtmyer) for a special case of Maxwell’s law. Mathematics of Computation, 32 (144):1123–1130, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hilbert:1973:MUA

[Hil73] Stephen Hilbert. A mollifier useful for approximations in Sobolev spaces and some applications to approximating solutions of differential equations. Mathematics of Computation, 27(121):81–89, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hill:1975:CAN

[Hil75] David R. Hill. On comparing Adams and natural spline multistep formulas. Mathematics of Computation, 29(131):741–745, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See table errata [Hil76].

Hill:1976:TEC

[Hil76] David R. Hill. Table errata: “On comparing Adams and natural spline multistep formulas” (Math. Comp. 29 (1975), 741–745). Mathematics of Computation, 30(134):381, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0386214-2;   http:
//www.ams.org/journals/mcom/1976-
30-134/S0025-5718-1976-0386214-2/
S0025-5718-1976-0386214-2.pdf; http:
//www.ams.org/mathscinet-getitem?

mr=386214. See [Hil75].

Hayes:1972:IMN

[HKK72] John K. Hayes, David K. Kahaner, and Richard G. Kellner. An improved method for numerical conformal mapping. Mathematics of Computation, 26(118):327–334, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hayes:1975:NCI

[HKK75] John K. Hayes, David K. Kahaner, and Richard G. Kellner. A numerical comparison of integral equations of the first and second kind for conformal mapping. Mathematics of Computation, 29(130):512–521, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1977:QAN

[HL77] Peter Hagis, Jr. and Graham Lord. Quasiamicable numbers. Mathematics of Computation, 31(138):608–611, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1973:LPD

[HM73] Peter Hagis, Jr. and Wayne L. McDaniel. On the largest prime divisor of an odd perfect number. Mathematics of Computation, 27(124):955–957, October 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1975:LPD

[HM75] Peter Hagis, Jr. and Wayne L. McDaniel. On the largest prime divisor of an odd perfect number. II. Mathematics of Computation, 29(131):922–924, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoskins:1977:INC

[HM77] W. D. Hoskins and D. S. Meek. The infinity norm of a certain type of symmetric circulant matrix. Mathematics of Computation, 31(139):733–737, July 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoff:1979:FDS

[Hof79] David Hoff. A finite difference scheme for a system of two conservation laws with artificial viscosity. Mathematics of Computation, 33(148):1171–1193, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoskins:1971:TTD

[Hos71] W. D. Hoskins. Table for thirddegree spline interpolation using equispaced knots. Mathematics of Computation, 25(116):797–801, October 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoskins:1973:BES

[Hos73] W. D. Hoskins. Boundary expansions for spline interpolation. Mathematics of Computation, 27(124):829–830, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Householder:1972:CNM

[Hou72] Alston S. Householder. On a condition for the nonsingularity of a matrix. Mathematics of Computation, 26(117):119–120, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Houstis:1977:AMC

[Hou77] E. N. Houstis. Application of method of collocation on lines for solving nonlinear hyperbolic problems. Mathematics of Computation, 31(138):443–456, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Howell:1973:AERa

[How73a] Jo Ann Howell. An algorithm for the exact reduction of a matrix to Frobenius form using modular arithmetic. I. Mathematics of Computation, 27(124):887–904, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Howell:1973:AERb

[How73b] Jo Ann Howell. An algorithm for the exact reduction of a matrix to Frobenius form using modular arithmetic. II. Mathematics of Computation, 27(124):905–920, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Howard:1977:NGR

[How77] F. T. Howard. Numbers generated by the reciprocal of e^{x} − x − 1. Mathematics of Computation, 31(138):581–598, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Howland:1978:RID

[How78] James Lucien Howland. The resultant iteration for determining the stability of a polynomial. Mathematics of Computation, 32(143):779–789, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoskins:1972:SPC

[HP72] W. D. Hoskins and P. J. Ponzo. Some properties of a class of band matrices. Mathematics of Computation, 26(118): 393–400, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Houstis:1978:CMF

[HP78] E. N. Houstis and T. S. Papatheodorou. A collocation method for Fredholm integral equations of the second kind. Mathematics of Computation, 32(141):159–173, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunter:1972:NEC

[HR72] D. B. Hunter and T. Regan. A note on the evaluation of the complementary error function. Mathematics of Computation, 26(118):539–541, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hager:1975:FBF

[HS75a] William W. Hager and Gilbert Strang. Free boundaries and finite elements in one dimension. Mathematics of Computation, 29(132):1020–1031, October 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hertling:1975:SRG

[HS75b] Jorg Hertling and Alexandru I. Schiop. On the stability of the Ritz–Galerkin method for Hammerstein equations. Mathematics of Computation, 29(130):484–488, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hinata:1974:NSP

[HSK74] Masahiro Hinata, Masaaki Shimasaki, and Takeshi Kiyono. Numerical solution of Plateau’s problem by a finite element method. Mathematics of Computation, 28(125):45–60, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Huddleston:1971:SRB

[Hud71] Robert E. Huddleston. Some relations between the values of a function and its first derivative at n abscissa points. Mathematics of Computation, 25(115):553–558, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hudson:1973:NDC

[Hud73] Richard H. Hudson. A note on Dirichlet characters. Mathematics of Computation, 27(124):973–975, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Huddleston:1974:CET

[Hud74] R. E. Huddleston. On the conditional equivalence of two starting methods for the second algorithm of Remez. Mathematics of Computation, 28(126):569–572, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hudson:1976:GCT

[Hud76] Richard H. Hudson. Generalizations of a classical theorem in number theory. Mathematics of Computation, 30(135):649–656, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hulme:1972:DGR

[Hul72a] Bernie L. Hulme. Discrete Galerkin and related one-step methods for ordinary differential equations. Mathematics of Computation, 26(120):881–891, October 1972.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hulme:1972:OSP

[Hul72b] Bernie L. Hulme. One-step piecewise polynomial Galerkin methods for initial value problems. Mathematics of Computation, 26(118):415–426, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunt:1974:CTE

[Hun74a] David C. Hunt. The character table of an eight-dimensional orthogonal group. Mathematics of Computation, 28(126): 659–660, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunt:1974:CTF

[Hun74b] David C. Hunt. The character table of Fischer’s simple group, M(23). Mathematics of Computation, 28(126):660–661, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunkins:1975:CPH

[Hun75a] Dalton R. Hunkins. Cubatures of precision 2k and 2k+1 for hyperrectangles. Mathematics of Computation, 29(132):1098– 1104, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunter:1975:SPO

[Hun75b] D. B. Hunter. Some properties of orthogonal polynomials. Mathematics of Computation, 29(130):559–565, April 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

I:1972:BRC

[I.72] E. I. Book review: Constructive Aspects of the Fundamental Theorem of Algebra, by B. Dejon and P. Henrici. Mathematics of Computation, 26(118): 588–589, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.jstor.org/
stable/2005189.

I:1975:BRA

[I.75] E. I. Book review: Alston S. Householder.

Principles of Numerical Analysis, Dover, New York, 1974, x + 274 pp., 21 cm. Price $4.00 (paperbound). Mathematics of Computation, 29(132):1152–1165, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1975-29-132/S0025-5718-75-99674-X;
 http://www.ams.org/journals/mcom/
1975-29-132/S0025-5718-75-99674-X/
S0025-5718-75-99674-X.pdf.

Ikebe:1975:ZRC

[Ike75] Yasuhiko Ikebe. The zeros of regular Coulomb wave functions and of their derivatives. Mathematics of Computation, 29(131):878–887, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ikeda:1979:SFA

[IN79] Tsutomu Ikeda and Tomoyasu Nakagawa.

On the SHASTA FCT algorithm for the equation \frac{∂ρ}{∂t} + \frac{∂}{∂x}(υ(ρ)ρ) = 0. Mathematics of Computation, 33(148):1157–1169, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iserles:1977:FFN

[Ise77] Arieh Iserles. Functional fitting — new family of schemes for integration of stiff O.D.E. Mathematics of Computation, 31 (137):112–123, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iserles:1978:SDP

[Ise78] Arieh Iserles. A-stability and dominating pairs. Mathematics of Computation, 32(141):19–33, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jain:1972:SSM

[Jai72] R. K. Jain. Some A-stable methods for stiff ordinary differential equations. Mathematics of Computation, 26(117):71–77, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jeltsch:1976:NCS

[Jel76] Rolf Jeltsch. A necessary condition for A-stability of multistep multiderivative methods. Mathematics of Computation, 30(136):739–746, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jeltsch:1977:MMU

[Jel77] Rolf Jeltsch. Multistep methods using higher derivatives and damping at infinity. Mathematics of Computation, 31(137): 124–138, January 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jeltsch:1978:CCM

[Jel78] Rolf Jeltsch. Complete characterization of multistep methods with an interval of periodicity for solving y′′ = f(x, y). Mathematics of Computation, 32(144):1108– 1114, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jerri:1977:CHF

[Jer77] A. J. Jerri. Computations of the Hill functions of higher order. Mathematics of Computation, 31(138):481–484, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jespersen:1977:LSD

[Jes77] Dennis C. Jespersen. A least squares decomposition method for solving elliptic equations. Mathematics of Computation, 31(140):873–880, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jones:1970:ACT

[JH70] William B. Jones and G. Hardy. Accelerating convergence of trigonometric approximations. Mathematics of Computation, 24 (111):547–560, July 1970. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jeans:1978:DFU

[JH78] N. S. Jeans and M. D. Hendy. Determining the fundamental unit of a pure cubic field given any unit. Mathematics of Computation, 32(143):925–935, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1973:VII

[Joh73] Wells Johnson. On the vanishing of the Iwasawa invariant μp for p < 8000. Mathematics of Computation, 27(122):387–396, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1974:ENP

[Joh74a] C. G. L. Johnson. Estimates near plane portions of the boundary for discrete elliptic boundary problems. Mathematics of Computation, 28(128):909–935, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1974:IPD

[Joh74b] Wells Johnson. Irregular prime divisors of the Bernoulli numbers. Mathematics of Computation, 28(126):653–657, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1975:IPC

[Joh75] Wells Johnson. Irregular primes and cyclotomic invariants. Mathematics of Computation, 29(129):113–120, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1977:CRF

[Joh77] Wells Johnson. On congruences related to the first case of Fermat’s last theorem. Mathematics of Computation, 31(138): 519–526, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1978:DFQ

[Joh78] Wells Johnson. On the p-divisibility of the Fermat quotients. Mathematics of Computation, 32(141):297–301, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnston:1972:NSH

[JP72] R. L. Johnston and S. K. Pal. The numerical solution of hyperbolic systems using bicharacteristics. Mathematics of Computation, 26(118):377–392, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnston:1974:NSL

[JP74] R. L. Johnston and S. K. Pal. Numerical solution of Lundquist equations of magnetohydrodynamics. Mathematics of Computation, 28(125):33–44, January 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jordan:1970:CPE

[JR70] J. H. Jordan and J. R. Rabung. A conjecture of Paul Erdős concerning Gaussian primes. Mathematics of Computation, 24(109):221–223, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1971:MQF

[JR71] L. W. Johnson and R. D. Riess. Minimal quadratures for functions of low-order continuity. Mathematics of Computation, 25(116):831–835, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jeltsch:1978:CMJ

[JS78] Rolf Jeltsch and Daniel Shanks. Corrigenda: Microfiche (january 1977) and Gauss’s Ternary Form Reduction and the 2-Sylow Subgroup. Mathematics of Computation, 32(144):1328–1329, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Sha71c].

Jones:1974:NSE

[JT74] William B. Jones and W. J. Thron. Numerical stability in evaluating continued fractions. Mathematics of Computation, 28 (127):795–810, July 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1975:EEF

[JT75] Claes Johnson and Vidar Thomee. Error estimates for a finite element approximation of a minimal surface. Mathematics of Computation, 29(130):343–349, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kadlec:1976:ESI

[Kad76] J. Kadlec. On the evaluation of some integrals occurring in scattering problems. Mathematics of Computation, 30(134): 263–277, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1976-30-134/
                                              S0025-5718-1976-0395164-7.

Kahaner:1970:CTQ

[Kah70] David K. Kahaner. Chebyshev type quadrature formulas. Mathematics of Computation, 24(111):571–574, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kahaner:1971:SPC

[Kah71] David K. Kahaner. Some polynomials for complex quadrature. Mathematics of Computation, 25(116):827–830, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kahaner:1972:MND

[Kah72a] David K. Kahaner. Minimum norm differentiation formulas with improved roundoff error bounds. Mathematics of Computation, 26(118):477–485, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kahaner:1972:NQA

[Kah72b] David K. Kahaner. Numerical quadrature by the ϵ-algorithm. Mathematics of Computation, 26(119):689–693, July 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kahng:1972:BA

[Kah72c] S. W. Kahng. Best L_{p} approximation. Mathematics of Computation, 26(118): 505–508, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kambo:1970:ENC

[Kam70] N. S. Kambo. Error of the Newton–Cotes and Gauss–Legendre quadrature formulas. Mathematics of Computation, 24(110): 261–269, April 1970. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Kammler:1973:CPC

[Kam73] David W. Kammler. Chebyshev polynomials corresponding to a semi-infinite interval and an exponential weight factor. Mathematics of Computation, 27(123):633–637, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Klopfenstein:1971:PAS

[KD71] R. W. Klopfenstein and C. B. Davis.

PECE algorithms for the solution of stiff systems of ordinary differential equations. Mathematics of Computation, 25(115): 457–473, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kearfott:1978:PCE

[Kea78] Baker Kearfott. A proof of convergence and an error bound for the method of bisection in R^{n}. Mathematics of Computation, 32(144):1147–1153, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1975:AMN

[Kel75] H. B. Keller. Approximation methods for nonlinear problems with application to two point boundary value problems. Mathematics of Computation, 29(130):464–474, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1979:SAE

[Kel79] Stuart R. Keller. On the surface area of the ellipsoid. Mathematics of Computation, 33(145):310–314, January 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kershaw:1970:IEI

[Ker70] D. Kershaw. Inequalities on the elements of the inverse of a certain tridiagonal matrix. Mathematics of Computation, 24 (109):155–158, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kershaw:1972:OAF

[Ker72] D. Kershaw. The orders of approximation of the first derivative of cubic splines at the knots. Mathematics of Computation, 26(117):191–198, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kim:1972:AEH

[Kim72] Shoon K. Kim. The asymptotic expansion of a hypergeometric function 2F_{2}(1, α; ρ_{1}, ρ_{2}; z). Mathematics of Computation, 26(120):963, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1971:INI

[Kin71] Richard F. King. Improved Newton iteration for integral roots. Mathematics of Computation, 25(114):299– 304, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0283981-9.

Kincaid:1972:NSO

[Kin72] David R. Kincaid. Norms of the successive overrelaxation method. Mathematics of Computation, 26(118):345–357, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1974:SLS

[Kin74] J. Thomas King. Semidiscrete leastsquares methods for second order parabolic problems with nonhomogeneous data. Mathematics of Computation, 28 (126):405–411, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kishore:1977:OPN

[Kis77] Masao Kishore. Odd perfect numbers not divisible by 3 are divisible by at least ten distinct primes. Mathematics of Computation, 31(137):274–279, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kishore:1978:OIF

[Kis78] Masao Kishore. Odd integers N with five distinct prime factors for which 2 − 10^{−12} < σ(N)/N < 2 + 10^{−12}. Mathematics of Computation, 32(141):303–309, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kumar:1974:QFS

[KJ74] Ravindra Kumar and M. K. Jain. Quadrature formulas for semi-infinite integrals. Mathematics of Computation, 28(126): 499–503, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Korn:1979:CFF

[KL79] David G. Korn and Jules J. Lambiotte, Jr.

Computing the Fast Fourier Transform on a vector computer. Mathematics of Computation, 33(147):977–992, July 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Klove:1975:SDE

[Klo75] Torleiv Klove. Sums of distinct elements from a fixed set. Mathematics of Computation, 29(132):1144–1149, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaufman:1978:CRD

[KLT78] Edwin H. Kaufman, Jr., David J. Leeming, and G. D. Taylor. A combined Remes-differential correction algorithm for rational approximation. Mathematics of Computation, 32(141):233–242, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kinderman:1977:CMS

[KMR77] A. J. Kinderman, J. F. Monahan, and J. G. Ramage. Computer methods for sampling from Student’s t distribution. Mathematics of Computation, 31(140):1009–1018, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knight:1970:TGQ

[KN70] C. J. Knight and A. C. R. Newbery. Trigonometric and Gaussian quadrature. Mathematics of Computation, 24(111): 575–581, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knuth:1970:NSP

[Knu70] Donald E. Knuth. A note on solid partitions. Mathematics of Computation, 24 (112):955–961, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knuth:1975:EEB

[Knu75] Donald E. Knuth. Estimating the efficiency of backtrack programs. Mathematics of Computation, 29(129):121–136, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1970:CZI

[Kol70] K. S. Kolbig. Complex zeros of an incomplete Riemann zeta function and of the incomplete gamma function. Mathematics of Computation, 24(111):679–696, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1972:CZT

[Kol72a] K. S. Kolbig. Complex zeros of two incomplete Riemann zeta functions. Mathematics of Computation, 26(118):551–565, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1972:ZIG

[Kol72b] K. S. Kolbig. On the zeros of the incomplete gamma function. Mathematics of Computation, 26(119):751–755, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolata:1979:EAF

[Kol79] William G. Kolata. Eigenvalue approximation by the finite element method: the method of Lagrange multipliers. Mathematics of Computation, 33(145):63–76, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kowalski:1971:FDM

[Kow71] Jan Krzysztof Kowalski. A finite-difference method for parabolic differential equations with mixed derivatives. Mathematics of Computation, 25(116):675–698, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1978:SGF

[KP78] H. B. Keller and V. Pereyra. Symbolic generation of finite difference formulas. Mathematics of Computation, 32(144):955–971, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kusterer:1979:SEP

[KR79] Roland Kusterer and Manfred Reimer.

Stable evaluation of polynomials in time log n. Mathematics of Computation, 33 (147):1019–1031, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kramarz:1978:GAS

[Kra78] Luis Kramarz. Global approximations to solutions of initial value problems. Mathematics of Computation, 32(141):35–59, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kreiss:1972:DAB

[Kre72a] Heinz-Otto O. Kreiss. Difference approximations for boundary and eigenvalue problems for ordinary differential equations. Mathematics of Computation, 26 (119):605–624, July 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kress:1972:GHC

[Kre72b] R. Kress. On the general Hermite cardinal interpolation. Mathematics of Computation, 26(120):925–933, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Krogh:1970:EAP

[Kro70] Fred T. Krogh. Efficient algorithms for polynomial interpolation and numerical differentiation. Mathematics of Computation, 24(109):185–190, January 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Krogh:1979:RRC

[Kro79] Fred T. Krogh. Recurrence relations for computing with modified divided differences. Mathematics of Computation, 33 (148):1265–1271, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1978:CET

[KS78] J. Thomas King and Steven M. Serbin.

Computational experiments and techniques for the penalty method with extrapolation. Mathematics of Computation, 32(141):111–126, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kopetzky:1979:TCB

[KS79] H. G. Kopetzky and W. Schwarz. Two conjectures of B. R. Santos concerning totitives. Mathematics of Computation, 33(146):841–844, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kellogg:1978:ASD

[KT78] R. Bruce Kellogg and Alice Tsan. Analysis of some difference approximations for a singular perturbation problem without turning points. Mathematics of Computation, 32(144):1025–1039, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kumar:1974:CQF

[Kum74] Ravindra Kumar. A class of quadrature formulas. Mathematics of Computation, 28 (127):769–778, July 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kuttler:1971:FOF

[Kut71] J. R. Kuttler. A fourth-order finitedifference approximation for the fixed membrane eigenproblem. Mathematics of Computation, 25(114):237–256, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/journals/
mcom/1971-25-114/S0025-5718-1971-
0301955-6.

Kincaid:1972:MSO

[KY72] David R. Kincaid and David M. Young. The modified successive overrelaxation method with fixed parameters. Mathematics of Computation, 26(119):705–717, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kyuno:1979:IAC

[Kyu79] Shoji Kyuno. An inductive algorithm to construct finite lattices. Mathematics of Computation, 33(145):409–421, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lakein:1974:CIC

[Lak74] Richard B. Lakein. Computation of the ideal class group of certain complex quartic fields. Mathematics of Computation, 28 (127):839–846, July 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lakein:1975:CIC

[Lak75] Richard B. Lakein. Computation of the ideal class group of certain complex quartic fields. II. Mathematics of Computation, 29(129):137–144, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1974:IUT

[Lal74] Mohan Lal. Iterates of the unitary totient function. Mathematics of Computation, 28(125):301–302, January 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lambert:1970:LMM

[Lam70] J. D. Lambert. Linear multistep methods with mildly varying coefficients. Mathematics of Computation, 24(109):81–93, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lam:1978:CQN

[Lam78] Binh Lam. On the convergence of a quasi-Newton method for sparse nonlinear systems. Mathematics of Computation, 32(142):447–451, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lander:1973:TT

[Lan73] L. J. Lander. Three thirteens. Mathematics of Computation, 27(122):397, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Larkin:1970:OAH

[Lar70] F. M. Larkin. Optimal approximation in Hilbert spaces with reproducing kernel functions. Mathematics of Computation, 24(112):911–921, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lardy:1973:EGS

[Lar73] L. J. Lardy. An extrapolated Gauss–Seidel iteration for Hessenberg matrices. Mathematics of Computation, 27(124):921–926, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Laurenzi:1973:DWF

[Lau73] Bernard J. Laurenzi. Derivatives of Whittaker functions W_{k,1/2} and M_{k,1/2} with respect to order K. Mathematics of Computation, 27(121):129–132, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LeRoux:1977:NCE

[Le 77] A. Y. Le Roux. A numerical conception of entropy for quasi-linear equations. Mathematics of Computation, 31(140):848–872, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LeRoux:1979:STP

[Le 79] Marie-Noëlle Le Roux. Semidiscretization in time for parabolic problems. Mathematics of Computation, 33(147):919–931, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1970:NDR

[Leh70] D. H. Lehmer. Note on the distribution of Ramanujan’s tau function. Mathematics of Computation, 24(111):741–743, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehman:1974:FLI

[Leh74] R. Sherman Lehman. Factoring large integers. Mathematics of Computation, 28(126):637–646, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1977:PFC

[Leh77] D. H. Lehmer. Prime factors of cyclotomic class numbers. Mathematics of Computation, 31(138):599–607, April 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leland:1975:AWN

[Lel75] Kenneth O. Leland. An algorithm for winding numbers for closed polygonal paths. Mathematics of Computation, 29(130):554–558, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lether:1970:CPC

[Let70] Frank G. Lether. Cross-product cubature error bounds. Mathematics of Computation, 24(111):583–592, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lether:1973:CEB

[Let73] F. G. Lether. Cubature error bounds for analytic functions. Mathematics of Computation, 27(123):655–668, July 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lether:1977:SCS

[Let77] F. G. Lether. Subtracting out complex singularities in numerical integration. Mathematics of Computation, 31(137):223–229, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewis:1970:CBO

[Lew70] James T. Lewis. Computation of best one-sided L_{1} approximation. Mathematics of Computation, 24(111):529–536, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewis:1972:CBM

[Lew72] James T. Lewis. Computation of best monotone approximations. Mathematics of Computation, 26(119):737–747, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewis:1975:RRA

[Lew75] James T. Lewis. Restricted range approximation and its application to digital filter design. Mathematics of Computation, 29(130):522–539, April 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1971:NCF

[LF71] M. Lal and A. Forbes. A note on Chowla’s function. Mathematics of Computation, 25(116):923–925, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lancaster:1972:NDS

[LF72] P. Lancaster and H. K. Farahat. Norms on direct sums and tensor products. Mathematics of Computation, 26(118):401–414, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1972:E

[LG72] M. Lal and P. Gillard. On the equation ϕ(n) = ϕ(n + k). Mathematics of Computation, 26(118):579–583, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1974:ECA

[LG74] M. Lal and P. Gillard. Evaluation of a constant associated with a parking problem. Mathematics of Computation, 28(126):561–564, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leung:1979:AFE

[LG79] K. V. Leung and S. S. Ghaderpanah. An application of the finite element approximation method to find the complex zeros of the modified Bessel function K_{n}(z). Mathematics of Computation, 33 (148):1299–1306, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Linz:1970:NCE

[Lin70] Peter Linz. On the numerical computation of eigenvalues and eigenvectors of symmetric integral equations. Mathematics of Computation, 24(112):905–910, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Linz:1972:MCB

[Lin72] Peter Linz. A method for computing Bessel function integrals. Mathematics of Computation, 26(118):509–513, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1978:FEH

[Lin78] Chih Bing Ling. Further evaluation of Howland integrals. Mathematics of Computation, 32(143):900–904, July 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Liron:1971:ISR

[Lir71] N. Liron. Infinite sums of roots for a class of transcendental equations and Bessel functions of order one-half. Mathematics of Computation, 25(116):769–781, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Livne:1975:SPD

[Liv75] Avishai Livne. Seven-point difference schemes for hyperbolic equations. Mathematics of Computation, 29(130):425–433, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Linz:1973:NCI

[LK73] Peter Linz and T. E. Kropp. A note on the computation of integrals involving products of trigonometric and Bessel functions. Mathematics of Computation, 27(124):871–872, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1971:NME

[LL71] Chih-Bing Ling and Jung Lin. A new method of evaluation of Howland integrals. Mathematics of Computation, 25(114):331–337, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0295537-2.

Ling:1972:EM

[LL72] Chih-Bing Ling and Jung Lin. On evaluation of moments of K_{ν}(t)/I_{ν}(t). Mathematics of Computation, 26(118):529–537, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1974:NFT

[LL74] D. H. Lehmer and Emma Lehmer. A new factorization technique using quadratic forms. Mathematics of Computation, 28(126):625–635, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1970:ISH

[LLS70] D. H. Lehmer, Emma Lehmer, and Daniel Shanks. Integer sequence having prescribed quadratic character. Mathematics of Computation, 24(110):433–451, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lavoie:1970:EED

[LM70] J. L. Lavoie and R. Michaud. Explicit expressions for the determinants of certain matrices. Mathematics of Computation, 24(109):151–154, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1978:TCC

[LM78] D. H. Lehmer and J. M. Masley. Table of the cyclotomic class numbers h^{∗}(p) and their factors for 200 < p < 521. Mathematics of Computation, 32(142):577–582, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lagarias:1979:CAF

[LO79] J. C. Lagarias and A. M. Odlyzko. On computing Artin L-functions in the critical strip. Mathematics of Computation, 33(147):1081–1095, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Long:1977:CDT

[Lon77] F. W. Long. Corrections to Dickson’s table of three dimensional division algebras over F_{5}. Mathematics of Computation, 31(140): 1031–1033, October 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1973:EME

[LP73] J. N. Lyness and K. K. Puri. The Euler–Maclaurin expansion for the simplex. Mathematics of Computation, 27 (122):273–293, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lentini:1974:VOF

[LP74] M. Lentini and V. Pereyra. A variable order finite difference method for nonlinear multipoint boundary value problems. Mathematics of Computation, 28 (128):981–1003, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lee:1973:CAR

[LR73] C. M. Lee and F. D. K. Roberts. A comparison of algorithms for rational l_{∞} approximation. Mathematics of Computation, 27(121):111–121, January 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigenda [LR79a].

Lee:1979:CCA

[LR79a] C. M. Lee and F. D. K. Roberts. Corrigenda: “A comparison of algorithms for rational l_{∞} approximation” (math. comp. 27 (1973), 111–121). Mathematics of Computation, 33(146):847, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [LR73].

Lesaint:1979:FEC

[LR79b] P. Lesaint and P. A. Raviart. Finite element collocation methods for first-order systems. Mathematics of Computation, 33 (147):891–918, July 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lipow:1972:HSC

[LS72] Peter R. Lipow and Frank Stenger. How slowly can quadrature formulas converge? Mathematics of Computation, 26(120): 917–922, October 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1975:GGQ

[LTK75] Yudell L. Luke, Bing Yuan Ting, and Marilyn J. Kemp. On generalized Gaussian quadrature. Mathematics of Computation, 29(132):1083–1093, October 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1970:FAE

[Luk70] Yudell L. Luke. Further approximations for elliptic integrals. Mathematics of Computation, 24(109):191–198, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1971:MTBa

[Luk71a] Yudell L. Luke. Miniaturized tables of Bessel functions. Mathematics of Computation, 25(114):323–330, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1971:MTBb

[Luk71b] Yudell L. Luke. Miniaturized tables of Bessel functions, II. Mathematics of Computation, 25(116):789–795, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1972:MTB

[Luk72] Yudell L. Luke. Miniaturized tables of Bessel functions. III. Mathematics of Computation, 26(117):237–240, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lunnon:1972:CEE

[Lun72] W. F. Lunnon. Corrigendum: Explicit expressions for the determinants of certain matrices (Math. Comp. 24 (1970), 151–

154) by J. L. Lavoie and R. Michaud. Mathematics of Computation, 26(117): 309, January 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1972-26-117/
S0025-5718-1972-0400682-5;   http:
//www.ams.org/journals/mcom/1972-
26-117/S0025-5718-1972-0400682-5/
S0025-5718-1972-0400682-5.pdf; https:
//www.ams.org/mathscinet-getitem?
mr=400682;   https://www.ams.org/
mathscinet/search/authors.html?authorName=
Lunnon%2C%20W.%20F..

Luskin:1979:CFE

[Lus79] Mitchell Luskin. Convergence of a finite element method for the approximation of normal modes of the oceans. Mathematics of Computation, 33(146):493–519, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lambin:1979:TGC

[LV79] P. Lambin and J. P. Vigneron. Tables for the Gaussian computation of ∫_{0}^{∞} x^{α}e^{−x}f(x)dx for values of α varying continuously between −1 and +1. Mathematics of Computation, 33(146):805–811, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1970:CFC

[Lyn70] J. N. Lyness. The calculation of Fourier coefficients by the Möbius inversion of the Poisson summation formula. part I: Functions whose early derivatives are continuous. Mathematics of Computation, 24 (109):101–135, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1971:AFF

[Lyn71a] J. N. Lyness. Adjusted forms of the Fourier coefficient asymptotic expansion and applications in numerical quadrature. Mathematics of Computation, 25(113):87–104, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1971:CFCa

[Lyn71b] J. N. Lyness. The calculation of Fourier coefficients by the Möbius inversion of the Poisson summation formula part II. piecewise continuous functions and functions with poles near the interval [0, 1]. Mathematics of Computation, 25(113):59–78, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1971:CFCb

[Lyn71c] J. N. Lyness. The calculation of Fourier coefficients by the Möbius inversion of the Poisson summation formula. part III: Functions having algebraic singularities. Mathematics of Computation, 25(115): 483–493, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1974:CTB

[Lyn74] J. N. Lyness. Computational techniques based on the Lanczos representation. Mathematics of Computation, 28 (125):81–123, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1976:EFE

[Lyn76] J. N. Lyness. An error functional expansion for N-dimensional quadrature with an integrand function singular at a point. Mathematics of Computation, 30(133):1– 23, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1979:ASI

[Lyn79a] J. N. Lyness. The affine scale invariance of minimization algorithms. Mathematics of Computation, 33(145):265–287, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1979:BME

[Lyn79b] J. N. Lyness. A bench mark experiment for minimization algorithms. Mathematics of Computation, 33(145):249–264, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

MacKinnon:1972:AEH

[Mac72a] Robert F. MacKinnon. The asymptotic expansions of Hankel transforms and related integrals. Mathematics of Computation, 26(118):515–527, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

MacKinnon:1972:TET

[Mac72b] Robert F. MacKinnon. Table errata:

Table of integrals, series, and products (English translation of the fourth Russian edition, Academic Press, New York, 1965) by I. S. Gradshteyn and I. M. Ryzhik. Mathematics of Computation, 26 (117):305, January 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1972-26-117/
S0025-5718-1972-0415970-6;   http:
//www.ams.org/journals/mcom/1972-

                                              26-117/S0025-5718-1972-0415970-6/
                                              S0025-5718-1972-0415970-6.pdf; https:
                                              //www.ams.org/mathscinet-getitem?
                                              mr=415970;   https://www.ams.org/
                                              mathscinet/search/authors.html?authorName=
                                              MacKinnon%2C%20Robert%20F..

MacLeod:1973:ICC

[Mac73] Malcolm A. MacLeod. Improved computation of cubic natural splines with equispaced knots. Mathematics of Computation, 27(121):107–109, January 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Madsen:1972:CSD

[Mad72] N. K. Madsen. Convergence of singular difference approximations for the discrete ordinate equations in x-y geometry. Mathematics of Computation, 26(117):45– 50, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Madsen:1973:PEB

[Mad73] Niel K. Madsen. A posteriori error bounds for numerical solutions of the neutron transport equation. Mathematics of Computation, 27(124):773–780, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mahler:1975:NSC

[Mah75] K. Mahler. A necessary and sufficient condition for transcendency. Mathematics of Computation, 29(129):145–153, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Makowski:1970:RCE

[Mak70] Andrzej Makowski. Remark on a conjecture of Erdős on binomial coefficients. Mathematics of Computation, 24(111):705, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Malcolm:1971:NCJ

[Mal71] Michael A. Malcolm. A note on a conjecture of L. J. Mordell. Mathematics of Computation, 25(114):375– 377, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0284395-8.

Mangasarian:1971:CGM

[Man71] O. L. Mangasarian. Convergent generalized monotone splitting of matrices. Mathematics of Computation, 25(116):649–653, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mansfield:1978:CTT

[Man78] Lois Mansfield. A Clough–Tocher type element useful for fourth order problems over nonpolygonal domains. Mathematics of Computation, 32(141):135–142, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Marder:1975:GPT

[Mar75] B. M. Marder. GAP — a PIC-type fluid code. Mathematics of Computation, 29(130):434–446, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morrison:1975:MFF

[MB75] Michael A. Morrison and John Brillhart.

A method of factoring and the factorization of F^{7}. Mathematics of Computation, 29(129):183–205, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McAllister:1971:DPC

[McA71] G. T. McAllister. The Dirichlet problem for a class of elliptic difference equations. Mathematics of Computation, 25 (116):655–673, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCabe:1974:CFE

[McC74] J. H. McCabe. A continued fraction expansion, with a truncation error estimate, for Dawson’s integral. Mathematics of Computation, 28(127):811–816, July 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCarthy:1977:OAE

[McC77] D. P. McCarthy. The optimal algorithm to evaluate x^{n} using elementary multiplication methods. Mathematics of Computation, 31(137):251–256, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCabe:1978:FCP

[McC78] J. H. McCabe. A further correspondence property of M fractions. Mathematics of Computation, 32(144):1303–1305, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McDaniel:1971:DOP

[McD71] Wayne L. McDaniel. On the divisibility of an odd perfect number by the sixth power of a prime. Mathematics of Computation, 25(114):383–385, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0296013-3.

McDaniel:1974:MPD

[McD74] Wayne L. McDaniel. On multiple prime divisors of cyclotomic polynomials. Mathematics of Computation, 28(127):847–850, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McKinney:1972:GRM

[McK72] Earl H. McKinney. Generalized recursive multivariate interpolation. Mathematics of Computation, 26(119):723–735, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McKay:1976:LDI

[McK76] John McKay. The largest degrees of irreducible characters of the symmetric group. Mathematics of Computation, 30 (135):624–631, July 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McKay:1979:TGF

[McK79] Brendan D. McKay. Transitive graphs with fewer than twenty vertices. Mathematics of Computation, 33(147):1101– 1121, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McNolty:1973:SPD

[McN73] Frank McNolty. Some probability density functions and their characteristic functions. Mathematics of Computation, 27 (123):495–504, July 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Merz:1972:RPH

[Mer72] Gerhard Merz. Remark on a paper by Huddleston. Mathematics of Computation, 26(120):901–902, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morris:1977:CPN

[MH77] A. G. Morris and T. S. Horner. Chebyshev polynomials in the numerical solution of differential equations. Mathematics of Computation, 31(140):881–891, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Micula:1973:ASD

[Mic73] G. Micula. Approximate solution of the differential equation y′′ = f(x, y) with spline functions. Mathematics of Computation, 27(124):807–816, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miel:1977:PEE

[Mie77] George Miel. On a posteriori error estimates. Mathematics of Computation, 31 (137):204–213, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mignotte:1974:IAF

[Mig74] M. Mignotte. An inequality about factors of polynomials. Mathematics of Computation, 28(128):1153–1157, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1971:ISN

[Mil71] R. K. Miller. On ignoring the singularity in numerical quadrature. Mathematics of Computation, 25(115):521–532, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1975:FSN

[Mil75a] J. C. P. Miller. On factorisation, with a suggested new approach. Mathematics of Computation, 29(129):155–172, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mills:1975:CFL

[Mil75b] W. H. Mills. Continued fractions and linear recurrences. Mathematics of Computation, 29(129):173–180, January 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miranker:1971:MDS

[Mir71] W. L. Miranker. Matricial difference schemes for integrating stiff systems of ordinary differential equations. Mathematics of Computation, 25(116):717–728, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monegato:1979:NEP

[ML79] G. Monegato and J. N. Lyness. On the numerical evaluation of a particular singular two-dimensional integral. Mathematics of Computation, 33(147):993–1002, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miranker:1974:SNS

[MM74] W. L. Miranker and J. P. Morreeuw.

Semianalytic numerical studies of turning points arising in stiff boundary value problems. Mathematics of Computation, 28 (128):1017–1034, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McGuire:1975:EIS

[MM75] G. R. McGuire and J. Ll. Morris. Explicitimplicit schemes for the numerical solution of nonlinear hyperbolic systems. Mathematics of Computation, 29(130):407–424, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Majda:1978:FMN

[MMO78] Andrew Majda, James McDonough, and Stanley Osher. The Fourier method for nonsmooth initial data. Mathematics of Computation, 32(144):1041–1081, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mock:1976:PMD

[Moc76] M. S. Mock. Projection methods with different trial and test spaces. Mathematics of Computation, 30(135):400–416, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moenck:1977:EAP

[Moe77] Robert T. Moenck. On the efficiency of algorithms for polynomial factoring. Mathematics of Computation, 31(137):235–250, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monegato:1976:NEG

[Mon76] Giovanni Monegato. A note on extended Gaussian quadrature rules. Mathematics of Computation, 30(136):812–817, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monegato:1978:PWE

[Mon78a] Giovanni Monegato. Positivity of the weights of extended Gauss–Legendre quadrature rules. Mathematics of Computation, 32(141):243–245, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monegato:1978:SRC

[Mon78b] Giovanni Monegato. Some remarks on the construction of extended Gaussian quadrature rules. Mathematics of Computation, 32(141):247–252, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monahan:1979:ENM

[Mon79] John F. Monahan. Extensions of von Neumann’s method for generating random variables. Mathematics of Computation, 33(147):1065–1069, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morris:1973:CPF

[Mor73] P. A. Morris. Computational problems in S-function theory. Mathematics of Computation, 27(124):965–971, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morrison:1975:NPT

[Mor75] Michael A. Morrison. A note on primality testing using Lucas sequences. Mathematics of Computation, 29(129):181–182, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morris:1979:DFR

[Mor79] Robert Morris. The dilogarithm function of a real argument. Mathematics of Computation, 33(146):778–787, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mossige:1972:TIP

[Mos72] Svein Mossige. Table of irreducible polynomials over GF[2] of degrees 10 through 20. Mathematics of Computation, 26(120): 1007–1009, October 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mosevich:1977:IDE

[Mos77] Jack W. Mosevich. Identifying differential equations by Galerkin’s method. Mathematics of Computation, 31(137):139–147, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McAllister:1977:ACS

[MPR77] David F. McAllister, Eli Passow, and John A. Roulier. Algorithms for computing shape preserving spline interpolations to data. Mathematics of Computation, 31 (139):717–725, July 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Matta:1971:UCE

[MR71] F. Matta and A. Reichel. Uniform computation of the error function and other related functions. Mathematics of Computation, 25(114):339–344, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0295538-4.

McAllister:1978:ICQ

[MR78] David F. McAllister and John A. Roulier. Interpolation by convex quadratic splines. Mathematics of Computation, 32(144): 1154–1162, October 1978. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McAllister:1973:DGF

[MS73a] G. T. McAllister and E. F. Sabotka. Discrete Green’s functions. Mathematics of Computation, 27(121):59–80, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Munteanu:1973:MCL

[MS73b] M. J. Munteanu and L. L. Schumaker. On a method of Carasso and Laurent for constructing interpolating splines. Mathematics of Computation, 27(122):317–325, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morgan:1975:NBP

[MS75] John Morgan and Ridgway Scott. A nodal basis for C^{1} piecewise polynomials of degree n ≥ 5. Mathematics of Computation, 29(131):736–740, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McNamee:1971:WCF

[MSW71] J. McNamee, F. Stenger, and E. L. Whitney. Whittaker’s cardinal function in retrospect. Mathematics of Computation, 25 (113):141–154, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

More:1976:GCB

[MT76] J. J. More and J. A. Trangenstein. On the global convergence of Broyden’s method. Mathematics of Computation, 30(135): 523–540, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meijerink:1977:ISM

[Mv77] J. A. Meijerink and H. A. van der Vorst.

An iterative solution method for linear systems of which the coefficient matrix is a symmetric M-matrix. Mathematics of Computation, 31(137):148–162, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miranker:1978:ME

[Mv78] W. L. Miranker and M. van Veldhuizen.

The method of envelopes. Mathematics of Computation, 32(142):453–496, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miranker:1976:AMS

[MW76] W. L. Miranker and G. Wahba. An averaging method for the stiff highly oscillatory problem. Mathematics of Computation, 30 (135):383–399, July 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Matthew:1977:SNP

[MW77] G. Matthew and H. C. Williams. Some new primes of the form k·2^{n}+1. Mathematics of Computation, 31(139):797–798, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McKay:1979:NSG

[MY79] John McKay and Kiang Chuen Young.

The nonabelian simple groups G, |G| < 10^{6}—minimal generating pairs. Mathematics of Computation, 33(146):812–814, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nasell:1974:IMB

[Nas74] Ingemar Nasell. Inequalities for modified Bessel functions. Mathematics of Computation, 28(125):253–256, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Natterer:1977:UCG

[Nat77] Frank Natterer. Uniform convergence of Galerkin’s method for splines on highly nonuniform meshes. Mathematics of Computation, 31(138):457–468, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nazareth:1976:GCD

[Naz76] Larry Nazareth. Generation of conjugate directions for unconstrained minimization without derivatives. Mathematics of Computation, 30(133):115–131, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ng:1970:CDF

[ND70] E. W. Ng and C. J. Devine. On the computation of Debye functions of integer orders. Mathematics of Computation, 24(110):405–407, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Netravali:1973:SAS

[Net73] Arun N. Netravali. Spline approximation to the solution of the Volterra integral equation of the second kind. Mathematics of Computation, 27(121):99–106, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nevanlinna:1978:CDA

[Nev78] Olavi Nevanlinna. On the convergence of difference approximations to nonlinear contraction semigroups in Hilbert spaces. Mathematics of Computation, 32(142): 321–334, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1970:TIC

[New70a] A. C. R. Newbery. Trigonometric interpolation and curve-fitting. Mathematics of Computation, 24(112):869–876, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1970:TFF

[New70b] Morris Newman. A table of the first factor for prime cyclotomic fields. Mathematics of Computation, 24(109):215–219, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1971:GIA

[New71] A. C. R. Newbery. A generalized interpolation algorithm. Mathematics of Computation, 25(115):549–552, July 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1973:EAF

[New73] A. C. R. Newbery. Error analysis for Fourier series evaluation. Mathematics of Computation, 27(123):639–644, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1974:EAP

[New74] A. C. R. Newbery. Error analysis for polynomial evaluation. Mathematics of Computation, 28(127):789–793, July 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1975:PES

[New75a] A. C. R. Newbery. Polynomial evaluation schemes. Mathematics of Computation, 29 (132):1046–1050, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1975:CSM

[New75b] Morris Newman. Congruence subgroups of the modular group. Mathematics of Computation, 29(129):207–213, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1976:AFR

[New76] Morris Newman. Asymptotic formulas related to free products of cyclic groups. Mathematics of Computation, 30(136): 838–846, October 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nadeau:1974:ANM

[NGF74] A. Nadeau, J. Guyard, and M. R. Feix. Algebraic-numerical method for the slightly perturbed harmonic oscillator. Mathematics of Computation, 28 (128):1057–1066, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nicolaides:1977:CAS

[Nic77] R. A. Nicolaides. On the l^{2} convergence of an algorithm for solving finite element equations. Mathematics of Computation, 31(140):892–906, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nicolaides:1978:MCI

[Nic78a] R. A. Nicolaides. On multigrid convergence in the indefinite case. Mathematics of Computation, 32(144):1082–1086, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nicolaides:1978:ORC

[Nic78b] R. A. Nicolaides. On the observed rate of convergence of an iterative method applied to a model elliptic difference equation. Mathematics of Computation, 32 (141):127–133, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nicolaides:1979:STP

[Nic79] R. A. Nicolaides. On some theoretical and practical aspects of multigrid methods. Mathematics of Computation, 33 (147):933–952, July 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1972:DPR

[Nie72] Harald Niederreiter. On the distribution of pseudo-random numbers generated by the linear congruential method. Mathematics of Computation, 26(119):793–795, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1974:DPR

[Nie74] Harald Niederreiter. On the distribution of pseudo-random numbers generated by the linear congruential method. II. Mathematics of Computation, 28(128):1117–1132, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1976:DPR

[Nie76] Harald Niederreiter. On the distribution of pseudo-random numbers generated by the linear congruential method. III. Mathematics of Computation, 30(135):571–597, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ninomiya:1970:BRS

[Nin70a] Ichizo Ninomiya. Best rational starting approximations and improved Newton iteration for the square root. Mathematics of Computation, 24(110):391–404, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ninomiya:1970:GRC

[Nin70b] Ichizo Ninomiya. Generalized rational Chebyshev approximation. Mathematics of Computation, 24(109):159–169, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1976:UDQ

[NO76] Harald Niederreiter and Charles F. Osgood. A uniform distribution question related to numerical analysis. Mathematics of Computation, 30(134):366– 370, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0398067-7.

Nosrati:1973:MBF

[Nos73] H. Nosrati. A modified Butcher formula for integration of stiff systems of ordinary differential equations. Mathematics of Computation, 27(122):267–272, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nicholson:1971:ENI

[NRRZ71] D. Nicholson, P. Rabinowitz, N. Richter, and D. Zeilberger. On the error in the numerical integration of Chebyshev polynomials. Mathematics of Computation, 25(113):79–86, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Neild:1974:RQF

[NS74a] Carol Neild and Daniel Shanks. On the 3-rank of quadratic fields and the Euler product. Mathematics of Computation, 28(125):279–291, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nevanlinna:1974:NTE

[NS74b] Olavi Nevanlinna and Aarne H. Sipila. A nonexistence theorem for explicit A-stable methods. Mathematics of Computation, 28 (128):1053–1055, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nitsche:1974:IER

[NS74c] Joachim A. Nitsche and Alfred H. Schatz.

Interior estimates for Ritz–Galerkin methods. Mathematics of Computation, 28 (128):937–958, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nashed:1974:CRA

[NW74] M. Z. Nashed and Grace Wahba. Convergence rates of approximate least squares solutions of linear integral and operator equations of the first kind. Mathematics of Computation, 28(125):69–80, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

ONeill:1978:NAC

[OD78] C. J. O’Neill and T. Downs. A numerical accuracy consideration in polynomial deflation. Mathematics of Computation, 32 (144):1144–1146, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ojala:1977:EAC

[Oja77] T. Ojala. Euclid’s algorithm in the cyclotomic field Q(ζ^{16}). Mathematics of Computation, 31(137):268–273, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Okamoto:1976:ANM

[Oka76] Masashi Okamoto. Asymptotic normality in Monte Carlo integration. Mathematics of Computation, 30(136):831–837, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Oliger:1974:FOD

[Oli74] Joseph Oliger. Fourth order difference methods for the initial boundary-value problem for hyperbolic equations. Mathematics of Computation, 28(125):15–25, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Oliver:1975:CLO

[Oli75] J. Oliver. A curiosity of low-order explicit Runge–Kutta methods. Mathematics of Computation, 29(132):1032–1036, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Oliger:1976:HDM

[Oli76] Joseph Oliger. Hybrid difference methods for the initial boundary-value problem for hyperbolic equations. Mathematics of Computation, 30(136):724–738, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Oren:1973:SSV

[Ore73] Shmuel S. Oren. Self-scaling variable metric algorithms without line search for unconstrained minimization. Mathematics of Computation, 27(124):873–885, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olver:1972:NBR

[OS72] F. W. J. Olver and D. J. Sookne. Note on backward recurrence algorithms. Mathematics of Computation, 26(120):941–947, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osborn:1975:SAC

[Osb75] John E. Osborn. Spectral approximation for compact operators. Mathematics of Computation, 29(131):712–725, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osher:1972:SPD

[Osh72] Stanley Osher. Stability of parabolic difference approximations to certain mixed initial boundary value problems. Mathematics of Computation, 26(117):13–39, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osler:1972:IAL

[Osl72a] Thomas J. Osler. The integral analog of the Leibniz rule. Mathematics of Computation, 26(120):903–915, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osler:1972:IAT

[Osl72b] Thomas J. Osler. An integral analogue of Taylor’s series and its use in computing Fourier transforms. Mathematics of Computation, 26(118):449–460, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osler:1975:ISC

[Osl75] Thomas J. Osler. An identity for simplifying certain generalized hypergeometric functions. Mathematics of Computation, 29(131):888–893, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

OLeary:1979:ELE

[OSV79] Dianne P. O’Leary, G. W. Stewart, and James S. Vandergraft. Estimating the largest eigenvalue of a positive definite matrix. Mathematics of Computation, 33 (148):1289–1292, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

OLeary:1979:CMM

[OW79] Dianne P. O’Leary and Olof Widlund. Capacitance matrix methods for the Helmholtz equation on general threedimensional regions. Mathematics of Computation, 33(147):849–879, July 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

P:1971:BRH

[P.71] B. N. P. Book review: A. S. Householder, KWIC Index for Matrices in Numerical Analysis, Volume I: Primary Authors A– J, viii + 124 pp., Volume II: Primary Authors K–Z, vii + 151 pp., 1969, Oak Ridge National Laboratory, Oak Ridge, Tennessee, 28 cm. Available from National Technical Information Service, U. S. Department of Commerce, Springfield, Virginia 22151. Price: Printed copy $3.00, Microfiche $0.65, each volume. Mathematics of Computation, 25(115):628, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1971-25-115/S0025-5718-71-99712-2;
 http://www.ams.org/journals/mcom/
1971-25-115/S0025-5718-71-99712-2/
S0025-5718-71-99712-2.pdf.

Page:1973:CSE

[Pag73] R. L. Page. On computing some extremal periodic positive-definite functions. Mathematics of Computation, 27(122):345–353, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paige:1973:EAM

[Pai73] C. C. Paige. An error analysis of a method for solving matrix equations. Mathematics of Computation, 27(122):355–359, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paige:1979:CSP

[Pai79] C. C. Paige. Computer solution and perturbation analysis of generalized linear least squares problems. Mathematics of Computation, 33(145):171–183, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parnes:1972:CZM

[Par72] R. Parnes. Complex zeros of the modified Bessel function K_{n}(Z). Mathematics of Computation, 26(120):949–953, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1974:RQI

[Par74] B. N. Parlett. The Rayleigh quotient iteration and some generalizations for nonnormal matrices. Mathematics of Computation, 28(127):679–693, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parry:1977:RQF

[Par77] Charles J. Parry. Real quadratic fields with class numbers divisible by five. Mathematics of Computation, 31(140):1019–1029, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parry:1978:CNR

[Par78] Charles J. Parry. On the class number of relative quadratic fields. Mathematics of Computation, 32(144):1261–1270, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paul:1972:AFA

[Pau72] A. K. Paul. Anharmonic frequency analysis. Mathematics of Computation, 26(118): 437–447, April 1972. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Piessens:1974:NOA

[PB74a] Robert Piessens and Maria Branders. A note on the optimal addition of abscissas to quadrature formulas of Gauss and Lobatto type. Mathematics of Computation, 28(125):135–139, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See supplement [PB74b].

Piessens:1974:SNO

[PB74b] Robert Piessens and Maria Branders. Supplement to A Note on the Optimal Addition of Abscissas to Quadrature Formulas of Gauss and Lobatto Type. Mathematics of Computation, 28(125):344–347, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [PB74a].

Pereyra:1970:HAN

[Per70] Victor Pereyra. Highly accurate numerical solution of casilinear elliptic boundaryvalue problems in n dimensions. Mathematics of Computation, 24(112):771–783, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Petty:1973:WNA

[Pet73] J. V. Petty. A winding number algorithm for closed polygonal paths. Mathematics of Computation, 27(122):333–337, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pexton:1970:CIC

[Pex70] Robert L. Pexton. Computer investigation of Coulomb wave functions. Mathematics of Computation, 24(110):409–411, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Piessens:1975:CFD

[PH75] Robert Piessens and Ann Haegemans. Cubature formulas of degree nine for symmetric planar regions. Mathematics of Computation, 29(131):810–815, July 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Phillips:1970:GLE

[Phi70] David L. Phillips. Generalized logarithmic error and Newton’s method for the m-th root. Mathematics of Computation, 24(110):383–389, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Phillips:1971:TDH

[Phi71] James L. Phillips. The triangular decomposition of Hankel matrices. Mathematics of Computation, 25(115):599–602, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Phillips:1973:EAD

[Phi73] James L. Phillips. Error analysis for direct linear integral equation methods. Mathematics of Computation, 27(124):849–859, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Philip:1979:NKP

[Phi79] J. R. Philip. Note on the Kelvin phase functions. Mathematics of Computation, 33(145):337–341, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pelios:1972:MEC

[PK72] A. Pelios and R. W. Klopfenstein. Minimal error constant numerical differentiation (N.D.) formulas. Mathematics of Computation, 26(118):467–475, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plemmons:1972:MIA

[Ple72] Robert J. Plemmons. Monotonicity and iterative approximations involving rectangular matrices. Mathematics of Computation, 26(120):853–858, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pollard:1971:FFT

[Pol71] J. M. Pollard. The Fast Fourier Transform in a finite field. Mathematics of Computation, 25(114):365–374, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0301966-0.

Pollard:1978:MCM

[Pol78] J. M. Pollard. Monte Carlo methods for index computation (modp). Mathematics of Computation, 32(143):918–924, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pomerance:1975:SLP

[Pom75] Carl Pomerance. The second largest prime factor of an odd perfect number. Mathematics of Computation, 29(131):914–921, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pomerance:1979:PNG

[Pom79] Carl Pomerance. The prime number graph. Mathematics of Computation, 33 (145):399–408, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Poole:1975:GMA

[Poo75] George D. Poole. Generalized M-matrices and applications. Mathematics of Computation, 29(131):903–910, July 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Penney:1974:SEC

[PP74] David E. Penney and Carl Pomerance.

A search for elliptic curves with large rank. Mathematics of Computation, 28 (127):851–853, July 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Penney:1975:TEC

[PP75] David E. Penney and Carl Pomerance.

Three elliptic curves with rank at least seven. Mathematics of Computation, 29 (131):965–967, July 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1977:MFIa

[PP77a] Wilhelm Plesken and Michael Pohst. On maximal finite irreducible subgroups of GL(n, Z): I. the five and seven dimensional cases. Mathematics of Computation, 31(138):536–551, April 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1977:MFIb

[PP77b] Wilhelm Plesken and Michael Pohst. On maximal finite irreducible subgroups of GL(n, Z): II. the six dimensional case. Mathematics of Computation, 31(138): 552–573, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pereyra:1977:HOF

[PPW77] Victor Pereyra, Wlodzimierz Proskurowski, and Olof Widlund. High order fast Laplace solvers for the Dirichlet problem on general regions. Mathematics of Computation, 31(137):1–16, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Prothero:1974:SAO

[PR74] A. Prothero and A. Robinson. On the stability and accuracy of one-step methods for solving stiff systems of ordinary differential equations. Mathematics of Computation, 28(125):145–162, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pruess:1973:SLB

[Pru73] Steven A. Pruess. Solving linear boundary value problems by approximating the coefficients. Mathematics of Computation, 27 (123):551–561, July 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pruess:1979:AES

[Pru79] Steven Pruess. Alternatives to the exponential spline in tension. Mathematics of Computation, 33(148):1273–1281, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pereyra:1973:ECM

[PS73] V. Pereyra and G. Scherer. Efficient computer manipulation of tensor products with applications to multidimensional approximation. Mathematics of Computation, 27(123):595–605, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pexton:1977:RTTb

[PS77a] Robert L. Pexton and Arno D. Steiger. Roots of two transcendental equations determining the frequency spectra of standing spherical electromagnetic waves. Mathematics of Computation, 31(140):1000– 1002, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pexton:1977:RTTa

[PS77b] Robert L. Pexton and Arno D. Steiger.

Roots of two transcendental equations involving spherical Bessel functions. Mathematics of Computation, 31(139):752–753, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Preparata:1977:CCF

[PS77c] F. P. Preparata and D. V. Sarwate. Computational complexity of Fourier transforms over finite fields. Mathematics of Computation, 31(139):740–751, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pexton:1978:RTT

[PS78] Robert L. Pexton and Arno D. Steiger.

Roots of two transcendental equations as functions of a continuous real parameter. Mathematics of Computation, 32(142): 511–518, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1979:LAS

[PS79a] B. N. Parlett and D. S. Scott. The Lánczos algorithm with selective orthogonalization. Mathematics of Computation, 33(145):217–238, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pexton:1979:DRT

[PS79b] Robert L. Pexton and Arno D. Steiger.

Degenerate roots of three transcendental equations involving spherical Bessel functions. Mathematics of Computation, 33(147):1041–1048, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Proskurowski:1976:NSH

[PW76] Włodzimierz Proskurowski and Olof Widlund. On the numerical solution of Helmholtz’s equation by the capacitance matrix method. Mathematics of Computation, 30(135):433–468, July 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Petschek:1979:MCF

[PWKW79] A. G. Petschek, R. E. Williamson, W. J. Krauser, and P. C. White. Monte Carlo functional expansion calculation of free molecular flows. Mathematics of Computation, 33(146):765–777, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pohst:1977:ENG

[PZ77] Michael Pohst and Hans Zassenhaus. An effective number geometric method of computing the fundamental units of an algebraic number field. Mathematics of Computation, 31(139):754–770, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rakotch:1975:NSE

[Rak75] E. Rakotch. Numerical solution for eigenvalues and eigenfunctions of a Hermitian kernel and an error estimate. Mathematics of Computation, 29(131):794–805, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rakotch:1978:IEE

[Rak78] E. Rakotch. Improved error estimates for numerical solutions of symmetric integral equations. Mathematics of Computation, 32(142):399–404, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rall:1976:QCI

[Ral76] L. B. Rall. A quadratically convergent iteration method for computing zeros of operators satisfying autonomous differential equations. Mathematics of Computation, 30(133):112–114, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ramos:1971:REA

[Ram71] George U. Ramos. Roundoff error analysis of the Fast Fourier Transform. Mathematics of Computation, 25(116):757–768, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rose:1973:MAF

[RB73] Kermit Rose and Simcha Brudno. More about four biquadrates equal one biquadrate. Mathematics of Computation, 27(123):491–494, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riddell:1978:SEB

[RC78] J. Riddell and C. Chan. Some extremal 2-bases. Mathematics of Computation, 32(142):630–634, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reddien:1976:FDA

[Red76] G. W. Reddien. Finite-difference approximations to singular Sturm–Liouville eigenvalue problems. Mathematics of Computation, 30(134):278–282, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0403235-1.

Reinsch:1971:SRA

[Rei71] Christian H. Reinsch. A stable, rational QR algorithm for the computation of the eigenvalues of an Hermitian, tridiagonal matrix. Mathematics of Computation, 25 (115):591–597, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riesel:1970:SCR

[RG70] Hans Riesel and Gunnar Gohl. Some calculations related to Riemann’s prime number formula. Mathematics of Computation, 24(112):969–983, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rainey:1972:TCN

[RH72] J. W. Rainey and G. J. Habetler. Tridiagonalization of completely nonnegative matrices. Mathematics of Computation, 26 (117):121–128, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rheinboldt:1972:BRH

[Rhe72] W. Rheinboldt. Book review: A. S. Householder, KWIC Index for the Numerical Treatment of Nonlinear Equations, Oak Ridge National Laboratory, Oak Ridge, Tennessee, 1970, vii + 129 pp., 28 cm. Available from U. S. Department of Commerce, Springfield, Va. 22151. Price: printed copy $3.00; microfiche $0.65. Mathematics of Computation, 26(117):287–303, January 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/1972-26-
117/S0025-5718-72-99110-7;   http:
//www.ams.org/journals/mcom/1972-
26-117/S0025-5718-72-99110-7/S0025-
5718-72-99110-7.pdf.

Rheinboldt:1976:MIC

[Rhe76] Werner C. Rheinboldt. On measures of ill-conditioning for nonlinear equations. Mathematics of Computation, 30(133):

104–111, January 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1971:MRN

[Ric71] John R. Rice. Matrix representations of nonlinear equation iterations–application to parallel computation. Mathematics of Computation, 25(116):639–647, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riley:1971:HKG

[Ril71] Robert Riley. Homomorphisms of knot groups on finite groups. Mathematics of Computation, 25(115):603–619, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riordan:1975:DCC

[Rio75] John Riordan. The distribution of crossings of chords joining pairs of 2n points on a circle. Mathematics of Computation, 29(129):215–222, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rissanen:1973:ATD

[Ris73] J. Rissanen. Algorithms for triangular decomposition of block Hankel and Toeplitz matrices with application to factoring positive matrix polynomials. Mathematics of Computation, 27(121):147–154, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabung:1970:CPR

[RJ70] J. R. Rabung and J. H. Jordan. Consecutive power residues or nonresidues. Mathematics of Computation, 24(111):737–740, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Robertson:1971:LTP

[Rob71] T. N. Robertson. The linear two-point boundary-value problem on an infinite interval. Mathematics of Computation, 25 (115):475–481, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Robinson:1975:SCT

[Rob75] Raphael M. Robinson. The structure of certain triple systems. Mathematics of Computation, 29(129):223–241, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rosman:1971:ECA

[Ros71a] Bernard H. Rosman. Exponential Chebyshev approximation on finite subsets of [0, 1]. Mathematics of Computation, 25 (115):575–577, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ross:1971:CSG

[Ros71b] George G. Ross. On the computation of some Grunsky coefficients relevant to the Bieberbach conjecture. Mathematics of Computation, 25(116):733–741, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rosenblatt:1975:MSS

[Ros75] M. Rosenblatt. Multiply schemes and shuffling. Mathematics of Computation, 29 (131):929–934, July 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rosenberger:1976:UMN

[Ros76] Gerhard Rosenberger. The uniqueness of the Markoff numbers. Mathematics of Computation, 30(134):361– 365, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0396413-1.

Rubin:1970:TDS

[RP70] Ephraim L. Rubin and Stanley Preiser.

Three-dimensional second-order accurate difference schemes for discontinuous hydrodynamic flows. Mathematics of Computation, 24(109):57–63, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rosen:1971:SNE

[RP71] D. Rosen and G. S. Patterson, Jr. Some numerical evidence concerning the uniqueness of the Markov numbers. Mathematics of Computation, 25(116):919–921, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1970:APM

[RR70a] Philip Rabinowitz and Nira Richter.

Asymptotic properties of minimal integration rules. Mathematics of Computation, 24(111):593–609, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1970:CTI

[RR70b] Philip Rabinowitz and Nira Richter.

Chebyshev-type integration rules of minimum norm. Mathematics of Computation, 24(112):831–845, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1970:NEC

[RR70c] Philip Rabinowitz and Nira Richter. New error coefficients for estimating quadrature errors for analytic functions. Mathematics of Computation, 24(111):561–570, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rosenberg:1975:LBA

[RS75a] Ivo G. Rosenberg and Frank Stenger. A lower bound on the angles of triangles constructed by bisecting the longest side. Mathematics of Computation, 29(130): 390–395, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rosser:1975:SBC

[RS75b] J. Barkley Rosser and Lowell Schoenfeld. Sharper bounds for the Chebyshev functions Θ(x) and Ψ(x). Mathematics of Computation, 29(129):243–269, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rowen:1978:SEM

[RS78] Louis Halle Rowen and Uri Schild. A scalar expression for matrices with symplectic involution. Mathematics of Computation, 32(142):607–613, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Raviart:1977:PHF

[RT77] P. A. Raviart and J. M. Thomas. Primal hybrid finite element methods for 2nd order elliptic equations. Mathematics of Computation, 31(138):391–413, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ruhe:1974:MEP

[Ruh74] Axel Ruhe. SOR-methods for the eigenvalue problem with large sparse matrices. Mathematics of Computation, 28 (127):695–710, July 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ruhe:1979:IAB

[Ruh79] Axel Ruhe. Implementation aspects of band Lánczos algorithms for computation of eigenvalues of large sparse symmetric matrices. Mathematics of Computation, 33(146):680–687, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rump:1979:PMR

[Rum79] Siegfried M. Rump. Polynomial minimum root separation. Mathematics of Computation, 33(145):327–336, January 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rowland:1972:ECS

[RV72] J. H. Rowland and Y. L. Varol. Exit criteria for Simpson’s compound rule. Mathematics of Computation, 26(119):699–703, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rheinboldt:1973:SAP

[RV73] Werner C. Rheinboldt and James S. Vandergraft. A simple approach to the Perron–Frobenius theory for positive operators on general partially-ordered finitedimensional linear spaces. Mathematics of Computation, 27(121):139–145, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Russell:1975:CGM

[RV75] R. D. Russell and J. M. Varah. A comparison of global methods for linear two-point boundary value problems. Mathematics of Computation, 29(132):1007–1019, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

S:1975:BRR

[S.75] J. S. Book review: R. W. Hamming, Numerical Methods for Scientists and Engineers, 2nd ed., McGraw-Hill Book Co., New York, 1973, ix + 721 pp., 24 cm. Mathematics of Computation, 29(130):

648–649, May 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1975-29-130/
S0025-5718-75-99678-7/S0025-5718-
75-99678-7.pdf.

S:1978:BRM

[S.78] H. J. S. Book review: Martin Gutknecht, Peter Henrici, Peter Lauchli, and Hans-Rudolf Schwarz, Heinz Rutishauser: Vorlesungen Uber Numerische Mathematik Vols. 1 and 2. Mathematics of Computation, 32(142):651–652, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

/www.jstor.org/stable/pdf/2006176.
pdf.

Sackett:1971:NSP

[Sac71] Gary G. Sackett. Numerical solution of a parabolic free boundary problem arising in statistical decision theory. Mathematics of Computation, 25(115):425–434, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sag:1970:CIM

[Sag70] T. W. Sag. Chebyshev iteration methods for integral equations of the second kind. Mathematics of Computation, 24(110):341–355, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salkauskas:1971:EQF

[Sal71a] K. Salkauskas. Existence of quadrature formulae with almost equal weights. Mathematics of Computation, 25(113):105–109, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1971:FBH

[Sal71b] Herbert E. Salzer. Formulas for bivariate hyperosculatory interpolation. Mathematics of Computation, 25(113):119–133, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salkauskas:1973:AIC

[Sal73] K. Salkauskas. Almost-interpolatory Chebyshev quadrature. Mathematics of Computation, 27(123):645–654, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1974:SPO

[Sal74] Herbert E. Salzer. Some problems in optimally stable Lagrangian differentiation. Mathematics of Computation, 28 (128):1105–1115, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salamin:1976:CUA

[Sal76a] Eugene Salamin. Computation of π using arithmetic-geometric mean. Mathematics of Computation, 30(135):565–570, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1976:CIS

[Sal76b] Herbert E. Salzer. Converting interpolation series into Chebyshev series by recurrence formulas. Mathematics of Computation, 30(134):295–302, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0395159-3.

Salzer:1977:OHC

[Sal77] Herbert E. Salzer. Osculatory and hyperosculatory cubature formulas. Mathematics of Computation, 31(137):230–234, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sameh:1971:JJL

[Sam71] Ahmed H. Sameh. On Jacobi and Jacobilike algorithms for a parallel computer. Mathematics of Computation, 25(115): 579–590, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Samaddar:1974:SII

[Sam74] S. N. Samaddar. Some integrals involving associated Legendre functions. Mathematics of Computation, 28(125):257–263, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sammon:1977:DLS

[Sam77] Peter H. Sammon. A discrete least squares method. Mathematics of Computation, 31(137):60–65, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schlegel:1970:EIT

[Sch70a] P. Schlegel. The explicit inverse of a tridiagonal matrix. Mathematics of Computation, 24(111):665, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schubert:1970:MQN

[Sch70b] L. K. Schubert. Modification of a quasi-Newton method for nonlinear equations with a sparse Jacobian. Mathematics of Computation, 24(109):27–30, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schultz:1970:ESF

[Sch70c] Martin H. Schultz. Elliptic spline functions and the Rayleigh–Ritz–Galerkin method. Mathematics of Computation, 24(109):65– 80, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schultz:1970:EBP

[Sch70d] Martin H. Schultz. Error bounds for polynomial spline interpolation. Mathematics of Computation, 24(111):507–515, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1974:OCR

[Sch74] Alfred H. Schatz. An observation concerning Ritz–Galerkin methods with indefinite bilinear forms. Mathematics of Computation, 28(128):959–962, October 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schett:1976:PTS

[Sch76a] Alois Schett. Properties of the Taylor series expansion coefficients of the Jacobian elliptic functions. Mathematics of Computation, 30(133):143–147, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schoenfeld:1976:SBC

[Sch76b] Lowell Schoenfeld. Sharper bounds for the Chebyshev functions θ(x) and ψ(x). II. Mathematics of Computation, 30(134): 337–360, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0457374-X.

Schett:1977:RFT

[Sch77] Alois Schett. Recurrence formula of the Taylor series expansion coefficients of the Jacobian elliptic functions. Mathematics of Computation, 31(140):1003–1005, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schmidt:1978:AAT

[Sch78a] Paul W. Schmidt. An asymptotic approximation for a type of Fourier integral. Mathematics of Computation, 32 (144):1171–1182, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schmidt:1978:EI

[Sch78b] Paul W. Schmidt. Evaluation of the integral ∫_{0}^{∞} ^{\frac{t^{2α−1}J_{ν}(x\sqrt{1+t^{2}})}{(1+t^{2})^{α+β−1}}}dt. Mathematics of Computation, 32(141):265–269, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schnabel:1978:MNS

[Sch78c] Robert B. Schnabel. Minimum norm symmetric quasi-Newton updates restricted to subspaces. Mathematics of Computation, 32(143):829–837, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schonfelder:1978:CEE

[Sch78d] J. L. Schonfelder. Chebyshev expansions for the error and related functions. Mathematics of Computation, 32(144):1232– 1240, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scott:1976:OEF

[Sco76] Ridgway Scott. Optimal L^{∞} estimates for the finite element method on irregular meshes. Mathematics of Computation, 30(136):681–697, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scott:1979:HML

[Sco79] D. S. Scott. How to make the Lánczos algorithm converge slowly. Mathematics of Computation, 33(145):239–247, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sacks-Davis:1977:EES

[SD77] R. Sacks-Davis. Error estimates for a stiff differential equation procedure. Mathematics of Computation, 31(140):939–953, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Serbin:1975:CIL

[Ser75] Steven M. Serbin. Computational investigations of least-squares type methods for the approximate solution of boundary value problems. Mathematics of Computation, 29(131):777–793, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Snider:1974:AAA

[SF74] Arthur David Snider and Gary Charles Fleming. Approximation by aliasing with application to “Certaine” stiff differential equations. Mathematics of Computation, 28(126):465–473, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Strecok:1972:HPE

[SG72] A. J. Strecok and J. A. Gregory. High precision evaluation of the irregular Coulomb wave functions. Mathematics of Computation, 26(120):955–961, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanno:1970:CQN

[Sha70] D. F. Shanno. Conditioning of quasi-Newton methods for function minimization. Mathematics of Computation, 24 (111):647–656, July 1970. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1971:CGI

[Sha71a] Daniel Shanks. Corrigendum to: On generating infinitely many solutions of the Diophantine equation A^{6} + B^{6} + C^{6} = D^{6} + E^{6} + F^{6}, (Math. Comp. 24 (1970), 453–454) by Simcha Brudno. Mathematics of Computation, 25(114):

409, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0400638-1.

Shanks:1971:CSS

[Sha71b] Daniel Shanks. Corrigendum to: “Statistical study of digits of some square roots of integers in various bases” (Math. Comp. 24 (1970), 455–473) by W. A. Beyer, N. Metropolis and J. R. Neergaard. Mathematics of Computation, 25 (114):409, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0400639-3.

Shanks:1971:GTF

[Sha71c] Daniel Shanks. Gauss’s ternary form reduction and the 2-Sylow subgroup. Mathematics of Computation, 25(116):837–853, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [JS78].

Shampine:1973:LES

[Sha73a] L. F. Shampine. Local extrapolation in the solution of ordinary differential equations. Mathematics of Computation, 27(121):91– 97, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shapiro:1973:IVI

[Sha73b] Harold N. Shapiro. Integer vectors with interprimed components. Mathematics of Computation, 27(123):455–462, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1974:LPD

[Sha74a] L. F. Shampine. Limiting precision in differential equation solvers. Mathematics of Computation, 28(125):141–144, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1974:SCF

[Sha74b] Daniel Shanks. The simplest cubic fields.

Mathematics of Computation, 28(128): 1137–1152, October 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1975:CAE

[Sha75a] Daniel Shanks. Calculation and applications of Epstein zeta functions. Mathematics of Computation, 29(129):271–287, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1975:CCA

[Sha75b] Daniel Shanks. Corrigenda: Calculation and applications of Epstein zeta functions (Math. Comp. 29 (1975), 271–287). Mathematics of Computation, 29(132): 1167, October 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1975-29-132/
S0025-5718-1975-0409358-4;   http:
//www.ams.org/journals/mcom/1975-
29-132/S0025-5718-1975-0409358-4/
S0025-5718-1975-0409358-4.pdf; https:
//www.ams.org/mathscinet-getitem?
mr=409358;   https://www.ams.org/
mathscinet/search/authors.html?authorName=
Shanks%2C%20Daniel.

Shapiro:1975:LF

[Sha75c] Ralph Shapiro. Linear filtering. Mathematics of Computation, 29(132):1094– 1097, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1976:TER

[Sha76a] D. Shanks. Table errata: “Regular continued fractions for π and γ”, (Math. Comp. 25 (1971), 403); “Rational approximations to π” (ibid. 25 (1971), 387– 392) by K. Y. Choong, D. E. Daykin and C. R. Rathbone. Mathematics of Computation, 30(134):381, 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1976-30-134/
                                              S0025-5718-1976-0386215-4.

Shanks:1976:CGQ

[Sha76b] Daniel Shanks. Class groups of the quadratic fields found by F. Diaz y Diaz. Mathematics of Computation, 30(133): 173–178, January 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1978:LPD

[Sha78] L. F. Shampine. Limiting precision in differential equation solvers. II. sources of trouble and starting a code. Mathematics of Computation, 32(144):1115–1122, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shippee:1978:FNF

[Shi78] D. E. Shippee. Four new factors of Fermat numbers. Mathematics of Computation, 32 (143):941, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shreve:1974:AFT

[Shr74] David C. Shreve. L^{p} approximation of Fourier transforms and certain interpolating splines. Mathematics of Computation, 28(127):779–787, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1977:UPA

[Sid77] Avram Sidi. Uniqueness of Padé approximants from series of orthogonal polynomials. Mathematics of Computation, 31(139): 738–739, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1979:CPS

[Sid79] Avram Sidi. Convergence properties of some nonlinear sequence transformations. Mathematics of Computation, 33(145): 315–326, January 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sikorski:1979:TDA

[Sik79] Krzysztof Sikorski. A three-dimensional analogue to the method of bisections for solving nonlinear equations. Mathematics of Computation, 33(146):722–738, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Silvester:1970:SQF

[Sil70] P. Silvester. Symmetric quadrature formulae for simplexes. Mathematics of Computation, 24(109):95–100, January 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Simpson:1972:EEE

[Sim72] R. B. Simpson. Existence and error estimates for solutions of a discrete analog of nonlinear eigenvalue problems. Mathematics of Computation, 26(118):359–375, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sincovec:1972:SEA

[Sin72] Richard F. Sincovec. On the solution of the equations arising from collocation with cubic B-splines. Mathematics of Computation, 26(120):893–895, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shaw:1976:ACE

[SJR76] J. K. Shaw, L. W. Johnson, and R. D. Riess. Accelerating convergence of eigenfunction expansions. Mathematics of Computation, 30(135):469–477, July 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanno:1970:OCQ

[SK70] D. F. Shanno and P. C. Kettler. Optimal conditioning of quasi-Newton methods. Mathematics of Computation, 24 (111):657–664, July 1970. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Seidman:1972:EFI

[SK72] Thomas I. Seidman and Robert J. Korsan. Endpoint formulas for interpolatory cubic splines. Mathematics of Computation, 26(120):897–900, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skeel:1979:EFM

[Ske79] Robert D. Skeel. Equivalent forms of multistep formulas. Mathematics of Computation, 33(148):1229–1250, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skollermo:1975:FMN

[Sko75] Gunilla Skollermo. A Fourier method for the numerical solution of Poisson’s equation. Mathematics of Computation, 29 (131):697–711, July 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skollermo:1979:EAF

[Skö79] Gunilla Sköllermo. Error analysis of finite difference schemes applied to hyperbolic initial-boundary value problems. Mathematics of Computation, 33(145):11–35, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1972:BCR

[SL72] Daniel Shanks and Mohan Lal. Bateman’s constants reconsidered and the distribution of cubic residues. Mathematics of Computation, 26(117):265–285, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sarkany:1974:EFM

[SL74] E. F. Sarkany and W. Liniger. Exponential fitting of matricial multistep methods for ordinary differential equations. Mathematics of Computation, 28(128):1035–1052, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sloan:1976:IIC

[Slo76] Ian H. Sloan. Improvement by iteration for compact operator equations. Mathematics of Computation, 30(136):758–764, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smyth:1978:CRI

[Smy78] W. F. Smyth. The construction of rational iterating functions. Mathematics of Computation, 32(143):811–827, July 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Snider:1976:EAS

[Sni76a] Arthur David Snider. Error analysis for a stiff system procedure. Mathematics of Computation, 30(134):216– 219, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0398111-7.

Snider:1976:NAR

[Sni76b] Arthur David Snider. Numerical applications of reflection to partial differential equations. Mathematics of Computation, 30(134):220–240, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Srivastava:1975:SAA

[SP75] H. M. Srivastava and Rekha Panda. Some analytic or asymptotic confluent expansions for functions of several variables. Mathematics of Computation, 29(132): 1115–1128, October 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Singh:1975:CHM

[SPB75] Inderjit Singh, George Poole, and Thomas Boullion. A class of Hessenberg matrices with known pseudoinverse and Drazin inverse. Mathematics of Computation, 29(130):615–619, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1971:CGF

[Spi71] Robert Spira. Calculation of the gamma function by Stirling’s formula. Mathematics of Computation, 25(114):317–322, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1973:CRD

[Spi73] Robert Spira. Calculation of the Ramanujan τ-Dirichlet series. Mathematics of Computation, 27(122):379–385, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1976:ZHZ

[Spi76] Robert Spira. Zeros of Hurwitz zeta functions. Mathematics of Computation, 30 (136):863–866, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Steiner:1976:ABF

[SR76] Ray Steiner and Ronald Rudman. On an algorithm of Billevich for finding units in algebraic number fields. Mathematics of Computation, 30(135):598–609, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Solomon:1970:IVP

[SS70] Alan Solomon and Faiza Solomon. The initial-value problem for the equation (tu_{t})_{t} = u_{xx}. Mathematics of Computation, 24(111):611–620, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Srivastava:1972:UPC

[SS72] H. M. Srivastava and J. P. Singhal. A unified presentation of certain classical polynomials. Mathematics of Computation, 26(120):969–975, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1973:QFF

[SS73] Daniel Shanks and Richard Serafin.

Quadratic fields with four invariants divisible by 3. Mathematics of Computation, 27(121):183–187, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schoenberg:1974:SQF

[SS74] I. J. Schoenberg and S. D. Silliman. On semicardinal quadrature formulae. Mathematics of Computation, 28(126):483–497, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stevens:1978:CGV

[SS78] R. S. Stevens and R. Shantaram.

Computer-generated van der Waerden partitions. Mathematics of Computation, 32(142):635–636, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1970:DMN

[ST70] L. F. Shampine and R. J. Thompson. Difference methods for nonlinear first-order hyperbolic systems of equations. Mathematics of Computation, 24(109):45–56, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stahnke:1973:PBP

[Sta73a] Wayne Stahnke. Primitive binary polynomials. Mathematics of Computation, 27(124):977–980, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stauduhar:1973:DGG

[Sta73b] Richard P. Stauduhar. The determination of Galois groups. Mathematics of Computation, 27(124):981–996, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stark:1975:CQF

[Sta75] H. M. Stark. On complex quadratic fields with class-number two. Mathematics of Computation, 29(129):289–302, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1969:LMF

[Ste69] G. W. Stewart, III. On Lehmer’s method for finding the zeros of a polynomial. Mathematics of Computation, 23(108):829–835, s24–s30, October 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Ste71c].

Stephens:1971:FTC

[Ste71a] N. M. Stephens. On the Feit–Thompson conjecture. Mathematics of Computation, 25(115):625, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stepleman:1971:DAQ

[Ste71b] Robert S. Stepleman. Difference analogues of quasi-linear elliptic Dirichlet problems with mixed derivatives. Mathematics of Computation, 25(114): 257–269, April 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0303756-1.

Stewart:1971:CLM

[Ste71c] G. W. Stewart, III. Corrigendum: “On Lehmer’s Method for Finding the Zeros of a Polynomial”. Mathematics of Computation, 25(113):203, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Ste69].

Stewart:1971:EAA

[Ste71d] G. W. Stewart, III. Error analysis of the algorithm for shifting the zeros of a polynomial by synthetic division. Mathematics of Computation, 25(113):135–139, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1974:MPE

[Ste74] G. W. Stewart. Modifying pivot elements in Gaussian elimination. Mathematics of Computation, 28(126):537–542, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1975:GTG

[Ste75] G. W. Stewart. Gershgorin theory for the generalized eigenvalue problem Ax = λBx. Mathematics of Computation, 29 (130):600–606, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stepleman:1976:TFO

[Ste76] Robert S. Stepleman. Tridiagonal fourth order approximations to general two-point nonlinear boundary value problems with mixed boundary conditions. Mathematics of Computation, 30(133):92–103, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stenger:1979:SGM

[Ste79] Frank Stenger. A “sinc-Galerkin” method of solution of boundary value problems. Mathematics of Computation, 33(145):85– 109, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stothers:1978:FSF

[Sto78] W. W. Stothers. Free subgroups of the free product of cyclic groups. Mathematics of Computation, 32(144):1274–1280, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Strecok:1971:TEB

[Str71] Anthony J. Strecok. Table errata: Handbook of mathematical functions with formulas, graphs, and mathematical tables (Nat. Bur. Standards, Washington, D. C., 1964), edited by Milton Abramowitz and Irene A. Stegun. Mathematics of Computation, 25(114):405, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/journals/mcom/1971-25-114/
                                              S0025-5718-1971-0415963-8.

Strom:1972:MRQ

[Str72] Torsten Strom. Monotonicity in Romberg quadrature. Mathematics of Computation, 26(118):461–465, April 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Straus:1975:CTG

[Str75] E. G. Straus. A combinatorial theorem in group theory. Mathematics of Computation, 29(129):303–309, January 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Streit:1976:TSM

[Str76a] R. F. Streit. The T^{+m} transformation. Mathematics of Computation, 30 (135):503–511, July 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Streit:1976:T

[Str76b] Roland F. Streit. The T_{+m} transformation. Mathematics of Computation, 30 (135):505–511, July 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1976:SFD

[Str76c] A. H. Stroud. Some fourth degree integration formulas for simplexes. Mathematics of Computation, 30(134):291– 294, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0391484-0.

Stynes:1979:FCB

[Sty79] Martin Stynes. On faster convergence of the bisection method for certain triangles. Mathematics of Computation, 33(146):717–721, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1971:CEE

[SW71] L. F. Shampine and H. A. Watts. Comparing error estimators for Runge–Kutta methods. Mathematics of Computation, 25 (115):445–455, July 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sag:1973:CMN

[SW73] T. W. Sag and J. W. Wamsley. On computing the minimal number of defining relations for finite groups. Mathematics of Computation, 27(122):361–368, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1974:BC

[SW74] Daniel Shanks and John W. Wrench, Jr. Brun’s constant. Mathematics of Computation, 28(125):293–299, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1977:IMN

[SW77] A. H. Schatz and L. B. Wahlbin. Interior maximum norm estimates for finite element methods. Mathematics of Computation, 31(138):414–442, April 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1978:MNE

[SW78] A. H. Schatz and L. B. Wahlbin. Maximum norm estimates in the finite element method on plane polygonal domains. part 1. Mathematics of Computation, 32(141): 73–109, January 1978. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1979:MNE

[SW79a] A. H. Schatz and L. B. Wahlbin. Maximum norm estimates in the finite element method on plane polygonal domains. part 2, refinements. Mathematics of Computation, 33(146):465–492, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schmidt:1979:MAS

[SW79b] Darrell Schmidt and Kenneth L. Wiggins.

Minimax approximate solutions of linear boundary value problems. Mathematics of Computation, 33(145):139–148, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Steihaug:1979:AAE

[SW79c] Trond Steihaug and Arne Wolfbrandt. An attempt to avoid exact Jacobian and nonlinear equations in the numerical solution of stiff differential equations. Mathematics of Computation, 33(146):521–534, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stepleman:1979:AND

[SW79d] R. S. Stepleman and N. D. Winarsky.

Adaptive numerical differentiation. Mathematics of Computation, 33(148):1257– 1264, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swarztrauber:1979:PAS

[Swa79] Paul N. Swarztrauber. A parallel algorithm for solving general tridiagonal equations. Mathematics of Computation, 33 (145):185–199, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Takacs:1978:SBC

[Tak78] Lajos Takács. A sum of binomial coefficients. Mathematics of Computation, 32 (144):1271–1273, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1973:NCD

[te 73] H. J. J. te Riele. A note on the Catalan– Dickson conjecture. Mathematics of Computation, 27(121):189–192, January 1973.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [CS73].

teRiele:1974:FLA

[te 74] H. J. J. te Riele. Four large amicable pairs. Mathematics of Computation, 28 (125):309–312, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tuan:1972:CSE

[TE72] P. D. Tuan and David Elliott. Coefficients in series expansions for certain classes of functions. Mathematics of Computation, 26(117):213–232, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temme:1975:UAE

[Tem75] N. M. Temme. Uniform asymptotic expansions of the incomplete gamma functions and the incomplete beta function. Mathematics of Computation, 29(132):1109– 1114, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thacher:1977:ACC

[Tha77] Henry C. Thacher. Book review: Applied and Computational Complex Analysis, by Peter Henrici. Mathematics of Computation, 31(137):325–326, January 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

/www.jstor.org/stable/pdf/2005805.
pdf.

Thacher:1979:NBR

[Tha79] Henry C. Thacher, Jr. New backward recurrences for Bessel functions. Mathematics of Computation, 33(146):744–764, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomas:1976:PIS

[Tho76] D. H. Thomas. Pseudospline interpolation for space curves. Mathematics of Computation, 30(133):58–67, January 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomee:1977:HOL

[Tho77] Vidar Thomee. High order local approximations to derivatives in the finite element method. Mathematics of Computation, 31 (139):652–660, July 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomee:1979:SIE

[Tho79] Vidar Thomée. Some interior estimates for semidiscrete Galerkin approximations for parabolic equations. Mathematics of Computation, 33(145):37–62, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Todenes:1970:NSD

[Tod70] Øystein Todenes. On the numerical solution of the diffusion equation. Mathematics of Computation, 24(111):621–627, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Toint:1977:SSM

[Toi77] Ph. L. Toint. On sparse and symmetric matrix updating subject to a linear equation. Mathematics of Computation, 31(140):954–961, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Toint:1978:SNR

[Toi78] Ph. L. Toint. Some numerical results using a sparse matrix updating formula in unconstrained optimization. Mathematics of Computation, 32(143):839–851, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Trench:1974:ITB

[Tre74] William F. Trench. Inversion of Toeplitz band matrices. Mathematics of Computation, 28(128):1089–1095, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Troesch:1973:EMS

[Tro73] B. Andreas Troesch. Elliptical membranes with smallest second eigenvalue. Mathematics of Computation, 27(124):767–772, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Troesch:1973:EEM

[TT73] B. A. Troesch and H. R. Troesch. Eigenfrequencies of an elliptic membrane. Mathematics of Computation, 27(124):755–765, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tuan:1971:RQR

[Tua71] P. D. Tuan. On the remainder in quadrature rules. Mathematics of Computation, 25(116):819–826, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tuckerman:1973:SPL

[Tuc73] Bryant Tuckerman. A search procedure and lower bound for odd perfect numbers. Mathematics of Computation, 27 (124):943–949, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tuckerman:1977:C

[Tuc77] Bryant Tuckerman. Corrigendum. Mathematics of Computation, 31(140):1051, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tung:1975:LUB

[Tun75] S. H. Tung. On lower and upper bounds of the difference between the arithmetic and the geometric mean. Mathematics of Computation, 29(131):834–836, July 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Turkel:1973:SFD

[Tur73] Eli Turkel. Symmetrization of the fluid dynamic matrices with applications. Mathematics of Computation, 27(124):729–736, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Turan:1975:PSM

[Tur75] Paul Turan. Power sum method and the approximative solution of algebraic equations. Mathematics of Computation, 29 (129):311–318, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Taylor:1974:EQP

[TW74a] G. D. Taylor and J. Williams. Existence questions for the problem of Chebyshev approximation by interpolating rationals. Mathematics of Computation, 28 (128):1097–1103, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomee:1974:CRP

[TW74b] Vidar Thomee and Lars Wahlbin. Convergence rates of parabolic difference schemes for non-smooth data. Mathematics of Computation, 28(125):1–13, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Usmani:1978:DVM

[Usm78] Riaz A. Usmani. Discrete variable methods for a boundary value problem with engineering applications. Mathematics of Computation, 32(144):1087–1096, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vaillancourt:1970:SFS

[Vai70] Remi Vaillancourt. On the stability of Friedrich’s scheme and the modified Lax– Wendroff scheme. Mathematics of Computation, 24(112):767–770, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanLinde:1974:HOF

[van74] H. J. van Linde. High-order finitedifference methods for Poisson’s equation. Mathematics of Computation, 28(126): 369–391, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1970:CIS

[Var70a] J. M. Varah. Computing invariant subspaces of a general matrix when the eigensystem is poorly conditioned. Mathematics of Computation, 24(109):137–149, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1970:MNS

[Var70b] J. M. Varah. Maximum norm stability of difference approximations to the mixed initial boundary-value problem for the heat equation. Mathematics of Computation, 24(109):31–44, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1972:SBT

[Var72] J. M. Varah. On the solution of blocktridiagonal systems arising from certain finite-difference equations. Mathematics of Computation, 26(120):859–868, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1974:CSN

[Var74] James M. Varah. A comparison of some numerical methods for two-point boundary value problems. Mathematics of Computation, 28(127):743–755, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1977:CNL

[Var77] J. M. Varah. On the condition number of local bases for piecewise cubic polynomials. Mathematics of Computation, 31(137): 37–44, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1979:EII

[Var79] J. M. Varah. On the efficient implementation of implicit Runge–Kutta methods. Mathematics of Computation, 33(146): 557–561, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vaughan:1978:GSC

[Vau78] Theresa P. Vaughan. A generalization of the simple continued fraction algorithm. Mathematics of Computation, 32(142): 537–558, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Venit:1973:MRP

[Ven73] Stewart Venit. Mesh refinements for parabolic equations of second order. Mathematics of Computation, 27(124):745–754, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1972:FSI

[Wag72] Samuel S. Wagstaff, Jr. On k-free sequences of integers. Mathematics of Computation, 26(119):767–771, July 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1975:ZAF

[Wag75] Samuel S. Wagstaff, Jr. Zeros of p-adic Lfunctions. Mathematics of Computation, 29(132):1138–1143, October 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1978:IP

[Wag78] Samuel S. Wagstaff, Jr. The irregular primes to 125000. Mathematics of Computation, 32(142):583–591, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1979:GLP

[Wag79a] Samuel S. Wagstaff, Jr. Greatest of the least primes in arithmetic progressions having a given modulus. Mathematics of Computation, 33(147):1073–1080, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1979:SNE

[Wag79b] Samuel S. Wagstaff, Jr. Solution of Nathanson’s exponential congruence. Mathematics of Computation, 33(147): 1097–1100, July 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wahlbin:1975:MGP

[Wah75] Lars Wahlbin. A modified Galerkin procedure with Hermite cubics for hyperbolic problems. Mathematics of Computation, 29(132):978–984, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wall:1972:DBE

[Wal72] Charles R. Wall. Density bounds for Euler’s function. Mathematics of Computation, 26(119):779–783, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Walker:1978:IS

[Wal78] P. L. Walker. On an integral summable to 2ξ(s)/s(s − 1). Mathematics of Computation, 32(144):1311–1316, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wang:1976:FMP

[Wan76] Paul S. Wang. Factoring multivariate polynomials over algebraic number fields. Mathematics of Computation, 30(134): 324–336, April 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0568283-X.

Wang:1978:IMP

[Wan78] Paul S. Wang. An improved multivariate polynomial factoring algorithm. Mathematics of Computation, 32(144):1215– 1231, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ward:1976:AHM

[War76] Robert C. Ward. The QR algorithm and Hyman’s method on vector computers. Mathematics of Computation, 30(133): 132–142, January 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Watson:1973:CBL

[Wat73] G. A. Watson. The calculation of best linear one-sided L^{p} approximations. Mathematics of Computation, 27(123):607–620, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Watanabe:1978:BIO

[Wat78] Daniel S. Watanabe. Block implicit onestep methods. Mathematics of Computation, 32(142):405–414, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1976:CTE

[WB76] H. C. Williams and J. Broere. A computational technique for evaluating L(1, chi ) and the class number of a real quadratic field. Mathematics of Computation, 30 (136):887–893, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1979:CRU

[WB79] H. C. Williams and P. A. Buhr. Calculation of the regulator of Q(\sqrt{D}) by use of the nearest integer continued fraction algorithm. Mathematics of Computation, 33(145):369–381, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Warming:1975:DSS

[WBH75] R. F. Warming, Richard M. Beam, and B. J. Hyett. Diagonalization and simultaneous symmetrization of the gas-dynamic matrices. Mathematics of Computation, 29 (132):1037–1045, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wall:1972:DBS

[WCJ72] Charles R. Wall, Phillip L. Crews, and Donald B. Johnson. Density bounds for the sum of divisors function. Mathematics of Computation, 26(119):773–777, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1974:CSA

[Wd74] Jack Williams and Frank de Hoog. A class of A-stable advanced multistep methods. Mathematics of Computation, 28 (125):163–177, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Webb:1973:NRW

[Web73] William A. Webb. Numerical results for Waring’s problem in GF[q, x]. Mathematics of Computation, 27(121):193–196, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weinstein:1972:UAT

[Wei72a] S. E. Weinstein. Uniform approximation through partitioning. Mathematics of Computation, 26(118):493–503, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weiss:1972:PIG

[Wei72b] Richard Weiss. Product integration for the generalized Abel equation. Mathematics of Computation, 26(117):177–190, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weintraub:1974:CPL

[Wei74a] Sol Weintraub. A compact prime listing. Mathematics of Computation, 28 (127):855–857, July 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weiss:1974:AIR

[Wei74b] Richard Weiss. The application of implicit Runge–Kutta and collocation methods to boundary-value problems. Mathematics of Computation, 28(126):449–464, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weinberger:1975:SZD

[Wei75a] Peter J. Weinberger. On small zeros of Dirichlet L-functions. Mathematics of Computation, 29(129):319–328, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weiss:1975:BDA

[Wei75b] Richard Weiss. Bifurcation in difference approximations to two-point boundary value problems. Mathematics of Computation, 29(131):746–760, July 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weintraub:1977:SPA

[Wei77] Sol Weintraub. Seventeen primes in arithmetic progression. Mathematics of Computation, 31(140):1030, October 1977. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Watanabe:1974:IFO

[WF74] Daniel S. Watanabe and J. Richard Flood. An implicit fourth order difference method for viscous flows. Mathematics of Computation, 28(125):27–32, January 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1977:CS

[WH77] H. C. Williams and R. Holte. Computation of the solution of x^{3}+Dy^{3} = 1. Mathematics of Computation, 31(139):778–785, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1978:SOP

[WH78] H. C. Williams and R. Holte. Some observations on primality testing. Mathematics of Computation, 32(143):905–917, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Widlund:1971:ESP

[Wid71] Olof B. Widlund. On the effects of scaling of the Peaceman–Rachford method. Mathematics of Computation, 25(113):33– 41, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wiederholt:1976:SMM

[Wie76] Lawrence F. Wiederholt. Stability of multistep methods for delay differential equations. Mathematics of Computation, 30(134):283–290, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0398132-4.

Wilson:1970:DLS

[Wil70] M. Wayne Wilson. Discrete least squares and quadrature formulas. Mathematics of Computation, 24(110):271–282, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1972:NCA

[Wil72] Jack Williams. Numerical Chebyshev approximation by interpolating rationals. Mathematics of Computation, 26(117): 199–206, January 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1974:QPP

[Wil74] Kenneth S. Williams. A quadratic partition of primes ≡ 1(mod 7). Mathematics of Computation, 28(128):1133–1136, October 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilhelmsen:1976:NPA

[Wil76a] Don R. Wilhelmsen. A nearest point algorithm for convex polyhedral cones and applications to positive linear approximation.

Mathematics of Computation, 30(133):48– 57, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Willett:1976:CS

[Wil76b] Michael Willett. Characteristic msequences. Mathematics of Computation, 30(134):306–311, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1976-30-134/
S0025-5718-1976-0389863-0.

Williams:1976:ECQ

[Wil76c] Kenneth S. Williams. Explicit criteria for quintic residuacity. Mathematics of Computation, 30(136):847–853, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1977:CPC

[Wil77] H. C. Williams. Certain pure cubic fields with class-number one. Mathematics of Computation, 31(138):578–580, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See errata [Wil79].

Williams:1978:SPI

[Wil78] H. C. Williams. Some primes with interesting digit patterns. Mathematics of Computation, 32(144):1306–1310, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1979:ECP

[Wil79] H. C. Williams. Errata: “Certain pure cubic fields with class-number one” [Math. Comp. 31 (1977), no. 138, 578–580, MR 55 #5578]. Mathematics of Computation, 33(146):847–848, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Wil77].

Wimp:1962:PEB

[Wim62] Jet Wimp. Polynomial expansions of Bessel functions and some associated functions. Mathematics of Computation, 16 (80):446–458, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Wim72a].

Wimp:1972:CPE

[Wim72a] Jet Wimp. Corrigendum: Polynomial expansions of Bessel functions and some associated functions (Math. Comp. 16 (1962), 446–458). Mathematics of Computation, 26(117):309, January 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1972-26-117/
S0025-5718-1972-0400659-X;   http:
//www.ams.org/journals/mcom/1972-
26-117/S0025-5718-1972-0400659-X/
S0025-5718-1972-0400659-X.pdf; https:
//www.ams.org/mathscinet-getitem?
mr=400659;   https://www.ams.org/
mathscinet/search/authors.html?authorName=
Wimp%2C%20Jet. See [Wim62].

Wimp:1972:STM

[Wim72b] Jet Wimp. Some transformations of monotone sequences. Mathematics of Computation, 26(117):251–254, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wimp:1975:DDP

[Wim75] Jet Wimp. Differential-difference properties of hypergeometric polynomials. Mathematics of Computation, 29(130):577–581, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Winslow:1970:BID

[Win70] Leon Winslow. “best” interpolation, differentiation, and integration approximations on the Hardy space H^{2}. Mathematics of Computation, 24(111):523–527, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Winograd:1978:CDF

[Win78] S. Winograd. On computing the discrete Fourier transform. Mathematics of Computation, 32(141):175–199, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1976:DPU

[WJ76a] H. C. Williams and J. S. Judd. Determination of the primality of N by using factors of N^{2} ± 1. Mathematics of Computation, 30(133):157–172, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1976:SAP

[WJ76b] H. C. Williams and J. S. Judd. Some algorithms for prime testing using generalized Lehmer functions. Mathematics of Computation, 30(136):867–886, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wolfe:1975:DRA

[Wol75] Jerry M. Wolfe. Discrete rational L_{p} approximation. Mathematics of Computation, 29(130):540–548, April 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wong:1973:AEL

[Won73] R. Wong. An asymptotic expansion of W_{k,m}(z) with large variable and parameters. Mathematics of Computation, 27(122):429–436, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wong:1975:LTN

[Won75] R. Wong. On Laplace transforms near the origin. Mathematics of Computation, 29(130):573–576, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wong:1971:SEI

[WR71] R. Wong and E. Rosenbloom. Series expansions of W_{k,m}(z) involving parabolic cylinder functions. Mathematics of Computation, 25(116):783–787, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wang:1975:FMP

[WR75] Paul S. Wang and Linda Preiss Rothschild. Factoring multivariate polynomials over the integers. Mathematics of Computation, 29(131):935–950, July 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrench:1973:ECT

[Wre73] John W. Wrench. Erratum: Concerning Two Series for the Gamma Function. Mathematics of Computation, 27 (123):681–682, July 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrigge:1973:EII

[Wri73] H. S. Wrigge. An elliptic integral identity. Mathematics of Computation, 27 (124):839–840, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1979:SPF

[WS79a] H. C. Williams and E. Seah. Some primes of the form (a^{n}−1)/(a−1). Mathematics of Computation, 33(148):1337–1342, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1979:NCN

[WS79b] H. C. Williams and Daniel Shanks. A note on class-number one in pure cubic fields. Mathematics of Computation, 33 (148):1317–1320, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wunderlich:1973:GPL

[Wun73] M. C. Wunderlich. On the Gaussian primes on the line IM(X) = 1. Mathematics of Computation, 27(122):399–400, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wuytack:1975:ORI

[Wuy75] Luc Wuytack. On the osculatory rational interpolation problem. Mathematics of Computation, 29(131):837–843, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1972:SPN

[WZ72] H. C. Williams and C. R. Zarnke. Some prime numbers of the forms 2A3^{n} + 1 and 2A3^{n} − 1. Mathematics of Computation, 26(120):995–998, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wolkowicz:1978:CBA

[WZ78] H. Wolkowicz and S. Zlobec. Calculating the best approximate solution of an operator equation. Mathematics of Computation, 32(144):1183–1213, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1971:HMC

[Yan71] C. H. Yang. On Hadamard matrices constructible by circulant submatrices. Mathematics of Computation, 25(113):181–186, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1976:MBM

[Yan76] C. H. Yang. Maximal binary matrices and sum of two squares. Mathematics of Computation, 30(133):148–153, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1979:HMF

[Yan79] C. H. Yang. Hadamard matrices, finite sequences, and polynomials defined on the unit circle. Mathematics of Computation, 33(146):688–693, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Young:1970:CPS

[You70] David M. Young. Convergence properties of the symmetric and unsymmetric successive overrelaxation methods and related methods. Mathematics of Computation, 24(112):793–807, October 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zwas:1971:TFO

[ZA71] Gideon Zwas and Saul Abarbanel. Third and fourth order accurate schemes for hyperbolic equations of conservation law form. Mathematics of Computation, 25 (114):229–236, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/journals/mcom/1971-25-114/
S0025-5718-1971-0303766-4.

Zaremba:1976:CPG

[Zar76] S. K. Zaremba. On Cartesian products of good lattices. Mathematics of Computation, 30(135):546–552, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zassenhaus:1978:RHF

[Zas78] Hans Zassenhaus. A remark on the Hensel factorization method. Mathematics of Computation, 32(141):287–292, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zill:1970:SEI

[ZC70] D. G. Zill and B. C. Carlson. Symmetric elliptic integrals of the third kind. Mathematics of Computation, 24(109):199–214, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zakian:1978:TCF

[ZE78] V. Zakian and M. J. Edwards. Tabulation of constants for full grade I_{MN} approximants. Mathematics of Computation, 32(142):519–531, April 1978. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zlamal:1974:FEM

[Zla74] Milos Zlamal. Finite element methods for parabolic equations. Mathematics of Computation, 28(126):393–404, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zlamal:1975:FEM

[Zla75] Milos Zlamal. Finite element multistep discretizations of parabolic boundary value problems. Mathematics of Computation, 29(130):350–359, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zlamal:1978:SRI

[Zla78] Milos Zlamal. Superconvergence and reduced integration in the finite element method. Mathematics of Computation, 32 (143):663–685, July 1978. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zajta:1975:CCF

[ZP75] A. J. Zajta and W. Pandikow. Conversion of continued fractions into power series.

Mathematics of Computation, 29(130): 566–572, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zwillinger:1979:GCU

[Zwi79] Dan Zwillinger. A Goldbach conjecture using twin primes. Mathematics of Computation, 33(147):1071, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).